### v1

Abstract
Library migration is the process of replacing a library with a similar one in a software project. Manual library migration is time consuming and error prone, as it requires developers to understand the Application Programming Interfaces (API) of both libraries, map equivalent APIs, and perform the necessary code transformations. Due to the difficulty of the library migration process, most of the existing automated techniques and tooling stop at the API mapping stage or support a limited set of libraries and code transformations. In this paper, we develop an end-to-end solution that can automatically migrate code between any arbitrary pair of Python libraries that provide similar functionality. Due to the promising capabilities of Large Language Models (LLMs) in code generation and transformation, we use LLMs as the primary engine for migration. Before building the tool, we first study the capabilities of LLMs for library migration on a benchmark of 321 real-world library migrations. We find that LLMs can effectively perform library migration, but some post-processing steps can further improve the performance. Based on this, we develop PyMigTool, a command line application that combines the power of LLMs, static analysis, and dynamic analysis to provide accurate library migration. We evaluate PyMigTool on 717 real-world Python applications that are not from our benchmark. We find that PyMigTool can migrate 32% of the migrations with complete correctness. Of the remaining migrations, only 14% of the migration-related changes are left for developers to fix for more than half of the projects.
1 Introduction
Library migration is the process of replacing a library with an alternative one in a software project. The process involves identifying the library usage in the codebase, finding the corresponding Application Programming Interfaces (APIs) in the new library, and replacing the old API usage with the new API usage. This can be tedious and error-prone, especially when the library is extensively used in a large codebase kula2018developers .
Researchers have proposed various techniques to automate the library migration processes. Most of these techniques, however, are limited to only finding the analogous APIs (API mapping) teyton2013automatic ; alrubaye2018automating ; alrubaye2019use ; miningAnalogicalAPIs ; zhang2020deep ; di2025deepmig , but do not replace the old API usage in the codebase with the new one (transformation). To find analogous APIs, some of these techniques mine existing migrations teyton2013automatic ; alrubaye2018automating ; alrubaye2019use while others use machine learning models miningAnalogicalAPIs ; di2025deepmig and natural language processing techniques ni2021soar . To the best of our knowledge, there are two techniques that perform transformation ni2021soar ; nikolov2025google but both techniques evaluate their approaches on a small set of library pairs without providing a tool that can be used in practice.
In recent years, Large Language Models (LLMs) have shown that they can be effective in various software engineering tasks ozkaya2023application ; fan2023large ; wang2024software , including the necessary upstream tasks for library migration, such as code generation nguyen2022empirical ; peng2023impact ; murali2023codecompose , code comprehension nam2024using , and code transformation wei2023copiloting ; xia2023automated ; zhou2023hybrid . This suggests that LLMs may be capable of performing library migration.
In this paper, we leverage LLMs to explore and address the challenges of automating library migration in two phases. First, we conduct an experiment using an existing library migration benchmark, PyMigBench pymigbench , to determine the capability and challenges of using LLMs for library migration. We find that LLMs are generally capable of performing migration; however, some post processing can further improve the migration results. Second, based on our findings in first phase, we build a CLI tool, PyMigTool, that automates the full library migration process by combining the power of LLMs with post processing using static and dynamic program analysis. We run an extensive evaluation of PyMigTool on a set of 717 new, simulated migrations.
Phase one – Explore LLMs:
In this phase, we investigate the capabilities and challenges of using LLMs for library migration. We use LLama 3.1, GPT-4o mini, and GPT-4o to perform 321 migrations from PyMigBench. We then measure the correctness of the migrations by comparing the LLM’s migrations with the original developer’s migrations. We find that the three LLMs can correctly migrate 26%, 49%, and 57% of the migrations respectively. Overall, we find that LLMs are capable of library migration, and as expected, the more powerful LLM (GPT-4o) performs better than the less powerful LLMs (LLama 3.1 and GPT-4o mini). We also identify two major challenges that make solely relying on LLMs in a practical tool ineffecive: (1) We provide the LLM with only code files that directly use the source library. However, for a given project, there may be other files that depend on the migrated code and also require changes during the migration. (2) All LLMs often return only part of the code, skipping parts of the code that do not require migration. These limitations, however, can be addressed by post-processing the LLM’s results. We use these insights to improve the migration process in the next phase.
Phase two – Build PyMigTool:
In this phase, we build PyMigTool, a Command Line Interface (CLI) that automates the full library migration process using LLMs and post processing using static and dynamic program analysis. PyMigTool takes as 5 inputs from the user (1) A project that requires migration, with an existing test suite with high coverage. (2) A requirements file that lists the dependencies of the project. (3) A source library that is used in the project. (4) A target library and its version that the user wants to migrate to. (5) Name of an LLM that will be used for the migration. By building a dynamic call graph, PyMigTool identifies all files that require migration and sends those to the LLM. PyMigTool then verifies the results of the migration by comparing the test results before and after migration. If any previously passing tests now fail, PyMigTool performs a Post-processing round where it checks if the LLM skipped any code and merges it back. It also detects if any of functions were made async during the migration, and if so, it finds other parts of the project that uses these functions and makes them async as well. Finally, it runs the unit tests again and saves the results
We evaluate PyMigTool using a new purely test-based evaluation setup, as opposed to the benchmark based evaluation we did in the first phase. This allows us to evaluate new migrations without the need for a ground truth developer migration. We evaluate PyMigTool on a set of 717 simulated migrations in 175 unique projects. By simulated migrations, we mean that this migration did not exist in the project’s history; instead, we find source libraries used in a project and migrate the project to an analogous target library. We ensure that the projects we choose have high test coverage, so that the evaluation is reliable.
We demonstrate twi different evaluation setups. First, we measure how well PyMigTool perform as it is on the simulated migrations, without any manual intervention. While the best case scenario is that PyMigTool can migrate the project without any issues, a more practical expectation is that PyMigTool would do most of the task correctly, but there are some parts that require manual intervention. Therefore, in a second evaluation, we measure how much manual work can fix the incorrect migrations. We randomly select a subset of 43 diverse migrations that were not fully correct, and we manually fix them.
We find that PyMigTool was able to successfully migrate 32% migrations without any manual intervention. We manually fix 43 of the remaining 484 migrations, and were able to fix 21 (49%) of them. In terms of modified lines of code, the manual work is only 14% of the total migration-related changes.
To summarize, the contributions of this paper are:
  1. 1.
    We conduct an initial detailed experiment to explore the use of LLMs for library migration on an existing migration benchmark that provides us a ground truth of previously performed developer migrations.
  2. 2.
    Based on the insights gained from our initial experiment, we build a CLI tool, PyMigTool, that automates the full library migration process using a combination of LLMs and post processing using static and dynamic program analysis.
  3. 3.
    We run an extensive test-based evaluation of PyMigTool on a set of 717 simulated migrations, and show that it works well with no or minimal manual editing for a diverse set of library pairs.
2 Background and Terminology
In this section, we provide background information and terminology used in the rest of the paper.
2.1 Library Migration
Library migration is the process of updating a software project to replace a used library with another one that provides similar functionality teyton2012mining . The source library is the one being replaced, and the target library is the one that replaces it teyton2012mining . One migration instance refers to a commit in a repository where a migration happened from a specific source library to a target library pymigbench , denoted using the notation source→target. For example, the commit b0607a26 in repository openstack/ironic is a retrying→tenacity migration.
A migration-related code change, or code change for brevity, is a minimal replacement of source library APIs with target library APIs that cannot be meaningfully reduced further without losing the semantics of the change pymigbench . A migration contains one or more code changes. The lines between the boxes in 1(c) show code changes. For example, segment P1 in 1(a) is replaced by segment D1 in 1(b).
We use subscripts pre, dev, and llm to denote data before migration, after developer’s migration, and after an LLM’s migration, respectively. For example, codepre, codedev, and codellm denote three states of a code. Changedev and Changellm denote code changes by developers and an LLM, respectively.
2.2 PyMigBench and PyMigTax
PyMigBench is a dataset of real-world Python library migrations pymigbench , mined from version control history. We use the latest available version, version 2.2.5 pymigtax in our experiments. The dataset has 321 migrations and 2,989 migration-related code changes. It includes a detailed description of each code change, including line numbers and API names, which facilitates our evaluation.
PyMigTax is a taxonomy of migration-related code changes pymigtax , built using the first version of PyMigBench. The authors validated its generalizability on additional third-party data. PyMigTax describes a code change based on three dimensions: (1) Program elements: the types of program elements involved in the change (e.g., function call and attribute); (2) Cardinality: how many source APIs are replaced with how many target APIs. For example, one-to-many cardinality means one source API is replaced with multiple target APIs; one-to-many, many-to-one, and many-to-many cardinalities are commonly referred to as higher-cardinality; and (3) Properties: Additional properties to describe the code change (e.g., element name change when the source and target APIs have different names.). The code change data in PyMigBench is annotated with the PyMigTax categories. We use PyMigTax categories (shown in Table 4) to understand the types of code changes LLMs can handle. The PyMigTax paper pymigtax has full category descriptions.
3 Phase one: Exploring the capabilities of LLMs for Python library migration
We use three off-the-shelf LLMs – LLama 3.1 llamapaper , GPT-4o mini openaimodels , and GPT-4o openaimodels – to perform migrations on the PyMigBench dataset. We refer to these models as LLama, Mini, and 4o in the rest of the paper, respectively. We evaluate the correctness of the migrations in two ways. First, we compare the LLM-migrated code with the developer-migrated code in PyMigBench and report correctness at two granularity levels: the full migration and the individual code changes. We use PyMigTax to identify which types of code changes the LLMs can/cannot handle. Second, for the subset of migrations with available unit tests, we run the available tests to assess the correctness of the migrated code.
We answer the following research questions in this phase:
  1. RQ-A1
    How similar are the LLM migrations to the benchmark migrations? We consider the code changes stored in PyMigBench as the ground truth. We automatically check if the LLM was able to correctly perform all expected changes, while accounting for refactoring and alternative correct changes through manual review.
  2. RQ-A2
    How many migrations pass unit tests? To evaluate run-time correctness of the migrated code, we use a second evaluation strategy where we run any available unit tests.
3.1 Experiment Setup
3.1.1 Models
Table 1: Description of the models we use in our experiment.
LLama 3.1 GPT-4o mini GPT-4o
Used version 3.1-70B-Instruct 2024-07-18 2024-08-06
Knowledge cutoff December 2023 October 2023 October 2023
Model Type Open-source Proprietary Proprietary
# training parameters 70 billion No official info No official info
Context length 128K tokens 128K tokens 128K tokens
Output limit 4096 tokens  [^1] LLama 3.1 has an input+output limit of 8192 tokens. Our prompt and response are approximately equal, therefore we consider the output limit to be half (4096). 16,384 tokens 16,384 tokens
Cost  [^2] Cost per million input token at the time of our experiments: July and August 2024 Free \$0.15 \$2.5
Table 1 shows the three off-the-shelf models we use for our library migration task. The first model is Meta’s LLama 3.1, a free and open-source LLM llamapaper . The other two models are proprietary: OpenAI GPT-4o mini and GPT-4o openaimodels . GPT-4o mini replaces OpenAI’s previous GPT-3.5 Turbo and is more cost-effective. GPT-4o is the most powerful of the OpenAI models, while also being the most expensive openaimodels . We use these models to migrate the code in 321 migrations from our PyMigBench dataset (pymigbench ).
3.1.2 Data preparation
At the time of conducting this experiment, two repositories from PyMigBench were no longer public, preventing us from using 2 migrations. We clone the remaining repositories containing 319 migrations. For each migration commit, we use PyDriller PyDriller to extract the content of each file that has recorded migration-related code changes, both before and after the migration. We use Python’s built-in ast.parse function docast to ensure that the files are syntactically valid code and discard 5 migrations having files with syntax errors. We conduct our experiments with the remaining 314 migrations containing 2,910 migration-related code changes.
3.1.3 Migration
For each file having code changes, we use the LLM prompt template in 1 to migrate the code. The prompt is designed to ensure that the LLM performs focused, controlled migrations while maintaining transparency. By asking the LLM to explain its changes, we ensure that the LLM justifies its modifications, which can be useful for a human evaluator reviewing the migration. The prompt also aims to restrict the LLM from making unrelated modifications such that we remain focused on the task of library migration.
Since LLMs are non-deterministic atil2024llmstabilitydetailedanalysis , we use a temperature of 0 and run each migration 10 times for each model. We calculate the difference between the runs to understand the variability in the generated code. We use git-diff gitdiff to compute a diff between each pair of the 10 runs and normalize the number of different lines by dividing it by the total number of lines in Filepre. We find that on average, two runs on the same file are only 6.6%, 4.0% and 4.8% different for LLama, Mini, and 4o, respectively. Given this low variability and the high effort involved in manual validation, we randomly select one run for each model to evaluate the LLMs’ results. Our artifact contains the LLM’s migrations for all 10 runs.
3.2 RQ-A1 How similar are the LLM migrations to the benchmark migrations?
3.2.1 RQ-A1 Approach
In this RQ, we assess each LLM’s migration correctness by comparing its code changes (Changellm) to those made by developers (Changedev), while manually judging potential alternative changes. A migration may involve multiple files, each with multiple migration-related code changes. We evaluate each code change individually, and aggregate the results to determine the correctness of the whole migration. We now explain the process using the example in 1(c) that shows a migration from the library requests to aiohttp.
Listing 1: LLM Migration Prompt
1The following Python code uses library <source-lib>.
2Migrate this code to use library <target-lib> instead.
3In the output, first explain the changes you made. Then, provide the modified code.
4
5Do not make any changes to the code that are not related to migrating between these two libraries.
6Do not refactor. Do not reformat. Do not optimize. Do not change coding style.
7
8Provided code:
9<premig-code>
(a) Before migration (Filepre)
(b) Developer’s migration (Filedev)
(c) LLM Migration (Filellm)
Figure 1: A sample requests→aiohttp migration.
The lines between the boxes show matching code changes.  denotes automatic matching, while  denotes manual matching. Dashed lines denote changes that are not recorded in PyMigBench. The blue check marks denote correct changes, while the orange crosses denote incorrect changes.
3.2.1.1 Match code changes
The goal of this step is to compare the code changes the LLM made (Changellm) with the developer ground truth changes (Changedev), as well as manually finding alternative correct changes.
Auto change matching: We first attempt to automatically match as many changes as possible while ensuring precision i.e., not falsely marking a Changellm as correct. Therefore, we consider only exact syntactic matches between Changedev and Changellm, ignoring formatting differences. We find exact matches by identifying the AST nodes related to Changedev in Filedev and then matching them to corresponding nodes in Filellm. If there are potentially multiple matches, we check the containing function and use proximity based heuristics demonstrated below. We ensure that the matched node sets translate to the same code string using ast.unparse docast .
Consider Changedev D1 in 1(c) which uses ClientSession() and get() from aiohttp. In Filellm, these APIs appear in multiple locations, but only Lines 15, 16, 21, and 22 are in the same containing function fetch_flights as Changedev. Based on line number proximity, we identify two node sets L2 and L4 as potential matches for this Changedev. By comparing the code strings, we find that L2 exactly matches Changedev. By the end of the auto change matching step, we get a set of matched pairs of Changedev and Changellm, a set of unmatched Changedev (uChangedev), and a set of unmatched Changellm (uChangellm).
Manual review: The manual review step focuses on reviewing uChangedev and uChangellm. We use a three-way diff viewer to manually compare Filepre, Filedev, and Filellm for each file with at least one uChangedev or uChangellm. We use the online API documentation as well as the source code of the source and target libraries to understand the correct API usage in the context of the file. For 1(c), we manually find that L4 is a correct alternative to D2, noting that the LLM omitted the default value (True) of the allow_redirects argument, unlike the developer. L8 is a correct replacement of D3, where the LLM performed a refactoring. However, we find that the LLM used the function get instead of post at line 33, making L0 an incorrect alternative of D4.
After trying to match all changes in Changedev, we may find that some Changellm remain unmatched. Our aim is to categorize these as refactoring (not altering program semantics) or non-refactoring (potentially affecting code behavior). If a Changellm is refactoring, we remove it from uChangellm, as it does not impact migration correctness. In the example, the LLM introduced handling of a ValueError (L3), which changes the program behavior, so we keep it in uChangellm. Similarly, swapping variables in L5 is also non-refactoring.
Note that while PyMigBench records all migration-related code changes, it does not record additional changes indirectly related to the migration, i.e., lines that do not have a target library API usage. For example, adding async to function definitions (L1, L6, and L9) due to async calls introduced by migration are not recorded. We remove these changes from uChangellm and mark them as correct. Overall, L3 and L5 stay in uChangellm, while L1, L6, L7, and L9 get removed.
We perform a multi-reviewer manual review to ensure the quality of the matching process. The thesis author first manually reviews 190 code changes from one model. Two collaborators then review about half of those code changes each, not knowing the thesis author’s decisions. Therefore, each code change is reviewed by two reviewers independently. We then measure the agreement using Cohen’s Kappa cohen1960coefficient to quantify the level of agreement between the reviewers. We do not reach substantial agreement, so we discuss and resolve disagreements and update our coding guideline. The two pairs of reviewers then independently review another 111 code changes, achieving a Cohen’s Kappa score of 0.83 (almost perfect) and 0.73 (substantial)  landis1977measurement . Accordingly, we proceed with only the thesis author reviewing the remaining changes.
3.2.1.2 Determine migration status
Based on the matching of Changedev and Changellm, we determine each code change status as follows (example changes refer to 1(c)).
  • Correct change: LLM’s change is correct, either exactly like the developer (e.g., D1L2), or with an alternative API (e.g., D2L4), or with same APIs but with some refactoring (e.g., D3L8).
  • Incorrect change: The LLM incorrectly implemented the migration change: Used an incorrect API (e.g., get() instead of post() in L0), did not attempt to migrate it at all, or incorrectly removed part of a code.
Based on the individual matched code changes, we now automatically determine an overall migration status:
  • Response failure: The LLM did not generate a Filellm for at least one Filepre (e.g., due to token limit or API timeout).
  • Syntax error: The LLM generated a Filellm for all Filepre, but at least one Filellm has syntax errors.
  • Incorrect: The LLM could not correctly migrate any of the changes marked in PyMigBench. We assign this status when all Changedev across the migration remain unmatched.
  • Partially correct: The LLM correctly migrated only some of Changedev. We assign this status when there are only some unmatched Changedev.
  • Correct with non-refactoring changes: The LLM correctly migrated all Changedev but also performed some non-refactoring changes. We assign this status when there are no unmatched Changedev in any file, but there are some remaining unmatched Changellm.
  • Correct: The LLM correctly migrated all Changedev for this migration, without any non-refactoring changes. We assign this status when there are no unmatched Changedev or Changellm left in any of the files after the matching process.
3.2.2 RQ-A1 Findings: migration level correctness
Table 2 shows the migration level results. We find that 4o performs the best with 94% of the migrations having at least one correct code change, closely followed by Mini with 93%. When considering fully correct migrations, 4o outperforms Mini considerably, with 57% compared to 49%. LLama has the lowest performance, with only 26% being fully correct and 51% having at least one correct code change. This is mainly due to a higher proportion of response failures where LLama ran into token limits. LLama also produces more syntax errors compared to the other two models (5.1% vs. 1.9% and 1.0%).
RQ-A1, Migration Level: All three LLMs were able to perform correct migrations with 4o performing best: 94% of its migrations were partially correct and 57% were fully correct.
3.2.3 RQ-A1 Findings: code change level correctness
Table 2: Correctness of PyMigBench migrations (RQ1, migration level)
Number (percentage) of migrations
Status LLama 3.1 GPT-4o mini GPT-4o
At least partially correct
Correct 83 (26%) 154 (49%) 179 (57%)
Correct w/ non-refactorings 28 (8.9%) 53 (17%) 55 (18%)
Partially correct 48 (15%) 84 (27%) 62 (20%)
Subtotal 159 (51%) 291 (93%) 296 (94%)
Fully incorrect
Incorrect 7 (2.2%) 6 (1.9%) 2 (0.6%)
Syntax error 16 (5.1%) 6 (1.9%) 3 (1.0%)
Response failure 132 (42%) 11 (3.5%) 13 (4.1%)
Subtotal 155 (49%) 23 (7.3%) 18 (5.7%)
Total migrations 314 (100%) 314 (100%) 314 (100%)
We now present the correctness results at the individual code change level. The 314 migrations we use for our evaluation contain a total of 2,910 code changes. However, when there is a response failure or syntax error in the LLM’s migration, we cannot evaluate the correctness of the corresponding code changes in that file. Accordingly, in this analysis level, we exclude 1,999 code changes from LLama, 944 from Mini, and 914 from 4o. This leaves us with 911, 1,966, and 1,996 code changes to evaluate for the three LLMs, respectively.
3.2.3.1 Overall code change correctness
Table 3: Correctness of PyMigBench migration-related code changes (RQ1, overall code change level)
Number (percentage) of code changes
Status LLama 3.1 GPT-4o mini GPT-4o
Correct 808 (89%) 1,744 (89%) 1,867 (94%)
Incorrect 103 (11%) 222 (11%) 129 (6.5%)
Evaluated code changes 911 (100%) 1,966 (100%) 1,996 (100%)
Excluded code changes 1,999 944 914
Total code changes 2,910 2,910 2,910
Table 3 shows the overall correctness of the LLMs’ migrations at the code change level. 4o performs best, with 94% of the code changes being correct. Interestingly, Mini and LLama both perform equally well at the code change level, with both having 89% correct code changes. This is in contrast to the migration level, where Mini performs better than LLama (49% vs. 26%). This suggests that while Mini attempted more code changes compared to LLama, the ones that LLama was able to attempt without failures or syntax errors were comparatively correct.
3.2.3.2 Code change correctness by category
Table 4: Correctness of code changes across PyMigTax categories (RQ-A1, code change level by category)
Migs = Migrations, LPs = Library Pairs, CCs = Code Changes, #Eval = Number of code changes evaluated, %Cor = percentage of correct code changes.
PyMigTax category Frequency in PyMigBench LLama 3.1 GPT-4o mini GPT-4o Migs LPs $C$s #Eval %Cor #Eval %Cor #Eval %Cor Program elements Function call 229 (71%) 105 (78%) 1,804 (60%) 518 88% 1,093 87% 1,092 93% Import 318 (99%) 132 (99%) 732 (24%) 276 95% 580 94% 593 98% Decorator 39 (12%) 13 (10%) 411 (14%) 197 87% 335 90% 353 87% Attribute 54 (17%) 33 (25%) 183 (6.1%) 63 83% 103 78% 103 89% Type 17 (5.3%) 13 (10%) 55 (1.8%) 12 75% 34 85% 34 88% Exception 16 (5.0%) 13 (10%) 29 (1.0%) 17 65% 29 86% 29 93% Function ref 4 (1.2%) 4 (3.0%) 12 (0.4%) 3 33% 12 58% 12 100% Properties NoProps 57 (18%) 36 (27%) 262 (8.8%) 43 84% 150 88% 150 97% ElemNC 188 (59%) 88 (66%) 1,025 (34%) 398 84% 773 85% 771 90% ArgAdd 93 (29%) 49 (37%) 578 (19%) 150 85% 244 80% 244 88% ArgDel 66 (21%) 37 (28%) 349 (12%) 175 85% 234 87% 234 89% ArgTrans 81 (25%) 38 (28%) 317 (11%) 115 77% 201 76% 220 77% ParamAdd 13 (4.0%) 1 (0.7%) 127 (4.2%) 110 87% 127 95% 127 90% ArgNC 17 (5.3%) 11 (8.2%) 112 (3.7%) 20 60% 58 59% 58 84% AsyncTrans 13 (4.0%) 4 (3.0%) 50 (1.7%) 35 83% 50 90% 50 92% OutTrans 16 (5.0%) 15 (11%) 49 (1.6%) 25 88% 49 78% 49 92% Cardinality One-to-One 200 (62%) 98 (73%) 1,552 (52%) 408 86% 877 84% 894 91% One-to-Zero 50 (16%) 25 (19%) 297 (10%) 99 99% 290 99% 290 100% Zero-to-One 35 (11%) 18 (13%) 198 (6.6%) 34 97% 55 87% 55 89% Higher Card. 91 (28%) 45 (34%) 210 (7.0%) 94 70% 164 79% 164 84% One-to-Many 52 (16%) 23 (17%) 142 (4.8%) 61 69% 103 78% 103 80% Many-to-One 30 (9.3%) 24 (18%) 41 (1.4%) 16 62% 34 82% 34 91% Many-to-Many 21 (6.5%) 11 (8.2%) 27 (0.9%) 17 82% 27 78% 27 89% Total 321 134 2,989 911 89% 1,966 89% 1,996 94%
To understand if some types of code changes are more difficult to migrate, Table 4 shows the distribution of code change correctness across the different PyMigTax categories. The three columns in group “Frequency in PyMigBench” shows how often each PyMigTax category appears in all migrations, library pairs and code changes in PyMigBench, which provides perspective on the category’s impact in practice. This information is the same for each LLM since it describes the existing data in PyMigBench. For example, 1,804 (60%) of all 2,989 code changes involve function calls. The next three column groups show the performance of each model. Since each LLM has a different number of evaluated code changes, we show the number of evaluated code changes from each category (#Eval) alongside the proportion of these code changes that are correct (%Cor). For example, the Function call row under the Program elements group shows that LLama’s syntactically correct migrations included 518 code changes that involve function calls, and that 88% of these code changes were correctly migrated.
Program elements: Function calls are the most common program elements in terms of code changes (60%), and second most common in terms of migrations (71%) and library pairs (78%). All three LLMs correctly migrate a high proportion of these changes, with 88%, 87%, and 93% of the function calls being correctly migrated by the three LLMs. Almost all migrations and library pairs in PyMigBench require migrating import statements, a task that all three LLMs perform well, especially 4o correctly migrates nearly all of them (98%).
Mini slightly outperforms the other models in migrating decorators (90% vs. 87% and 87%); however, it does worst in attributes (78% vs. 83% and 89%). 4o performs better than other models in migrating types, exceptions and function references, followed by Mini. LLama and Mini both struggle with migrating function references (33% and 58% correct), while 4o correctly performs all such code changes. While this last category shows most discrepancy between the models, this program element is present in only 0.4% of the code changes, therefore does not affect the overall performance significantly.
Properties: The property no properties indicates that the source and target APIs are identical, and hence no changes are required. While 4o correctly leaves almost all of these APIs unchanged (97%), LLama and Mini incorrectly changed several of them, resulting in 84% and 88% correct, respectively.
Element name change is the most common property in PyMigBench. This indicates that the target APIs commonly bear different names from the source APIs. All LLMs perform well in making this change, with LLama, Mini, and 4o correctly migrating 84%, 85%, and 90% changes, respectively.
All three LLMs performed better or equally well in argument deletion compared to those with argument addition, which is expected as adding an argument requires the LLM to find a suitable replacement, while deleting an argument is a simpler task. The Argument transformation property indicates that an argument requires various changes, such as changing the type or value. The three LLMs correctly migrate only 76-77% of the argument transformation changes. We discuss specific failure instances in Section 3.4.
The property parameter addition to decorated function is found in migrations in just one library, click library, where some decorators require adding a parameter to the decorated function libclick . All three LLMs perform well in migrating these changes; interestingly, Mini performs better than 4o (95% vs. 90%). Mini also performed better than 4o in migrating decorators, which are common in the click library. These are the only two categories where 4o does not perform the best.
Argument name change applies when an argument representing the same semantics has different names in the source and target APIs. While 4o performs reasonably well in migrating these changes (84%), LLama and Mini fail frequently with them (only 60% and 59% correct).
Cardinality: One-to-One code changes are the most common in PyMigBench, and all three LLMs perform well in migrating them. One-to-Zero code changes are the ones where a source API needs to be removed, but no target needs to be added. Because it is a simple delete operation, all three LLMs migrate all or almost all of these changes correctly. Its counterpart, Zero-to-One code changes, are the ones where a target needs to be added, but no source needs to be removed. Interestingly, LLama performs the best in migrating these changes (97% correct), making this the only category where it performs better than the other models (87% and 89% correct). While previous research showed that higher-cardinality code changes are generally difficult compared to one-to-one code changes alrubaye2018automating ; wang2016transforming ; zhang2020deep ; HuangMappingAPI2024 , the PyMigBench dataset shows that higher cardinality changes are frequent with 34% of the library pairs requiring at least one higher-cardinality code change. Overall, we find that the LLMs are reasonably successful at migrating them, with 70%, 79%, and 84% correctness rates.
RQ-A1: Code Change Level LLMs correctly migrate a high proportion of most code changes, with 4o achieving 94% correct code changes. The models perform reasonably well in migrating difficult higher-cardinality code changes (4o gets 84% correct), but struggle relatively more with argument transformation and argument name change.
3.3 RQ-A2 How many migrations pass unit tests?
We now assess the same LLM PyMigBench migrations from RQ1 but using the unit tests available in the corresponding repositories. A test-based evaluation that runs the code ensures that the migration achieves the expected behavior. Note that we should not expect that an LLM migration improves the rate of passing tests; a successful migration means that each test that previously passed on codedev must still pass on codellm. We now explain our evaluation set up.
3.3.1 Preparing the code
We make a copy of the repository after the developers’ migration (codedev). We then make another copy of developer’s migration, but replace all Filedev with Filellm; we refer to this as codellm. The idea is that codellm is basically the version of codedev where the LLM did the migration on behalf of the developer.
3.3.2 Setting up the virtual environment
To set up a virtual environment, we need to identify the Python version used in the client repository. If this information is available in the setup.py file, we use that version. Otherwise, we resolve the version based on the migration commit date as follows. We first identify the release dates of all minor versions of Python (3.6, 3.7 etc). Then, we find the latest release date that is before the migration commit date. This is the latest Python version that was available at the time of the migration; we use this version to create the virtual environment. Next, we install the code dependencies using pyproject.toml, setup.py and requirements files. In cases where specific dependency versions are not specified in those files, we look up the version history of the dependency on PyPI and install the latest version available at the migration commit date, similar to how we resolve the Python version. This ensures that the dependencies are compatible with the code at the time of migration.
3.3.3 Running the tests and coverage
Once the virtual environment is ready, we run all the unit tests while measuring coverage in the repository on codedev. If the run has errors, we read the error log and try to fix the errors, and run the tests again. We maintain a configuration file where we record any project-specific configurations or commands we used. We include this information in our artifact.
We find that 218 out of the 314 migrations have at least one test. Among these, 172 migrations had unresolvable errors. The failures are primarily due to missing dependencies, specially for older projects. The remaining 46 migrations have tests that we are able to successfully run on codedev.
However, for the tests to be useful to validate migration, they must cover the migration-related code changes. Using the coverage report, we find that only 27 migrations have at least one test that covers the code changes, and among them, 25 migrations have at least one test that passes on codedev. We run the entire test suite on codellm for these 25 migrations for each model using the same randomly selected LLM run used in RQ-A1.
3.3.4 Determining migration status
We compare the test results between codedev and codellm and assign the following statuses to the migrations:
  • Correct: All passing tests in codedev also pass in codellm.
  • Partially correct: Some of the tests that pass in codedev pass in codellm, but the others fail or raise run-time errors.
  • Incorrect: None of the passing tests in codedev pass in codellm; they all fail or raise run-time errors.
3.3.5 RQ-A2 Findings
Table 5: Correctness of 25 PyMigBench migrations using their available unit tests (RQ2)
Number (percentage) of migrations
Status LLama 3.1 GPT-4o mini GPT-4o
Correct 9 (36%) 13 (52%) 16 (64%)
Partially correct 1 (4.0%) 5 (20%) 2 (8.0%)
Incorrect 15 (60%) 7 (28%) 7 (28%)
Evaluated 25 25 25
Table 5 shows the results for the 25 migrations. 4o has the highest percentage of correct migrations, with 64%, followed by Mini with 52%, and LLama with 36%. Mini has more partially correct migrations than 4o and LLama, leading it to have an equal number of incorrect migrations as 4o (28%). LLama, on the other hand, has a much higher proportion of incorrect migrations (60%).
RQ-A2 Out of 25 PyMigBench migrations with unit tests covering the migration-related code changes, the LLMs correctly migrated 36%-64%, with 4o being the highest.
3.4 Implications and discussion
3.4.1 Are LLMs suitable for library migration?
We look into some specific examples to understand the promises and limitations of LLMs for library migration.
3.4.1.1 The good
We start with the promising side of using LLMs for library migration. In addition to the high overall correctness of individual code changes and migrations in RQ1, our evaluation reveals that the evaluated LLMs can migrate code changes that were traditionally known as difficult to migrate xu2019meditor . We discuss some examples.
Different API styles: The library argparse libargparse provides API functions for parsing command line arguments, while click libclick provides decorators. Figure 2 shows a migration example between these libraries. Despite the completely different API styles, the three LLMs correctly migrate 94% of code changes between this library pair.
Figure 2: 4o correctly migrating from argparse to click, libraries using different API styles.
Higher cardinality migrations As shown in RQ2, LLama, Mini, and 4o successfully migrated 70%, 79%, and 84% of the higher cardinality code changes, respectively, which the literature always viewed as complex migrations xu2019meditor ; alrubaye2020learning . Figure 3 shows an example of a many-to-many migration from twitter to tweepy done by LLama. In addition to correctly replacing the functions Twitter(), and OAuth() with OAuthHandler(), set_access_token(), and API() with correct arguments, LLama also splits the code into multiple statements, making the code more readable.
twitter to tweepy, cloudbotirc/cloudbot/f8243223/twitter.py
Figure 3: LLama correctly migrating a many-to-many code change
Inferring changes outside of the source and target APIs The migration in Figure 4 adds the async_mode argument to the SocketIO() function. This is interesting because the API is from the Flask-SocketIO library, which is not the source (eventlet) or target (gevent) libraries. However, the SocketIO function works with several libraries, eventlet being the default flasksocketio . This means, if 4o did not explicitly set the async_mode argument to gevent, the code would break after the migration.
Figure 4: 4o performing a correct inferred change
3.4.1.2 The Bad
We also analyze cases where the LLMs did not correctly migrate the code and present notable observations.
Failing to handle argument transformation Figure 5 shows an example where the source library retrying libretrying expects the time arguments in milliseconds, while the target library tenacity libtenacity expects them in seconds. The dataset includes 28 instances of this transformation, and all three LLMs failed to handle it in most cases, except for 4o, which handled three instances correctly, while the other two models correctly handled one instance each. This suggests that while the LLMs can map parameters, differences in expected format or units may not be properly managed.
Figure 5: Mini filing in argument transformation.
Replacing exception type We find cases where the LLM replaces generic exception types with specific ones, which can potentially fail to catch the same exceptions as the original code. This occurs in both migration-related and non-migration-related changes. For example, the LLM replaces an original Exception with ValueError. While using more specific exceptions is considered good practice cacho2014does , it does change the behavior of the code. Since modern development tools are often equipped to suggest better exception handling, it may be safer for an LLM-based migration tool to avoid changing existing exception types during a migration.
3.4.1.3 The ugly
We also observe the following problematic LLM behavior.
Extra changes LLMs are known for not always following the instructions provided in the prompt ye2023cognitive . Despite our prompt explicitly asking the LLMs to avoid any refactoring or non-migration changes, we find cases where an LLM, particularly LLama, completely removes parts of the code that do not include any source APIs. We speculate that this behavior may stem from the LLM attempting to optimize output size due to token limitations. Silently deleting code can be dangerous for the overall functionality of the system, and led to test failures in RQ2 for many cases.
Similarly, the LLMs sometimes refactor code that is not related to the migration, adopting a different style. While these improvements can be beneficial, they are unrelated to the migration task at hand. Future work could explore the use of an interactive IDE plugin that confirms the code changes before applying them, potentially allowing the developer to directly edit the changes.
While less “dangerous”, we also find that the LLMs remove code comments and docstrings from the source code, even though the comments remain relevant after migration.
Large file handling We observe that large files pose a challenge to LLama, given its smaller context size. Additionally, we observe that even the other two models often migrate an initial part of a large file but then incorrectly state that the remaining part does not require any changes.
3.4.2 Differences in evaluation setup
While the scope of RQ-A2 is limited due to the availability of the tests, we do find cases where the migrations are correct in RQ-A1 but incorrect in RQ-A2, or vice versa. To better understand these discrepancies, we manually inspect migrations with different results and provide some examples.
3.4.2.1 Library version compatibility
When evaluating RQ-A1, we mark a code change to be correct if it is correct as of the latest version of the library. However, in RQ-A2, we install the version the developers specified, and if it is not available, we resolve the library based on the migration commit date (see Section 3.3). Since we do not ask for a specific version number in the prompt, we notice that the LLMs usually pick APIs from the latest version of the target library. In some cases, this version discrepancy led to runtime errors when running tests in LLM’s code, which we report as incorrect in RQ-A2. Note that we consider version numbers moving forward when designing PyMigTool in Section 4.
3.4.2.2 Function parameter shadowing import name
ipaddr to ipaddress, google/capirca/eb768ea7/capirca/lib/nacaddr.py
Figure 6: Example of function parameter shadowing import name.
Notice in Figure 6 that the function parameter in the function IP on line 6 (Filepre) is named ipaddress, the same as the target library. The developer changed the parameter name to ip in Filedev. However, while Mini correctly replaces function IPNetwork with ip_network, it retained the original parameter name, causing it to shadow the new import name and leading to test failures. 4o changed the parameter name to ipaddress_str to avoid the shadowing issue.
3.4.2.3 Developer doing non-refactoring changes
Recall that in RQ-A1, we find cases where the developer makes non-refactoring changes in the same commit as the migration while the LLM only does the migration it is expected to do. We count this as a correct migration in RQ-A1. However, given the changed behavior (which is often reflected in the tests), the tests fail in RQ-A2.
Overall, the discrepancies show that both evaluation techniques provide different perspectives on the migration correctness.
3.4.3 Understanding the cost of LLM-based migration
The 314 migrations we use in our experiments involve a total of 258,198 lines of code, or 258.0 KLOC. The 10 runs per model makes it 2582.0 KLOC per model. The total cost of running our experiments was approximately \$15 for GPT-4o mini and \$420 for GPT-4o, which equates to \$0.01 and \$0.16 per KLOC, respectively. OpenAI has since reduced prices, so current costs would be lower. LLama 3.1 is free and open-source, though we used a pro subscription of Hugging Face (US\$9/month [^3] https://huggingface.co/pricing) to avoid any installation issues. While library migration is an infrequent task in a project’s lifetime, it requires significant developer effort when it happens. The cost of a developer’s time can easily exceed the cost of using an LLM, even GPT-4o, which can justify the cost of LLM-based tooling.
3.5 Threats to validity
Moha: I tried merging the two threats sections. But it because repetations of ”in phase one” and ”in phase two”. So I kept them separate for now.
3.5.1 Internal validity
Our evaluation involves two manual steps: validating correctness of code changes and labeling a non-migration related change as refactoring or non-refactoring. Manual activities are inherently subjective and may contain errors. We minimize this by having two authors independently perform the manual activities and resolving disagreements through discussion.
3.5.2 Construct validity
In RQ-A1, we analyze the code change categories based on Changedev, not on Changellm. The categories can differ if the LLM produces a different code change than the developer. Categorizing the LLM changes would require additional manual effort following the our previous work in PyMigTax pymigtax , which is beyond the scope of this work. That said, our observation is that the code/API differences in most cases are minor, and would not yield in a different PyMigTax category.
Tests passing may not always guarantee migration correctness (e.g., the tests do not cover certain edge cases). We mitigate this, to the extent possible, by selecting repositories with high test coverage in LABEL:rq:unseen and ensuring that the tests cover the source API usages.
3.5.3 External validity
We use three off the shelf LLMs. While using an LLM specifically tuned for the library migration task could potentially yield better results, the goal of this initial empirical study is not to develop a migration tool, but to understand the challenges and opportunities of using LLMs for library migration. Our results may be viewed as a lower bound of the performance of potentially more specialized models.
The libraries included in PyMigBench are all old-enough libraries where the LLMs are likely familiar with their APIs and API usage. This familiarity with APIs and API usage is precisely our hope when using LLMs for migration. However, being familiar with the APIs alone does not always guarantee being able to map them to each other or apply the right transformations on all code bases, which is why our evaluation is necessary. For newer libraries, additional information, e.g., API documentation, may be needed to guide the LLM in the migration process. This can be an interesting avenue for future work.
4 Phase two: PyMigTool: a tool for end-to-end Python library migration
In Section 3, we showed that LLMs can perform library migrations. However, we see two major challenges that make solely relying on LLMs in a practical tool ineffective: (1) We provide the LLM with only code files that directly use the source library. However, for a given project, there may be other files that depend on the migrated code and also require changes during the migration. (2) All LLMs often return only part of the code, skipping parts of the code that do not require migration. In this phase, we build and evaluate PyMigTool, a CLI tool that automates the full library migration process for a full project by combining the power of LLMs with pre and post processing to address the above challenges.
PyMigTool first prepares the project that needs migration by setting up a virtual environment. It then runs an LLM to migrate the code, and finally performs post processing to improve the migration results. We run an extensive evaluation of PyMigTool on a set of 717 new, simulated migrations. Overall, PyMigTool is able to successfully migrate 32% of the migrations without any manual intervention. For a sample of the remaining not-fully-successful migrations, we find that a median of only 14% of the code lines in the project require manual changes to complete the migration, while the rest of the code is correctly migrated by PyMigTool.
4.1 Building PyMigTool
PyMigTool takes as input the following: (1) the path to a python project where the migration should happen, (2) the source library name, (3) the target library name and version, (4) the python version in which the code should run, (5) paths to one or more requirements files, (6) and an LLM to use for the migration. A key criteria for the input project is that it contains tests with high coverage, as PyMigTool uses the tests to identify the source library usage as well as to verify the migration. While the migration will run regardless of the test coverage, migration results without test validation may not be as reliable.
PyMigTool works in three main steps: Preparation round, LLM migration round, and Post-processing round. The preparation round sets up a virtual environment using the requirements files, and runs the tests to get the baseline results. The LLM migration round migrates the project to use the target library instead of the source library using the LLM the user specified. Finally, if the LLM migration is not fully successful, PyMigTool performs some non-LLM post processing steps to improve the migration results.
4.1.1 Preparation round
In the preparation round, PyMigTool first creates a virtual environment for the provided python version using the command python -m venv .venv, and then installs the required libraries using the command pip install -r <requirements-file-path> for each of the requirements file. It then installs the target library at the provided version using the pip install command. PyMigTool requires some additional packages, such as pytest, pytest-cov, which it also installs during this step.
Once the environment is set up, PyMigTool runs the tests using the command python -m pytest on the project and saves the test report. We refer to this version of the test report as TestReportpremig, which is used as the baseline for the migration.
4.1.2 LLM migration round
In our preliminary exploration of LLMs (Section 3), we use PyMigBench where the files that require migration were already known. However, in a real-world scenario for a full project, we do not know which files require migration. Therefore, PyMigTool first identifies the files that require migration, and then uses the LLM to migrate these files.
4.1.2.1 Identifying files that require migration
PyMigTool migrates one whole file at a time, therefore, it is sufficient to identify which files use the source library, without necessarily identifying the exact source library usage within these files. PyMigTool combines two techniques to identify the files that require migration: it uses static analysis to identify import statements and uses dynamic analysis to generate a run-time-generated call graph of the project using profiling data.
geolib (library code)
1class Square:
2 def __init__(self, l):
3 self.l = l
4 self.name = "Square"
5
6 def area(self):
7 return self.l ** 2
serialize.py (uses library w/o import)
1def to_dict(shape):
2 return {
3 "name": shape.name,
4 "area": shape.area()
5 }
main.py (uses library with import)
1from serialize import to_dict
2from geo import Square
3import pprint
4
5sq = Square(5)
6pprint(to_dict(sq))
serialize_test.py (test code)
1from serialize import to_dict
2from geo import Square
3
4def test_to_dict():
5 square = Square(4)
6 result = to_dict(square)
7 assert result == {"name": "Square", "area": 16}
Figure 7: Example of using a library with/without importing
Detecting import statements: If a file in the project imports a library, PyMigTool considers that the file uses the library, and thus require migration. In Python, a library may be imported using a name different than the library name. For example, the library PyYAML can be installed using the command pip install pyyaml, but it has to be imported as import yaml. The former, pyyaml, is the library name, and the latter, yaml, is the import name. Furthermore, a library can have more than one import names. PyMigTool only takes the source library name as input, therefore, it needs to find the import names of the source library in the project. To do this, PyMigTool uses the Johnnydep tool libjohnnydep , which extracts the import names from the wheel file of the library.
Next, PyMigTool uses the ast module to parse each Python file in the project and identify the import statements. It then compares the imported modules with the import names of the source library to identify the source library usage in the project.
In the example in Figure 7, the file main.py directly imports the library geolib, thus PyMigTool will identify this file as requiring migration. However, the file serialize.py does not import the library, therefore, will not be identified as requiring migration by the static analysis.
Detecting source library calls: In Figure 7, the file serialize.py does not import the library geolib, but both the main.py and serialize_test.py pass the geolib.Square object to the serialize.to_dict() function. That means when the code runs, the serialize.py file will use the geolib library, even though it does not import it. This shows that, identifying the import statements alone is not sufficient to identify all of the source library usage in the project. Accordingly, PyMigTool uses a run-time call graph to identify further source library usage in the project.
PyMigTool runs the tests in the project with profiling using the pprofile package libpprofile . The profiling data includes all the calls made during the test execution, and generates a report in the CallGrind format doccallgrind . A call information includes the caller function and the callee function, as well as the line number where the call happened. A function is identified by the path of the file in which the function is defined, and the qualified name which is unique within the file. PyMigTool uses this data to build a call graph of the project.
To know if a file requires migration, PyMigTool needs to determine whether any part of the file directly calls a source library function. A function can be defined in one of the three places: in python built-in modules, in the project, or in a third-party library. A Python system module file path is enclosed within angle brackets in the profile report; for example, <frozen os> indicates that the function is defined in the built-in os module. If a function is defined in a file under the project path (which PyMigTool takes as input), it considers this function to be in the project. PyMigTool also knows the path of the virtual environment where the third-party libraries are installed. Python virtual environments have a specific directory structure that PyMigTool leverages. For example, in Windows systems, the import name yaml would be in the <venv_path>/Lib/site-packages/yaml directory. Therefore, if a call is to a function that is defined in a file in this directory, PyMigTool can conclude that the function belongs to the yaml import name, i.e., the pyyaml library. In the call graph, we store the owner of each function, which can be project, system, or the import name of the library.
Given this constructed call graph, PyMigTool iterates over each call. If the caller is owned by the project and the callee is owned by the source library, then the file that contains the caller function requires migration and PyMigTool adds it to the list of files that will be passed to the LLM.
Overall, in Figure 7, PyMigTool will identify the file main.py requiring migration using static analysis, and serialize.py requiring migration using dynamic analysis.
4.1.2.2 Migrating using an LLM
PyMigTool can be configured to use any LLM that adheres to OpenAI’s APIs openaimodels , which has become a standard API for most models. For each file that requires migration, PyMigTool generates a prompt using the template in 2 and sends it to the LLM. The prompt is based on the prompt used in Section 4 (1), but enhanced with more specific instructions, and also includes the versions of the source and target libraries.
Listing 2: Prompt template used by PyMigTool
1The following Python code currently uses the library <source-lib> version<source-ver>.
2Migrate this code to use the library <target-lib> version <target-ver> instead.
3
4**Instructions:**
51. **Explain the Changes**: Begin the output with a brief explanation of the specific changes you made to migrate from "<source-lib>" to "<target-lib>".
62. **Provide the Modified Code**: After the explanation, present the modified code. Provide the entire code after migration even if only a part of it is changed.
7
8**Important Guidelines**:
9- Only make changes directly related to migrating between "<source-lib>" and "<target-lib>".
10- Do not refactor, reformat, optimize, or alter the original coding style.
11- The code given to you is part of a larger application. Do not change the names of classes, functions, or variables, because it can break the application.
12
13Original code:
14<premig-code>
The LLM response contains the migrated code, which PyMigTool extracts and saves. When the LLM finishes migrating all files, PyMigTool replaces the original files in the project with the migrated files. It then runs the tests and generates a test report, which we refer to as TestReportllmmig.
4.1.3 Post-processing round
PyMigTool performs two post-processing steps we describe below.
4.1.3.1 Merge skipped code
During the the preliminary experiments (Section 3), we found that it is common for the LLMs to skip some parts of the code in the response, especially if those parts do not require migration, despite being instructed to return the full migrated code (see paragraph 3.4.1.3). In this post-processing step, PyMigTool identifies such skipped code, and merges it with the migrated code.
For each file that was migrated in the LLM migration round, PyMigTool generates a diff between the original file and the migrated file using python’s difflib module. Then it attempts to detect which of the chunks were entirely removed from the original file. We make the following assumption to identify skipped code: 1. A skipped code must be mostly removal of lines. Because if the hunk has similar amount of additions and removals, it is likely that the hunk is replacing some code with other code, likely performing migration. 2. The skipped code should not use source library APIs, because this makes the hunk migration-related otherwise. 3. A skipped code should not be too small. This assumption avoids incorrectly considering small migration-related removals as skipped code.
We find the below specific baseline values for the above assumptions by trial-and-error to find a good balance between incorrectly merging migration related removals and incorrectly not merging skipped code:
  • Small (0-9 lines of removals): this hunk is not considered skipped irrespective of the number of additions.
  • Medium (10-19 lines of removals): if this hunk has at least one addition of line or use any source library APIs, it is not considered skipped.
  • Large (20 or more lines of removals): if this hunk has less than 90% of removed lines compared to total added+removed lines, it is not considered skipped.
The hunks contain the line numbers in the previous and new version of the file. Using the line numbers, PyMigTool inserts the removed lines in the migrated file, starting with the last removed hunk and going to the first removed hunk. The reverse order is important because inserting the lines would otherwise change the line numbers of the hunks that are inserted later. After inserting the skipped code, PyMigTool runs the tests again and generates a new test report, which we refer to as TestReportmerge.
4.1.3.2 Infer async transformation
In Python, a call to an asynchronous function must be awaited, i.e., preceded by the await keyword. Again, if a function call is awaited, the caller function must also be made asynchronous, i.e., it must be preceded by the async keyword.
When there is a migration from a synchronous library to an asynchronous library, some functions in the migrated code may become asynchronous, which requires their transitive calls to be awaited, and transitive callers to be made asynchronous as well. In this step, PyMigTool identifies the functions that were made asynchronous by the LLM (asynced functions), then find all other functions that needs to be made asynchronous (to-async functions) and all calls that need to be made await (to-await calls), and applies the changes to the project. Additionally, PyMigTool also decorates the corresponding test functions with the pytest.mark.asyncio decorator, and installs the pytest-asyncio package, which are required to correctly run asynchronous tests. Below we describe the steps in detail using the example in Figure 8.
service.py
1import requests
2
3
4def fetch_data(year):
5 url = f"https://api.io/data/{year}"
6 res = requests.get(url)
7 res.raise_for_status()
8 data = req.json()
9 return data
service_test.py
1from service import fetch_data
2
3def test_fetch_data():
4 year = 2023
5 data = fetch_data(year)
6 assert isinstance(data, dict)
7 assert "days" in data
8 assert isinstance(data["days"], list)
9 assert len(data["days"]) == 365
$\Downarrow$ LLM migration $\Downarrow$
1from aiohttp import ClientSession
2
3async def fetch_data(year):
4 url = f"https://api.io/data/{year}"
5
6 async with ClientSession() as session:
7 async with session.get(url) as res:
8 res.raise_for_status()
9 data = HLCawait res.json()
10 return data
1from service import fetch_data
2
3
4def test_fetch_data():
5 year = 2023
6 data = fetch_data(year)
7 assert isinstance(data, dict)
8 assert "days" in data
9 assert isinstance(data["days"], list)
10 assert len(data["days"]) == 365
$\Downarrow$ Infer async transformation $\Downarrow$
1from aiohttp import ClientSession
2
3
4async def fetch_data(year):
5 url = f"https://api.io/data/{year}"
6
7 async with ClientSession() as session:
8 async with session.get(url) as res:
9 res.raise_for_status()
10 data = await res.json()
11 return data
1import pytest
2from service import fetch_data
3
4@pytest.mark.asyncio
5async def test_fetch_data():
6 year = 2023
7 data = await fetch_data(year)
8 assert isinstance(data, dict)
9 assert "days" in data
10 assert isinstance(data["days"], list)
11 assert len(data["days"]) == 365
Figure 8: Example of applying LLM migration and then infer async transformation
The first row presents the original code before migration. The second row displays the code after LLM migration. The third row shows the after applying infer async transformation. The left column contains the main code, while the right column contains its associated test. Changes introduced by each migration step are highlighted in red.
Identifying asynced functions For a given file that was migrated in the LLM migration round (Section 4.1.2), PyMigTool builds a abstract syntax tree (AST) of the code before migration and after migration. PyMigTool then finds the asynchronous function definitions in the code after migration. For each such function, PyMigTool finds functions with the same qualified name in the AST of the code before migration. If such a function exists in the AST before migration, but was not asynchronous, we consider the function as asynced. For example, in Figure 8, the function fetch_data is asynced, because it was synchronous before migration, but is asynchronous after migration. If a corresponding function is not found at all, it means that the function is newly added due to the migration, and we do not consider it as asynced, because it is introduced by the LLM and does not have callers in other parts of the project.
Identifying to-async functions and to-await calls If a function a calls a function b, and b calls a function c, then a is said to transitively call b and c. For a given asynced function, PyMigTool finds all its transitive callers using the call graph, and marks them as to-async. For example, in Figure 8, the function fetch_data is asynced, and it is called by the function test_fetch_data, therefore, PyMigTool marks test_fetch_data as to-async. It then again uses the call graph to find all call sites of the asynced and to-async functions, and marks them as to-await. For example, in Figure 8, the call to fetch_data in the function test_fetch_data is marked as to-await, because it is a call to an asynced function. The to-async functions need the async keyword to be added to their signature, and the to-await calls need the await keyword to be added to them.
Applying the changes PyMigTool uses LibCST liblibcst to parse the code, search code elements to be modified, and applies the following changes:
  1. 1.
    If a function definition is marked to-async, and it is not already asynchronous, then PyMigTool adds the async keyword to the function definition, for example the function definition test_fetch_data in Figure 8.
  2. 2.
    If a call is marked to-await, and it is already not awaited, PyMigTool adds the await keyword to the call, for example the call to fetch_data in test_fetch_data in Figure 8.
  3. 3.
    If a test function is asynced or is marked to-async, PyMigTool decorates the test function with pytest.mark.asyncio decorator, and installs the pytest-asyncio package to the virtual environment, which is required to run asynchronous tests. For example, in Figure 8, the function test_fetch_data is decorated with pytest.mark.asyncio.
After finishing the async transfomations, PyMigTool runs the tests again and generates a new test report, which we refer to as TestReportasync.
4.2 Evaluating PyMigTool
We evaluate PyMigTool on a set of 717 simulated migrations in 175 unique projects. By simulated migrations, we mean that this migration did not exist in the project’s history; instead, we find source libraries used in a project and migrate the project to an analogous target library. We ensure that the projects we choose have high test coverage, so that the evaluation is reliable. We run each migration just one time, unlike in Section 3, because we found in that section that the LLM does not produce highly divergent migration results when run multiple times on the code. The setup we use is same as the setup in LABEL:rq:unseen (LABEL:sec:rq-llm-unseen), where we evaluate the LLMs on a set of migrations that were not used in training the LLMs.
We aim to answer the following research questions:
  1. RQ-B1
    How many migrations can PyMigTool perform correctly? We measure how well PyMigTool performs on the simulated migrations without any manual intervention.
  2. RQ-B2
    How much manual effort is required to complete the incomplete migrations? While the best case scenario is that PyMigTool can migrate the project without any issues, a more practical expectation is that PyMigTool would do most of the task correctly, but there are some parts that require manual intervention.
4.2.1 Data collection
4.2.1.1 Selecting repositories
We want to find repositories that are well covered by unit tests, so that we can automatically evaluate the migration correctness using the tests. We start by downloading a list of 95,472 Python GitHub repository metadata using SEART seart service. We use the filters to select Python repositories that have at least one commit after January 1, 2024 to discard repositories that are not actively maintained. We also filter out repositories with less than 100 lines of code to ensure that the repositories have some code to migrate.
For each repository, SEART metadata includes many details, including the repository name, the latest commit, the lines of code in the repository and repository size in KB. Before cloning the repository, we use some heuristics to filter out the repositories that may not be suitable for evaluation.
  1. i.
    Large repositories: Extra-large repositories would take too long to clone and analyze, so we filter out 30,563 repositories larger than 10MB, leaving us with 64,909 repositories.
  2. ii.
    No requirements files: We use requirements files to build the virtual environment. We filter out 39,916 repositories that do not have a requirements.txt file in the root directory, leaving us with 24,993 repositories. We check this by making an HTTP head request to the GitHub raw file URL. If the response code is not 200, we conclude that the file does not exist.
  3. iii.
    Filtering by test files: Python test file names usually contain the term test. We use GitHub code search API [^4] GitHub code search API: https://docs.github.com/en/rest/search/search#search-code to query files that contains the term test, and filter out a repository if it does not have any such files. We filter out 12,558 repositories that do not have any test files, leaving us with 12,435 repositories.
When a repository passes the above heuristics, we clone it and attempt to run the tests. In this process, we filter out the repositories with the following criteria:
  1. iv.
    Could not run tests: We attempt to setup a virtual environment for the repository using the requirements file and run the tests following Section 3.3.2. Due to various repository specific requirements; e.g., python version, system requirements, incomplete requirements file; we are often unable to run the tests. We filter out 10,104 repositories that we could not run the tests for, leaving us with 2,314 repositories.
  2. v.
    Low passing rate: We filter out the repositories that have less than 90% of the tests passing, which filters out 957 repositories, leaving us with 1,357 repositories.
  3. vi.
    Low coverage: We filter out repositories that have less than 80% line coverage by the tests, which filters out 557 repositories, leaving us with 800 repositories.
  4. vii.
    Tests require user interaction: Some repositories require user interaction for the test to continue. For example, one repository requires a user to select an application to run a file. We identify 2 repositories that require manual intervention and exclude them, leaving us with 798 repositories, which we use for evaluation. We store the commit SHA for the version of the repository we use to ensure reproducibility.
Selecting library pairs We collect the dependencies of the 798 selected repositories by parsing their requirements files. These dependent libraries are the potential source libraries. Then for each source library, we find a list of candidate analogous libraries using GPT 4o mini, and then manually we manually verify them. We describe the process below.
Collecting source libraries For each repository, we parse the requirements file to find the dependencies. This gives us an initial list of 1,421 unique source libraries. We then look up the library metadata on libraries.io libraries.io . We exclude 12 libraries that are not found on libraries.io, leaving us with 1,409 libraries. To ensure that the libraries are commonly known, we filter out 602 libraries that have less than 100 dependents, and filter out 2 libraries that do not have a description, leaving us with 805 potential source libraries.
Listing 3: Prompt template for finding candidate analogous libraries
1Identify alternative Python libraries that offer similar functionality as <lib_name>,
2allowing for potential replacement in applications.
3The alternative libraries should be available on PyPI.
4Produce the result as a JSON list, having only the library names.
5The produced library names should match the PyPI package names.
Finding analogous library pairs For each of the 805 source libraries, we find a list of candidate analogous libraries using gpt-4o-mini-2024-07-18 with the prompt template in 3. This gives us a total of 4,856 candidate library pairs. All of these pairs may not actually be analogous, i.e., they may not be suitable for migration, so we need to verify them. This manual task can be very time consuming, therefore, we first automatically filter out some of the pairs. First, we filter out 4,215 pairs where the source library is used in less than 10 of the selected repositories. We do this to ensure that every included library pair has at least 10 repositories to evaluate on, so that we can get a good estimate of the performance. Then, we filter out 192 pairs with unmaintained potential target library. Specifically, we exclude the target libraries that have not been updated since January 1, 2024. This leaves us with 449 pairs to manually verify.
During manual evaluation, we look into the library documents and confirm 60 library pairs that are indeed analogous and can be potentially migrated. We exclude the remaining because they fall into one of the below categories:
  1. i.
    Not analogous: Some libraries are not analogous, e.g., they provide different functionality or have different APIs.
  2. ii.
    Test related libraries: For example, pytest, mock, etc. We exclude these libraries because our evaluation process relies on the tests. Therefore, migrating the test libraries themselves would may break our automated evaluation.
  3. iii.
    Non-API tools: Some dependencies are command line interfaces. Since they are not used in the code, we ignore them. For example, black, flake8, etc.
  4. iv.
    Documentation tools: Some dependencies are used for documentation generation. Since they are not used in the code, we ignore them. For example, sphinx, mkdocs, etc.
4.2.1.2 Collecting migrations
From the above process, we have a list of 798 repositories and 60 analogous library pairs. For each of these repositories, we now find which of the source libraries from the analogous library pairs are used in the repository. To find actual usage, not just listing in the requirements file, we parse the source code of the repository and look for import statements that use the source libraries. We find that 175 repositories use at least one of the source libraries. Each source library can have one or more analogous target libraries. A repository, a commit, a source library and a target library thus form a unique migration instance. We get a total of 717 unique migration instances, which is the product of 175 repositories and 48 library pairs. These numbers are less than the 798 repositories and 60 analogous library pairs we started with, because we find many instances where a repository declares a dependency on a library, but does not actually use it in the code.
To summarize, our evaluation is based on 717 migrations between 48 library pairs in 175 repositories. The 48 library pairs include 24 unique source libraries and 37 unique target libraries, and 55 unique libraries in total.
4.2.2 Running migrations
In our initial LLM experiment (Section 3), we see that GPT-4o performs better than the other two models. Therefore, we use GPT-4o (gpt-4o-2024-11-20) to run the migrations in this evaluation.
4.2.3 RQ-B1: How many migrations can PyMigTool perform correctly?
4.2.3.1 RQ-B1: Approach
We describe in Section 4.1 that PyMigTool generates TestReportpremig before the migration, and three test reports after each migration step: TestReportllmmig after the LLM migration round, TestReportmerge after the merge skipped code, and TestReportasync after the infer async transformation. For a given post migration test report, we define correctness of a migration after that step as the proportion of tests that pass in the test report compared to the tests that passed in the test report before the migration.
Table 6: Example of migration correctness calculation.
Test case TestReportpremig TestReportllmmig TestReportmerge TestReportasync
Test 1 Pass Pass Pass Pass
Test 2 Pass Fail Fail Fail
Test 3 Pass Fail Fail Pass
Test 4 Pass Fail Pass Pass
Test 5 Fail - - -
# passing 4 1 2 3
correctness 25% 50% 75%
We use Table 6 to explain our correctness calculation. Assume that there are fives tests in the project. TestReportpremig has 4 passing tests and 1 failing test. We only take into account the tests that passed in TestReportpremig, because the failing tests cannot cannot indicate whether the migration is correct or not. In TestReportllmmig, only 1 test passes out of 4, therefore, the correctness is 25%. Similarly, in TestReportmerge and TestReportasync, 2 and 3 tests pass respectively, therefore, the correctness is 50% and 75% respectively.
4.2.3.2 RQ-B1: Results
Figure 9: Distribution of correctness of migrations.
The solid line indicate the median correctness, and the dashed line indicates the mean correctness.
Table 7: Comparison between migration steps
Metric LLM Only LLM+Merge LLM+Merge+Async
% fully correct migrations 27% 30% 32%
Median correctness 65% 80% 86%
# migrations applied to 717 260 129
# Migrations improved 88 (34%) 45 (35%)
Average improvement in correctness 46% 65%
Table 7 summarizes our results, while comparing the different migration steps. The table shows multiple metrics: the percentage of fully correct migrations is the proportion of migrations where all tests that passed before migration all pass after this migration step. Median correctness is the median correctness value across all evaluated migrations, where correctness is calculated according to Table 6. Number of migrations applied to is the number of migrations where this step was applicable to and got applied. For example, not all migrations require merging unchanged code or applying async transformations. Number of migrations improved shows the number of migrations where correctness improved in comparison to the previous step. Average improvement in correctness is calculated for those migrations where correctness improved. Figure 9 also shows the distribution of correctness of migrations after the each migration step.
All three violins have concentrated values at 0% and 100%, indicating that most of the migrations either pass all tests or fail all tests. From this figure and Table 7, we see that the median correctness of migration increased from 65% in the LLM migration round to 80% after the merge skipped code, and further increased to 86% after the infer async transformation. The distributions show that after post-processing, there are fewer migration with zero or low correctness, and more migrations with high correctness. This indicates that the post-processing steps are effective in improving the correctness of migrations.
Note that the post processing steps are applied only when the migrations are not fully correct, and when that specific post processing step is applicable. For example, merge skipped code is applied to 260 migrations, among which 88 (34%) migrations have improved correctness with an average improvement of 46% per migration. On the other hand, infer async transformation is applied to 129 migrations, among which 45 (35%) migrations have improved correctness with an average improvement of 65% per migration.
4.2.4 RQ-B2: How much manual effort is required to complete the incomplete migrations?
In RQ-B1, we see that PyMigTool is able to fully migrate 32% of the migrations; however, this means that 68% of the migrations would still require the developer to perform edits or fixes to complete the migrations. In this RQ, we look into how much additional manual effort is required to correct the incomplete migrations.
4.2.4.1 RQ-B2: Approach
RQ-B1 shows that 233 migrations are completely correct after the automated migration process. The remaining 484 migrations will require fixes for the tests to pass. These 484 migrations are between 46 unique library pairs in 147 repositories. Attempting to fix all these migrations manually would be very time consuming and impractical. Therefore, we select one random migration from each library pair, while maximizing the number of unique repositories. This gives us a total of 43 migrations between 43 library pairs in 39 unique repositories.
For each of these migrations, we open the code using the PyCharm IDE at the version after PyMigTool migration. We refer to library documents and other online resources to understand each test error and how to fix it. After applying the edits, we runs the tests again and save the test report, which we refer to as TestReportmanual.
We estimate the development effort required for a migration by measuring the number of lines changed for the migration. For a migration, we compute a diff between the code before manual edit (i.e., the code after automating migration by PyMigTool) and the code after manual edits. From the diff, we calculate the number of changed lines of code by adding number of removed lines and number of added lines. We refer to this number as $MigLOC_{manual}$. Now, while this number gives us the absolute number of line changed, it does not give us the relative effort in comparison to the automated migration by PyMigTool. For example, a migration that required 50 changed lines where the tool performed 48 of them and we only have to edit 2 lines should not be considered the same as a migration that required 4 lines of code and we still have to edit 2 lines. Therefore, we also measure the number of lines changed by PyMigTool during the automated migration from the diff between the code before automatic migration and the code after migration by PyMigTool. We refer to this number as $MigLOC_{auto}$. From these two numbers, we calculate the manual migration effort as the ratio of manual changes to the total changes.
$\mathit{Effort}_{\mathit{manual}}=\frac{\mathit{MigLOC}_{\mathit{manual}}}{\mathit{MigLOC}_{\mathit{auto}}+\mathit{MigLOC}_{\mathit{manual}}}$ (1)
Similarly, automatic migration effort is calculated as the ratio of automatic changes to the total changes.
$\mathit{Effort}_{\mathit{auto}}=\frac{\mathit{MigLOC}_{\mathit{auto}}}{\mathit{MigLOC}_{\mathit{auto}}+\mathit{MigLOC}_{\mathit{manual}}}$ (2)
Note that the total migration effort is the sum of $\mathit{Effort}_{\mathit{manual}}$and $\mathit{Effort}_{\mathit{auto}}$, i.e., $\mathit{Effort}_{\mathit{manual}}+\mathit{Effort}_{\mathit{auto}}=1$.
4.2.4.2 RQ-B2: Results
Figure 10: Distribution of effort savings across manually fixed migrations.
The yellow dotted line shows median effort saving across all 24 migrations.
We apply manual edits to the selected 43 migrations. The manual edit effort fixes at least one failing test for 24 (56%) migrations, including 21 (49%) migrations that were completely fixed.
For the 21 migrations that we were able to fully correct through the manual edits, we measure $\mathit{Effort}_{\mathit{manual}}$and $\mathit{Effort}_{\mathit{auto}}$using Equation 1 and Equation 2, respectively. Figure 10 shows the proportion of automatic and manual changes made for each migration. The blue part of the stacked bar is the $\mathit{Effort}_{\mathit{auto}}$, while red part is $\mathit{Effort}_{\mathit{manual}}$. The yellow dotted line shows the median of $\mathit{Effort}_{\mathit{auto}}$  across all 21 migrations, which is 86%. This means that for a median migration, PyMigTool performed 86% of the changes, while the developers only needed to perform 14% of the changes.
4.3 Implications and discussion
We now discuss some of the issues we noticed with the LLM-based migration in PyMigTool and potential extensions of PyMigTool to resolve these issues in the section after this (Section 4.3.2).
4.3.1 Reoccurring issues in PyMigTool migrations
To understand current limitations of PyMigTool, we look at the manual fixes we had to perform. The examples we show here are from the original migrations from our experiments, but often adopted for clarity and conciseness. In the left-to-right or top-to-bottom comparisons, the changes are highlighted in red.
Original code
1class EncAlchemyField:
2 impl = BINARY
3
4 def __init__(self, key):
5 self.key = key
6 self.crypto=Crypto(key)
PyMigTool migration
1class EncTortoiseField:
2 SQL_TYPE = "BLOB"
3
4 def __init__(self, key):
5 self.key=key
6 self.crypto=Crypto(key)
Manual fix
1class EncAlchemyField:
2 SQL_TYPE = "BLOB"
3
4 def __init__(self, key):
5 self.key=key
6 self.crypto=Crypto(key)
Figure 11: Example of PyMigTool renaming a public class, and we manually fixing it back to original name in a sqlalchemy→tortoise-orm migration.
  1. Issue-1
    The LLM refactors public program elements: In the LLM migration round, PyMigTool migrates each file in isolation, which can lead to incorrect migrations. For example, in Figure 11, PyMigTool renames the public class EncryptedAlchemyBinaryField defined in the client project to EncryptedTortoiseBinaryField to match the target library’s name. This class is used in other parts of the code. Renaming the file prevents its callers from finding it, causing the errors in tests. Since PyMigTool does not pass the code that contains usage of the class in other files to the LLM, the LLM does not understand that it should not rename the class. Note that the prompt (2) does ask not to change any API names, however, the LLMs do not always strictly follow the instructions.
    Original code
    1gh_actor = os.getenv("GH_ACTOR", "nb")
    2gh_app_id = get_int_env_var("GH_APP_ID")
    PyMigTool migration
    1gh_actor = settings.get("GH_ACTOR", "nb")
    2gh_app_id = get_int_env_var("GH_APP_ID")
    Manual fix
    1settings.reload()
    2gh_actor = settings.get("GH_ACTOR", "nb")
    3gh_app_id = get_int_env_var("GH_APP_ID")
    Test code
    1@patch.dict(
    2 os.environ,
    3 {"GH_ACTOR": "testactor"}
    4)
    5def test_get_env_vars_optional(self):
    6 expected_result = EnvVars(
    7 "testactor",
    8 )
    9 result = get_env_vars(True)
    10 actual = str(result)
    11 expected = str(expected_result)
    12 self.assertEqual(actual, expected)
    Figure 12: Example of requiring to call an additional API in a python-dotenv→dynaconf migration.
  2. Issue-2
    The LLM does not have enough context: Figure 12 shows an example where the underlying LLM was not able to add an additional API call (settings.reload()) required for correct migration. This is required because the target library, dynaconf, loads the environment variables when the library is imported for the first time, and a reload is required if the environment variables are changed. The original code uses python-dotenv to load the environment variables, which reads the environment variables at runtime. Notice how the test, which is in another file thus does not get passed to the LLM, updates the environment variables before running the tests, which is after the libraries are imported, causing the tests to fail without the settings.reload() call in the migrated code. If the LLM was given this test code that affects the migration-related code, it may have been able to understand that the reload is required.
    Original code
    1try:
    2 d.update(toml.loads(r))
    3except TypeError:
    4 try:
    5 d.update(toml.load(r))
    6 except TypeError:
    7 d.update(r)
    PyMigTool migration
    1try:
    2 d.update(tomli.loads(r))
    3except TypeError:
    4 try:
    5 d.update(tomli.load(r))
    6 except TypeError:
    7 d.update(r)
    Manual fix
    1try:
    2 d.update(tomli.loads(r))
    3except TypeError:
    4 try:
    5 d.update(tomli.load(r))
    6 except AttributeError:
    7 d.update(r)
    Figure 13: Example of manually replacing exception type in a toml→tomli migration.
  3. Issue-3
    The LLM does not always adopt to the target library’s behavior: Figure 13 shows an example where the LLM did not recognize that the target library raises a different exception in a scenario, and therefore it was unable to replace the exception type correctly.
  4. Issue-4
    The LLM skips part of the code: Despite explicitly mentioning in the prompt not to skip parts of the code (2), the LLM still skips part of the code in migrations, specially when the code pattern is repetitive or does not have source API usage. We observe the same behavior in Section 3, which inspired us to add the Merge skipped code post-processing step in PyMigTool to merge the skipped code with the migrated code (see paragraph 4.1.3.1). Because this step uses some heuristics that often cannot identify all skipped code, we still had to manually fix some skipped code.
4.3.2 Possible extensions of PyMigTool
The problems we discuss above can potentially be mitigated by the following extensions of PyMigTool:
  1. Extension-1
    Post-processing for reverse refactoring: We can add a post processing step that detects renamed public APIs, and renames them back to the original names, potentially solving Issue-1. Refactoring detection tools, such as RefactoringMiner RefactoringMiner2 , can be used to detect such refactoring changes, and refactoring tools, such as RefMerge refmerge , can be used to apply the changes.
  2. Extension-2
    Including all dependant code with the prompt: This will allow the LLMs to know whether and how the code to migrate is used in other parts of the project, including the test code, potentially resolving Issue-2. This approach, however, can be impractical for the files that have a large number of dependencies, due to the LLM’s context length limitations.
  3. Extension-3
    Reprompting with test error messages: When we manually fixed the migration in Figure 13 related to Issue-3, we were able to quickly identify the issue just by looking at the test error messages. This means, systematically reprompting PyMigTool with the test error messages may help it to understand the context better and produce a correct migration, similar to how SOAR ni2021soar uses error messages to improve their program synthesis process (see LABEL:sec:relatedwork:transformation). While this can potentially help PyMigTool to produce more correct migrations, the challenge is processing the error messages and generating a meaningful prompt for the LLM. Note that SOAR considers only two pairs of libraries, and therefore the error messages are not as diverse as in our case.
  4. Extension-4
    Designing a sophisticated merge skipped code algorithm: our merge skipped code algorithm only looks at the size of the skipped code and usages of source APIs (see paragraph 4.1.3.1). A more sophisticated algorithm that considers the context and the structure of the code, such as removed functions/classes, may potentially improve this step reducing the effect of Issue-4.
While the above steps can potentially mitigate the problems we identified, the positive side is that manually fixing the incomplete migrations was also simple, so further effort in more automation may not be necessary. For example, in Figure 13 (Issue-3), we were able to catch the issue easily just by looking at test error messages, and the fix was just replacing the exception type in the code. While reprompting with error messages may potentially fix the problem, it would also increase the cost of migration. The skipped code cases (Issue-4) are also fairly simple to resolve with modern IDEs. We use PyCharm’s change viewer during our manual fix, which allows us to easily identify the skipped code, and merge them with a click of mouse.
The case in Figure 12 (Issue-2) required us to look deep into the library documentation and was time consuming. However, the original developers are likely to be more familiar with the libraries and their behavior. Therefore, they may have been be able to fix the issues easily.
In some cases where migrations are initially marked as incorrect, the issues are related to configuration rather than code changes. For example, in a migration from chardet→cchardet, adjusting the Python version from 3.11 to 3.9 resolved the test failures. Similarly, we made a migration from flask→sanic successful by installing an additional dependency, sanic-ext, with no changes to the migrated code. The developers of the projects are likely to have been aware of such configuration requirements, therefore, they may be been able to fix the issues easily.
Overall, we find that PyMigTool automatically completes migration for a large portion of cases (32% migrations) and does most of the complex migration work (86% changes) for the remaining migrations, often leaving simple edits for the developers to perform (14% changes). Further improvements in migration process may potentially help in reducing the manual effort. However, the manual effort seems fairly minimal in most cases, specially for the developers who are familiar with the libraries and the project. We stress that the author of this thesis, who has no prior knowledge of the projects, performed all the manual edits alone without help of any other developers. This indicates that the manual edit part can potentially be done by someone new to the project, or even be outsourced outside of the project team.
4.4 Threats to validity
This section discusses the potential threats to the validity of our study and the steps we take to mitigate them.
4.4.1 Internal validity
One author performs all manual fixes and their evaluation, which may introduce subjectivity or bias. To mitigate this, we follow a consistent process for manual fixes and rely on objective test outcomes as the primary measure of correctness.
Configuration and environment setup also pose internal threats. Some migrations initially fail due to environment or dependency issues rather than code errors. We address these by adjusting configurations when possible, but environment-specific factors, such as Python and library version, operating system, may still affect the results.
4.4.2 External validity
We focus on Python projects that are well-tested and actively maintained, as selected through specific filters. This focus may limit the generalizability of our findings to projects with less test coverage, different maintenance status, or other programming languages. The set of library pairs and migration scenarios is also limited by the availability of analogous libraries and the manual validation process. As a result, the observed effectiveness of PyMigTool may not fully generalize to all migration contexts. That said, the size of the dataset (717 migrations) and the diversity of selected projects (175 repositories) provide a reasonable basis for evaluating the tool’s effectiveness in a variety of migration scenarios.
4.4.3 Construct validity
A key threat to construct validity is the correctness of the automated and manual evaluation process. We use automated test suites to assess migration correctness, but these tests may not cover all possible behaviors or edge cases in the migrated projects. To minimize this, we select projects with at least 80% lines of code covered by tests, which helps ensure that the tests are comprehensive enough to provide a reliable measure of correctness.
Our manual migration effort calculation metric (Equation 1) is based on lines of code changed, which may not capture all aspects of migration effort, such as time spent understanding or debugging the migration. We use it as a measurable objective proxy for the effort required.
5 Related Work
There has been a significant amount of research aimed at automating library migration. We discuss these techniques in this section.
The majority of the existing library migration techniques focus on identifying API mappings between two analogous libraries. Some of these techniques use mine existing library migrations to identify the API mappings teyton2012mining ; teyton2013automatic ; alrubaye2019use ; alrubaye2019migrationminer ; alrubaye2020learning . Some techniques see the requirement of existing migrations as a limitation, and therefore use other techniques, such as machine learning alrubaye2018automating ; miningAnalogicalAPIs and natural language processing ni2021soar . These techniques only identify the API mappings, but do not perform client code transformation, which is a separate challenge.
SOAR ni2021soar uses API documentation to identify API mappings, and then uses the mappings and unit tests to guide program synthesis for client code transformation. SOAR has several limitations that prevent it from being a general solution for library migration. SOAR is evaluated on only two pairs of deep learning libraries, focusing on their migration of neural network models. These APIs are usually called in a sequence, where the output of one call is used as input to the next call. The nature of these APIs allows SOAR to check the program state after each API call, which is not the case for most libraries. Additionally, given the limited number of libraries, their technique relies on the library-specific error message format to further guide the synthesis process, requiring extensive changes and technique adaptations for running SOAR on other libraries. SOAR supports one-to-many code changes for a small set of pre-determined APIs, whose information is hard-coded for the transformation process. SOAR also times out for 20% of the evaluated migrations, even for small code snippets, the largest being 183 lines of code.
Zhou et al. hapim proposed HaPiM, which first trains an unsupervised machine learning model capable of API mapping, and then uses the API mappings to guide an LLM in code transformation. Similar to Chen et al. miningAnalogicalAPIs , HaPiM learns from API usage patterns of source and target libraries in their respective client projects, therefore does not require any existing migration examples. For the code transformation, they use the API mappings to prompt an LLM to replace source API calls with target API calls. Their approach outperforms MigrationMapper alrubaye2019use and GPT-3.5 Turbo on the BLEU bleu and CodeBLEU codebleu metrics evaluated on 5 Java library pairs. A significant limitation of HaPiM is that it works with small independent code snippets that use the source library APIs, and results in an equivalent snippet that uses the target library APIs. This is as opposed to real-life code where the source API usages are interleaved with other types of code, such as variable declarations, control statements, and other library usages. This means that to use HaPiM, developers need to first extract code snippets that contains only the source library API usages from the client code, then use HaPiM to replace them with target library APIs, and then re-integrate the transformed code back into the client code, which is not practical for large-scale library migrations.
Nikolov et al. nikolov2025google describe Google’s experiences in using LLMs for several internal code transformation tasks, including Joda Time→Java Time migrations. They do not attempt to provide a general solution for library migration; rather, they use LLMs to perform the migration they need. Therefore, their prompts contains instructions and contextual information tailored specifically for Joda Time→Java Time migrations. Nonetheless, they show that using LLMs reduced migration time by at least 50% compared to manual migration.
The above shows that there have been different approaches to automate the library migration process. However, these techniques come with one or more limitations that prevent them from being a general solution for library migration. Our tool, PyMigTool, resolves these limitations and can automate the entire library migration process for any arbitrary analogous library pair (Section 4). We evaluate PyMigTool on a large and diverse dataset of 717 migrations between 48 analogous library pairs, and show that PyMigTool can successfully perform most of the migration with no or minimal manual intervention from the developer.
6 conclusion
In this paper, we explore the capabilities of LLMs for automating Python library migration and build PyMigTool, a practical tool that combines LLMs with static and dynamic program analysis to automate the full migration process. In the first phase of the work, we conduct an empirical study using three LLMs on the PyMigBench dataset to understand how well LLMs can perform library migration. We find that the best performing LLM, 4o, correctly migrates 94% of the individual code changes. At the migration level, it perfectly migrates 57% of the migrations, and 94% of them have at least one correctly migrated code change. We also run unit tests for a subset of the migrations, and find that 64% of the migrations by 4o have the same sets of tests passing in the developer’s migration and the LLM’s migration. A much lower cost LLM Mini and the free LLama also perform relatively well, with both correctly migrating 89% of code changes. Overall, we find that LLMs can handle complex code changes, therefore, we conclude that LLMs are a promising approach for automating library migration. However, we also identify some challenges that need to be addressed to make LLM-based migration practical.
We use these insights to build PyMigTool, a CLI tool that automates the full library migration process in phase two. Our evaluation on 717 simulated migrations across 175 diverse projects and 48 library pairs shows that PyMigTool can fully automate a large portion of migrations (32%), with most remaining cases requiring only minor manual intervention (median 14% changes). The tool’s layered approach, including post-processing steps, improves LLM migration correctness (from 65% to 86%) and reduces developer effort. While PyMigTool is not able to handle every migration scenario perfectly, it lowers the barrier for large-scale library migration in real-world projects. Overall, PyMigTool represents a meaningful step toward making automated library migration accessible and effective for the software engineering community.
Acknowledgements.
We thank Dr. Ildar Akhmetov, Northeastern University, for his contributions in to the early to the ideas of the work.

### v2

@@
 Library migration is the process of replacing a library with a similar one in a software project.
-Manual library migration is time consuming and error prone, as it requires developers to understand the Application Programming Interfaces (API) of both libraries, map equivalent APIs, and perform the necessary code transformations.
+Manual library migration is time-consuming and error-prone, as it requires developers to understand the Application Programming Interfaces (API) of both libraries, map equivalent APIs, and perform the necessary code transformations.
 Due to the difficulty of the library migration process, most of the existing automated techniques and tooling stop at the API mapping stage or support a limited set of libraries and code transformations.
 In this paper, we develop an end-to-end solution that can automatically migrate code between any arbitrary pair of Python libraries that provide similar functionality.
-Due to the promising capabilities of Large Language Models (LLMs) in code generation and transformation, we use LLMs as the primary engine for migration.
-Before building the tool, we first study the capabilities of LLMs for library migration on a benchmark of 321 real-world library migrations.
-We find that LLMs can effectively perform library migration, but some post-processing steps can further improve the performance.
-Based on this, we develop PyMigTool, a command line application that combines the power of LLMs, static analysis, and dynamic analysis to provide accurate library migration.
-We evaluate PyMigTool on 717 real-world Python applications that are not from our benchmark.
-We find that PyMigTool can migrate 32% of the migrations with complete correctness.
-Of the remaining migrations, only 14% of the migration-related changes are left for developers to fix for more than half of the projects.
+Given the promising capabilities of Large Language Models (LLMs) in code generation and transformation, we use them as the primary engine for migration.
+Before building the tool, we first study the capabilities of LLMs for library migration on a benchmark of 314 real-world library migrations.
+We find that the best performing LLM, GPT-4o, can correctly perform 57% of the migrations.
+We also identify post-processing steps that can further improve the performance.
+Based on this, we develop MigrateLib, a command-line application that combines the power of LLMs, static analysis, and dynamic analysis to provide accurate library migration.
+We evaluate MigrateLib on 717 migration scenarios in 175 real-world Python applications that are not from our benchmark.
+We find that MigrateLib can migrate 32% of the migrations with complete correctness without any developer intervention.
+Of the remaining migrations, 73% of the migration-related changes are automatically done by the tool for more than half of the projects.
 1 Introduction
@@
 The process involves identifying the library usage in the codebase, finding the corresponding Application Programming Interfaces (APIs) in the new library, and replacing the old API usage with the new API usage.
-This can be tedious and error-prone, especially when the library is extensively used in a large codebase kula2018developers .
-Researchers have proposed various techniques to automate the library migration processes.
-Most of these techniques, however, are limited to only finding the analogous APIs (API mapping) teyton2013automatic ; alrubaye2018automating ; alrubaye2019use ; miningAnalogicalAPIs ; zhang2020deep ; di2025deepmig , but do not replace the old API usage in the codebase with the new one (transformation).
-To find analogous APIs, some of these techniques mine existing migrations teyton2013automatic ; alrubaye2018automating ; alrubaye2019use while others use machine learning models miningAnalogicalAPIs ; di2025deepmig and natural language processing techniques ni2021soar .
-To the best of our knowledge, there are two techniques that perform transformation ni2021soar ; nikolov2025google but both techniques evaluate their approaches on a small set of library pairs without providing a tool that can be used in practice.
-In recent years, Large Language Models (LLMs) have shown that they can be effective in various software engineering tasks ozkaya2023application ; fan2023large ; wang2024software , including the necessary upstream tasks for library migration, such as code generation nguyen2022empirical ; peng2023impact ; murali2023codecompose , code comprehension nam2024using , and code transformation wei2023copiloting ; xia2023automated ; zhou2023hybrid .
+This can be tedious and error-prone, especially when the library is extensively used in a large codebase Kula et al.
+(2018).
+Researchers have proposed various techniques to automate the library migration process.
+Most of these techniques, however, are limited to only finding the analogous APIs (API mapping) Teyton et al.
+(2013); Alrubaye and Mkaouer (2018); Alrubaye et al.
+(2019b); Chen et al.
+(2019); Zhang et al.
+(2020); Di Rocco et al.
+(2025), but do not replace the old API usage in the codebase with the new one (code transformation).
+To find analogous APIs, some of these techniques mine existing migrations Teyton et al.
+(2013); Alrubaye and Mkaouer (2018); Alrubaye et al.
+(2019b) while others use machine learning models Chen et al.
+(2019); Di Rocco et al.
+(2025) and natural language processing techniques Ni et al.
+(2021).
+To the best of our knowledge, there are two techniques that actually perform code transformation Ni et al.
+(2021); Nikolov et al.
+(2025), but both techniques evaluate their approaches on a small set of selected library pairs without providing a tool that can be used in practice.
+In recent years, Large Language Models (LLMs) have shown that they can be effective in various software engineering tasks Ozkaya (2023); Fan et al.
+(2023); Wang et al.
+(2024), including the necessary upstream tasks for library migration, such as code generation Nguyen and Nadi (2022); Peng et al.
+(2023); Murali et al.
+(2023), code comprehension Nam et al.
+(2024), and code transformation Wei et al.
+(2023); Xia et al.
+(2023); Zhou et al.
+(2023a).
 This suggests that LLMs may be capable of performing library migration.
-In this paper, we leverage LLMs to explore and address the challenges of automating library migration in two phases.
-First, we conduct an experiment using an existing library migration benchmark, PyMigBench pymigbench , to determine the capability and challenges of using LLMs for library migration.
-We find that LLMs are generally capable of performing migration; however, some post processing can further improve the migration results.
-Second, based on our findings in first phase, we build a CLI tool, PyMigTool, that automates the full library migration process by combining the power of LLMs with post processing using static and dynamic program analysis.
-We run an extensive evaluation of PyMigTool on a set of 717 new, simulated migrations.
-Phase one – Explore LLMs:
-In this phase, we investigate the capabilities and challenges of using LLMs for library migration.
-We use LLama 3.1, GPT-4o mini, and GPT-4o to perform 321 migrations from PyMigBench.
-We then measure the correctness of the migrations by comparing the LLM’s migrations with the original developer’s migrations.
-We find that the three LLMs can correctly migrate 26%, 49%, and 57% of the migrations respectively.
-Overall, we find that LLMs are capable of library migration, and as expected, the more powerful LLM (GPT-4o) performs better than the less powerful LLMs (LLama 3.1 and GPT-4o mini).
-We also identify two major challenges that make solely relying on LLMs in a practical tool ineffecive: (1) We provide the LLM with only code files that directly use the source library.
-However, for a given project, there may be other files that depend on the migrated code and also require changes during the migration.
+However, studies have also identified limitations of LLMs in other software engineering tasks, especially generating incorrect or incomplete code Fan et al.
+(2023).
+These previous efforts suggest that LLMs may be capable of performing library migration and that LLM-based migration tooling could be effective.
+However, we still need further empirical evidence to support this claim.
+In this paper, our goal is to investigate the effectiveness of LLMs in library migration to allow informed design decisions for LLM-based migration tooling.
+We choose Python as the programming language for our study, as it is one of the most popular programming languages and has a rich ecosystem of libraries.
+Accordingly, we approach this in three phases as illustrated in Figure 1.
+In Phase 1, our goal is to systematically investigate the capabilities and challenges of using LLMs for library migration.
+We conduct an experiment using 314 migrations from PyMigBench Islam et al.
+(2023), a Python library migration benchmark we previously built.
+We experiment with three LLMs (Llama 3.1, GPT-4o mini, and GPT-4o) and measure the correctness of the migrations by comparing the LLM’s migrations with the original developer’s migrations.
+We find that the three LLMs can correctly migrate 26%, 49%, and 57% of the migrations, respectively.
+Using an alternative test-based evaluation on a subset of the projects that have unit tests, we find that 36%, 52%, and 64% of the LLM migrations maintain the same test results as the original developer migrations for the three LLMs, respectively.
+We also study the types of migration-related code changes LLMs excel at vs.
+struggle with.
+Overall, we find that LLMs are capable of library migration, and as expected, the more powerful LLM (GPT-4o) performs best.
+We also identify four major challenges that make solely relying on LLMs in a practical tool less effective: (1) The files that need migration are not always known beforehand.
 (2) All LLMs often return only part of the code, skipping parts of the code that do not require migration.
-These limitations, however, can be addressed by post-processing the LLM’s results.
-We use these insights to improve the migration process in the next phase.
-Phase two – Build PyMigTool:
-In this phase, we build PyMigTool, a Command Line Interface (CLI) that automates the full library migration process using LLMs and post processing using static and dynamic program analysis.
-PyMigTool takes as 5 inputs from the user (1) A project that requires migration, with an existing test suite with high coverage.
-(2) A requirements file that lists the dependencies of the project.
-(3) A source library that is used in the project.
-(4) A target library and its version that the user wants to migrate to.
-(5) Name of an LLM that will be used for the migration.
-By building a dynamic call graph, PyMigTool identifies all files that require migration and sends those to the LLM.
-PyMigTool then verifies the results of the migration by comparing the test results before and after migration.
-If any previously passing tests now fail, PyMigTool performs a Post-processing round where it checks if the LLM skipped any code and merges it back.
-It also detects if any of functions were made async during the migration, and if so, it finds other parts of the project that uses these functions and makes them async as well.
-Finally, it runs the unit tests again and saves the results
-We evaluate PyMigTool using a new purely test-based evaluation setup, as opposed to the benchmark based evaluation we did in the first phase.
-This allows us to evaluate new migrations without the need for a ground truth developer migration.
-We evaluate PyMigTool on a set of 717 simulated migrations in 175 unique projects.
-By simulated migrations, we mean that this migration did not exist in the project’s history; instead, we find source libraries used in a project and migrate the project to an analogous target library.
-We ensure that the projects we choose have high test coverage, so that the evaluation is reliable.
-We demonstrate twi different evaluation setups.
-First, we measure how well PyMigTool perform as it is on the simulated migrations, without any manual intervention.
-While the best case scenario is that PyMigTool can migrate the project without any issues, a more practical expectation is that PyMigTool would do most of the task correctly, but there are some parts that require manual intervention.
-Therefore, in a second evaluation, we measure how much manual work can fix the incorrect migrations.
-We randomly select a subset of 43 diverse migrations that were not fully correct, and we manually fix them.
-We find that PyMigTool was able to successfully migrate 32% migrations without any manual intervention.
-We manually fix 43 of the remaining 484 migrations, and were able to fix 21 (49%) of them.
-In terms of modified lines of code, the manual work is only 14% of the total migration-related changes.
+(3) For a given project, a file that we migrate may have other files that depend on it which may require changes during migration.
+Specifically, if any functions are made asynchronous during the migration, all parts of the code that use those functions must also be made asynchronous.
+(4) The inconsistencies between the versions between the libraries used in the project and those assumed by the LLM can lead to incorrect migrations.
+We find that the identified limitations from Phase 1 can be addressed by improving the prompt, some pre-processing and post-processing the LLM results.
+Accordingly, in Phase 2, we build a CLI tool, MigrateLib, that automates the full library migration process using LLMs and post-processing using static and dynamic program analysis.
+MigrateLib takes in a complete Python project with an existing test suite and automatically performs the migration for the specified library pair.
+By building a dynamic call graph, MigrateLib first identifies all files that require migration and sends those to the LLM via a prompt that instructs the LLM to perform the migration.
+MigrateLib then verifies the results of the migration by comparing the test results before and after migration.
+If any previously passing tests now fail, MigrateLib performs a post-processing step where it checks if the LLM skipped any code and merges it back.
+It also detects whether any functions were made async during the migration, and if so, finds other parts of the project that use those functions and makes them async as well.
+Finally, it runs the unit tests again to determine the final migration status.
+PyMigBenchPhase 1 Explore LLMs for Library MigrationLlama 3.1GPT-4o miniGPT-4oPhase 2 Build MigrateLib Phase 3 Evaluate MigrateLibGPT-4oSimulated migrations314 migrations717 migrations
+Figure 1: Overview of the contributions.
+In Phase 3, we evaluate MigrateLib using a new purely test-based evaluation setup, as opposed to the benchmark-based evaluation we did in Phase 1.
+This allows us to conduct a larger-scale evaluation and also evaluate new migrations without the need for ground truth developer changes.
+We run MigrateLib on a set of 717 simulated migrations in 175 unique projects.
+We define a simulated migration as one that did not exist in the project’s history.
+Specifically, we find source libraries used in a project and migrate the project to an analogous target library.
+We ensure that the projects we choose have high code coverage (above 80%), so that the evaluation is reliable.
+Given that GPT-4o had the best performance in Phase 1, we conduct this evaluation only using GPT-4o.
+We find that MigrateLib is able to successfully complete 32% migrations without any manual intervention (i.e., all tests that pass before migration still pass after migration).
+To understand how much manual work is required to fix the remaining 484 migrations that the LLM left in an incomplete stage, we randomly select a subset of 43 diverse migrations and attempt to manually fix them.
+We measure the amount of manual work required to complete these migrations using the number of lines of code we need to manually edit.
+We find that we are able to fix 24 (56%) of our sampled migrations, where the manual work is only 27% of the total changes needed for the migration.
+The main reason we could not fix some of the remaining migrations is that the projects had complex logic that required a deep understanding of the project context, which was difficult to achieve.
+Overall, our work provides a combination of empirical and engineering contributions.
+This paper is an extended version of our prior conference paper Islam et al.
+(2025) that only focused on Phase 1.
+Phases 2 and 3 are novel contributions of this work.
 To summarize, the contributions of this paper are:
   1. 1.
-    We conduct an initial detailed experiment to explore the use of LLMs for library migration on an existing migration benchmark that provides us a ground truth of previously performed developer migrations.
+    An empirical study to explore the use of LLMs for library migration by comparing LLM changes to 314 benchmark developer changes.
+Our analysis systematically highlights the types of migration-related code changes LLMs excel at vs.
+struggle with.
   2. 2.
-    Based on the insights gained from our initial experiment, we build a CLI tool, PyMigTool, that automates the full library migration process using a combination of LLMs and post processing using static and dynamic program analysis.
+    We design and build an open-source end-to-end CLI tool, MigrateLib, that automates the full library migration process using a combination of LLMs and post-processing using static and dynamic program analysis.
   3. 3.
-    We run an extensive test-based evaluation of PyMigTool on a set of 717 simulated migrations, and show that it works well with no or minimal manual editing for a diverse set of library pairs.
+    We design a new evaluation setup that allows us to evaluate MigrateLib on a large scale and also on newer projects.
+We run an extensive test-based evaluation of MigrateLib on a set of 717 simulated migrations, and show that it works well with no or minimal manual editing for a diverse set of library pairs.
+The source code of MigrateLib, the experimental data, and the detailed results of our experiments are available at https://figshare.com/account/articles/29602517.
 2 Background and Terminology
@@
 2.1 Library migration
-Library migration is the process of updating a software project to replace a used library with another one that provides similar functionality teyton2012mining .
-The source library is the one being replaced, and the target library is the one that replaces it teyton2012mining .
-One migration instance refers to a commit in a repository where a migration happened from a specific source library to a target library pymigbench , denoted using the notation source→target.
+Library migration is the process of updating a software project to replace a used library with another one that provides similar functionality Teyton et al.
+(2012).
+The source library is the one being replaced, and the target library is the one that replaces it Teyton et al.
+(2012).
+One migration instance refers to a commit in a repository where a migration happened from a specific source library to a target library Islam et al.
+(2023), denoted using the notation source→target.
 For example, the commit b0607a26 in repository openstack/ironic is a retrying→tenacity migration.
-A migration-related code change, or code change for brevity, is a minimal replacement of source library APIs with target library APIs that cannot be meaningfully reduced further without losing the semantics of the change pymigbench .
+A migration-related code change, or code change for brevity, is a minimal replacement of source library APIs with target library APIs that cannot be meaningfully reduced further without losing the semantics of the change Islam et al.
+(2023).
 A migration contains one or more code changes.
-The lines between the boxes in 1(c) show code changes.
-For example, segment P1 in 1(a) is replaced by segment D1 in 1(b).
+The lines between the boxes in 2(c) show code changes.
+For example, segment P1 in 2(a) is replaced by segment D1 in 2(b).
 We use subscripts pre, dev, and llm to denote data before migration, after developer’s migration, and after an LLM’s migration, respectively.
-For example, codepre, codedev, and codellm denote three states of a code.
+For example, Codepre, Codedev, and Codellm denote three states of a piece of code.
 Changedev and Changellm denote code changes by developers and an LLM, respectively.
 2.2 PyMigBench and PyMigTax
-PyMigBench is a dataset of real-world Python library migrations pymigbench , mined from version control history.
-We use the latest available version, version 2.2.5 pymigtax in our experiments.
+We use two of our prior artifacts: PyMigBench Islam et al.
+(2023) and PyMigTax Islam et al.
+(2024), in this work which we briefly describe next.
+PyMigBench is a dataset of real-world Python library migrations, which we mined from open source GitHub repositories.
+We use the latest version, version 2.2.5, in our experiments in this paper.
 The dataset has 321 migrations and 2,989 migration-related code changes.
 It includes a detailed description of each code change, including line numbers and API names, which facilitates our evaluation.
-PyMigTax is a taxonomy of migration-related code changes pymigtax , built using the first version of PyMigBench.
-The authors validated its generalizability on additional third-party data.
+PyMigTax is a taxonomy of migration-related code changes Islam et al.
+(2024), built using the first version of PyMigBench.
+In our prior work, we validated its generalizability on additional third-party data.
 PyMigTax describes a code change based on three dimensions: (1) Program elements: the types of program elements involved in the change (e.g., function call and attribute); (2) Cardinality: how many source APIs are replaced with how many target APIs.
 For example, one-to-many cardinality means one source API is replaced with multiple target APIs; one-to-many, many-to-one, and many-to-many cardinalities are commonly referred to as higher-cardinality; and (3) Properties: Additional properties to describe the code change (e.g., element name change when the source and target APIs have different names.).
-The code change data in PyMigBench is annotated with the PyMigTax categories.
+The code change data in PyMigBench is already annotated with the PyMigTax categories.
 We use PyMigTax categories (shown in Table 3) to understand the types of code changes LLMs can handle.
-The PyMigTax paper pymigtax has full category descriptions.
-3 Phase one: Exploring the capabilities of LLMs for Python library migration
-We use three off-the-shelf LLMs – LLama 3.1 llamapaper , GPT-4o mini openaimodels , and GPT-4o openaimodels – to perform migrations on the PyMigBench dataset.
-We refer to these models as LLama, Mini, and 4o in the rest of the paper, respectively.
-We evaluate the correctness of the migrations in two ways.
-First, we compare the LLM-migrated code with the developer-migrated code in PyMigBench and report correctness at two granularity levels: the full migration and the individual code changes.
-We use PyMigTax to identify which types of code changes the LLMs can/cannot handle.
-Second, for the subset of migrations with available unit tests, we run the available tests to assess the correctness of the migrated code.
+The PyMigTax paper Islam et al.
+(2024) has full category descriptions.
+3 Phase 1: Exploring LLMs for library migration
+We first systematically explore the capabilities of LLMs for library migration, while understanding how they handle different types of migration-related code changes.
+We first explain our setup and methods, and then present the results.
+3.1 Experiment setup
 We answer the following research questions in this phase:
-  1. RQ-A1
+  1. RQ-1.1
     How similar are the LLM migrations to the benchmark migrations?
@@
 We automatically check if the LLM was able to correctly perform all expected changes, while accounting for refactoring and alternative correct changes through manual review.
-  2. RQ-A2
+  2. RQ-1.2
     How many migrations pass unit tests? To evaluate run-time correctness of the migrated code, we use a second evaluation strategy where we run any available unit tests.
-3.1 Experiment Setup
-3.1.1 Models
-Table 1: Description of the models we use in our experiment.
-LLama 3.1 GPT-4o mini GPT-4o
-Used version 3.1-70B-Instruct 2024-07-18 2024-08-06
-Knowledge cutoff December 2023 October 2023 October 2023
-Model Type Open-source Proprietary Proprietary
-# training parameters 70 billion No official info No official info
-Context length 128K tokens 128K tokens 128K tokens
-Output limit 4096 tokens  [^1] LLama 3.1 has an input+output limit of 8192 tokens.
-Our prompt and response are approximately equal, therefore we consider the output limit to be half (4096).
-16,384 tokens 16,384 tokens
-Cost  [^2] Cost per million input token at the time of our experiments: July and August 2024 Free \$0.15 \$2.5
-Table 1 shows the three off-the-shelf models we use for our library migration task.
-The first model is Meta’s LLama 3.1, a free and open-source LLM llamapaper .
-The other two models are proprietary: OpenAI GPT-4o mini and GPT-4o openaimodels .
-GPT-4o mini replaces OpenAI’s previous GPT-3.5 Turbo and is more cost-effective.
-GPT-4o is the most powerful of the OpenAI models, while also being the most expensive openaimodels .
-We use these models to migrate the code in 321 migrations from our PyMigBench dataset (pymigbench ).
-3.1.2 Data preparation
-At the time of conducting this experiment, two repositories from PyMigBench were no longer public, preventing us from using 2 migrations.
-We clone the remaining repositories containing 319 migrations.
-For each migration commit, we use PyDriller PyDriller to extract the content of each file that has recorded migration-related code changes, both before and after the migration.
-We use Python’s built-in ast.parse function docast to ensure that the files are syntactically valid code and discard 5 migrations having files with syntax errors.
+We start with the 321 migrations in PyMigBench for our experiments, then discard 2 migrations whose repositories are no longer public, and 5 migrations where either the pre-migration or post-migration file has syntax errors.
 We conduct our experiments with the remaining 314 migrations containing 2,910 migration-related code changes.
-3.1.3 Migration
-For each file having code changes, we use the LLM prompt template in 1 to migrate the code.
+Listing 1: LLM Migration Prompt
+The following Python code uses library <source-lib>. Migrate this code to use library <target-lib> instead. In the output, first explain the changes you made. Then, provide the modified code.
+Do not make any changes to the code that are not related to migrating between these two libraries. Do not refactor. Do not reformat. Do not optimize. Do not
+change coding style.
+Provided code:
+<premig-code>
+3.1.1 Performing migration
+We use three off-the-shelf LLMs – Llama 3.1-70B-Instruct Dubey et al.
+(2024), GPT-4o-mini-2024-07-18, and GPT-4o-2024-08-06 OpenAI (2024) – to perform migrations on the PyMigBench dataset.
+We refer to these models as Llama, Mini, and 4o in the rest of the paper, respectively.
+For each file having code changes in a given migration, we use the LLM prompt template in 1 to migrate the code.
 The prompt is designed to ensure that the LLM performs focused, controlled migrations while maintaining transparency.
@@
 The prompt also aims to restrict the LLM from making unrelated modifications such that we remain focused on the task of library migration.
-Since LLMs are non-deterministic atil2024llmstabilitydetailedanalysis , we use a temperature of 0 and run each migration 10 times for each model.
-We calculate the difference between the runs to understand the variability in the generated code.
-We use git-diff gitdiff to compute a diff between each pair of the 10 runs and normalize the number of different lines by dividing it by the total number of lines in Filepre.
+Since LLMs are non-deterministic Atil et al.
+(2024), we use a temperature of 0 and run each migration 10 times for each model.
+To understand the extent of variability in the generated code, we calculate the difference between the runs.
+We use git-diff to compute a diff between each pair of the 10 runs and normalize the number of different lines by dividing it by the total number of lines in Filepre.
 We find that on average, two runs on the same file are only 6.6%, 4.0% and 4.8% different for Llama, Mini, and 4o, respectively.
 Given this low variability and the high effort involved in manual validation, we randomly select one run for each model to evaluate the LLMs’ results.
-Our artifact contains the LLM’s migrations for all 10 runs.
-3.2 RQ-A1 How similar are the LLM migrations to the benchmark migrations?
-3.2.1 RQ-A1 Approach
-In this RQ, we assess each LLM’s migration correctness by comparing its code changes (Changellm) to those made by developers (Changedev), while manually judging potential alternative changes.
-A migration may involve multiple files, each with multiple migration-related code changes.
-We evaluate each code change individually, and aggregate the results to determine the correctness of the whole migration.
-We now explain the process using the example in 1(c) that shows a migration from the library requests to aiohttp.
-Listing 1: LLM Migration Prompt
-1The following Python code uses library <source-lib>.
-2Migrate this code to use library <target-lib> instead.
-3In the output, first explain the changes you made. Then, provide the modified code.
-4
-5Do not make any changes to the code that are not related to migrating between these two libraries.
-6Do not refactor. Do not reformat. Do not optimize. Do not change coding style.
-7
-8Provided code:
-9<premig-code>
+Our artifact contains all LLM migrations for all 10 runs.
+3.1.2 Matching LLM migrations with developer migration
+In RQ-1.1, we compare the LLM migrations with the developer migrations in PyMigBench.
+We do this for each code change, and then aggregate the results at the migration level, while also condidering the additional non-migration-related changes made by the developer and the LLM.
+We now explain the process using the example in 2(c) that shows a migration from the library requests to aiohttp.
 (a) Before migration (Filepre)
@@
 The blue check marks denote correct changes, while the orange crosses denote incorrect changes.
-3.2.1.1 Match code changes
-The goal of this step is to compare the code changes the LLM made (Changellm) with the developer ground truth changes (Changedev), as well as manually finding alternative correct changes.
-Auto change matching: We first attempt to automatically match as many changes as possible while ensuring precision i.e., not falsely marking a Changellm as correct.
+A. Auto change matching
+We first attempt to automatically match as many LLM changes (Changellm) as possible with developer change (Changedev), while ensuring precision i.e., not falsely marking an incorrect Changellm as correct.
 Therefore, we consider only exact syntactic matches between Changedev and Changellm, ignoring formatting differences.
@@
 If there are potentially multiple matches, we check the containing function and use proximity based heuristics demonstrated below.
-We ensure that the matched node sets translate to the same code string using ast.unparse docast .
-Consider Changedev D1 in 1(c) which uses ClientSession() and get() from aiohttp.
+We ensure that the matched node sets translate to the same code string using ast.unparse.
+Consider Changedev D1 in 2(c) which uses ClientSession() and get() from aiohttp.
 In Filellm, these APIs appear in multiple locations, but only Lines 15, 16, 21, and 22 are in the same containing function fetch_flights as Changedev.
@@
 By the end of the auto change matching step, we get a set of matched pairs of Changedev and Changellm, a set of unmatched Changedev (uChangedev), and a set of unmatched Changellm (uChangellm).
-Manual review: The manual review step focuses on reviewing uChangedev and uChangellm.
-We use a three-way diff viewer to manually compare Filepre, Filedev, and Filellm for each file with at least one uChangedev or uChangellm.
+B. Manually reviewing unmatched changes
+The automatic matching ensures that no change is incorrectly matched, but it may miss some correct matches due to syntactic differences, or alternative correct implementations.
+The manual review step focuses on reviewing uChangedev and uChangellm.
 We use the online API documentation as well as the source code of the source and target libraries to understand the correct API usage in the context of the file.
-For 1(c), we manually find that L4 is a correct alternative to D2, noting that the LLM omitted the default value (True) of the allow_redirects argument, unlike the developer.
+For 2(c), we manually find that L4 is a correct alternative to D2, noting that the LLM omitted the default value (True) of the allow_redirects argument, unlike the developer.
 L8 is a correct replacement of D3, where the LLM performed a refactoring.
@@
 We perform a multi-reviewer manual review to ensure the quality of the matching process.
-The thesis author first manually reviews 190 code changes from one model.
-Two collaborators then review about half of those code changes each, not knowing the thesis author’s decisions.
-Therefore, each code change is reviewed by two reviewers independently.
-We then measure the agreement using Cohen’s Kappa cohen1960coefficient to quantify the level of agreement between the reviewers.
+The first author first manually reviews 190 code changes from one model.
+Two other authors then review about half of those code changes each independently.
+We then measure the agreement using Cohen’s Kappa Cohen (1960) to quantify the level of agreement between the reviewers.
 We do not reach substantial agreement, so we discuss and resolve disagreements and update our coding guideline.
-The two pairs of reviewers then independently review another 111 code changes, achieving a Cohen’s Kappa score of 0.83 (almost perfect) and 0.73 (substantial)  landis1977measurement .
-Accordingly, we proceed with only the thesis author reviewing the remaining changes.
-3.2.1.2 Determine migration status
-Based on the matching of Changedev and Changellm, we determine each code change status as follows (example changes refer to 1(c)).
+The two pairs of reviewers then independently review another 111 code changes, achieving a Cohen’s Kappa score of 0.83 (almost perfect) and 0.73 (substantial)  Landis and Koch (1977).
+Accordingly, we proceed with only the first author reviewing the remaining changes.
+C. Determining code change status
+Based on the matching of Changedev and Changellm, we determine each code change status as follows (example changes refer to 2(c)).
   • Correct change: LLM’s change is correct, either exactly like the developer (e.g., D1L2), or with an alternative API (e.g., D2L4), or with same APIs but with some refactoring (e.g., D3L8).
   • Incorrect change: The LLM incorrectly implemented the migration change: Used an incorrect API (e.g., get() instead of post() in L0), did not attempt to migrate it at all, or incorrectly removed part of a code.
+D. Determining migration status
 Based on the individual matched code changes, we now automatically determine an overall migration status:
@@
 We assign this status when there are no unmatched Changedev or Changellm left in any of the files after the matching process.
-3.2.2 RQ-A1 Findings: migration level correctness
+3.1.3 Evaluating LLM migrations using unit tests
+In RQ-1.2, we assess the same LLM PyMigBench migrations from RQ-1.1, but using the unit tests available in the corresponding repositories.
+A test-based evaluation that runs the code ensures that the migration preserves the original behavior.
+We now explain the process below.
+We make a copy of the repository after the developers’ migration (Codedev).
+We then make another copy of developer’s migration, but replace all Filedev with Filellm; we refer to this as Codellm.
+The idea is that Codellm is basically the version of Codedev where the LLM did the migration on behalf of the developer.
+To set up a virtual environment, we need to identify the Python version used in the client repository.
+If this information is available in the setup.py file, we use that version.
+Otherwise, we resolve the version based on the migration commit date as follows.
+We first identify the release dates of all minor versions of Python (3.6, 3.7 etc).
+Then, we find the latest release date that is before the migration commit date.
+This is the latest Python version that was available at the time of the migration; we use this version to create the virtual environment.
+Next, we install the code dependencies using pyproject.toml, setup.py and requirements files.
+In cases where specific dependency versions are not specified in those files, we look up the version history of the dependency on PyPI and install the latest version available at the migration commit date, similar to how we resolve the Python version.
+This ensures that the dependencies are compatible with the code at the time of migration.
+Once the virtual environment is ready, we run all the unit tests while measuring coverage in the repository on Codedev.
+If the run has errors, we read the error log and try to fix the errors, and run the tests again.
+Note that we only fix configuration or environment-related errors, not code errors.
+For example, if a dependency is missing, we install it; if the project requires a specific Python version, we set up the virtual environment with that version.
+We maintain a configuration file where we record any project-specific configurations or commands we used.
+We include this information in our artifact.
+We find that 218 out of the 314 migrations have at least one test.
+Among these, 172 migrations had unresolvable errors.
+The failures are primarily due to missing dependencies, especially for older projects.
+The remaining 46 migrations have tests that we are able to successfully run on Codedev.
+However, for the tests to be useful to validate migration, they must cover the migration-related code changes.
+Using the coverage report, we find that only 27 migrations have at least one test that covers the code changes, and among them, 25 migrations have at least one test that passes on Codedev.
+Thus, the evaluation in RQ-1.2 focuses only on these 25 migrations.
+We run the entire test suite on Codellm for these 25 migrations for each model using the same randomly selected LLM run used in RQ-1.1.
+We compare the test results between Codedev and Codellm and assign the following statuses to the migrations:
+  • Correct: All passing tests in Codedev also pass in Codellm.
+  • Partially correct: Some of the tests that pass in Codedev pass in Codellm, but the others fail or raise run-time errors.
+  • Incorrect: None of the passing tests in Codedev pass in Codellm; they all fail or raise run-time errors.
+3.2 RQ-1.1 How similar are the LLM migrations to the benchmark migrations?
+We now discuss the results of RQ-1.1.
+We first present the migration level results, then present the overall code change level results, and finally analyze the correctness of code changes by PyMigTax categories.
+3.2.1 Migration level correctness
 Table 1 shows the migration level results.
@@
 1.9% and 1.0%).
-RQ-A1, Migration Level: All three LLMs were able to perform correct migrations with 4o performing best: 94% of its migrations were partially correct and 57% were fully correct.
-3.2.3 RQ-A1 Findings: code change level correctness
-Table 2: Correctness of PyMigBench migrations (RQ1, migration level)
+RQ-1.1: Migration level All three LLMs were able to perform correct migrations with 4o performing best: 94% of its migrations were partially correct and 57% were fully correct.
+3.2.2 Overall code change level correctness
+Table 1: Correctness of PyMigBench migrations (RQ-1.1, migration level)
 Number (percentage) of migrations
@@
 Total migrations 314 (100%) 314 (100%) 314 (100%)
-We now present the correctness results at the individual code change level.
 The 314 migrations we use for our evaluation contain a total of 2,910 code changes.
@@
 This leaves us with 911, 1,966, and 1,996 code changes to evaluate for the three LLMs, respectively.
-3.2.3.1 Overall code change correctness
-Table 3: Correctness of PyMigBench migration-related code changes (RQ1, overall code change level)
+Table 2: Correctness of PyMigBench migration-related code changes (RQ-1.1, overall code change level)
 Number (percentage) of code changes
@@
 This suggests that while Mini attempted more code changes compared to Llama, the ones that Llama was able to attempt without failures or syntax errors were comparatively correct.
-3.2.3.2 Code change correctness by category
-Table 4: Correctness of code changes across PyMigTax categories (RQ-A1, code change level by category)
+3.2.3 Code change correctness by category
+Table 3: Correctness of code changes across PyMigTax categories (RQ-1.1, code change level by category)
 Migs = Migrations, LPs = Library Pairs, CCs = Code Changes, #Eval = Number of code changes evaluated, %Cor = percentage of correct code changes.
-PyMigTax category Frequency in PyMigBench LLama 3.1 GPT-4o mini GPT-4o Migs LPs $C$s #Eval %Cor #Eval %Cor #Eval %Cor Program elements Function call 229 (71%) 105 (78%) 1,804 (60%) 518 88% 1,093 87% 1,092 93% Import 318 (99%) 132 (99%) 732 (24%) 276 95% 580 94% 593 98% Decorator 39 (12%) 13 (10%) 411 (14%) 197 87% 335 90% 353 87% Attribute 54 (17%) 33 (25%) 183 (6.1%) 63 83% 103 78% 103 89% Type 17 (5.3%) 13 (10%) 55 (1.8%) 12 75% 34 85% 34 88% Exception 16 (5.0%) 13 (10%) 29 (1.0%) 17 65% 29 86% 29 93% Function ref 4 (1.2%) 4 (3.0%) 12 (0.4%) 3 33% 12 58% 12 100% Properties NoProps 57 (18%) 36 (27%) 262 (8.8%) 43 84% 150 88% 150 97% ElemNC 188 (59%) 88 (66%) 1,025 (34%) 398 84% 773 85% 771 90% ArgAdd 93 (29%) 49 (37%) 578 (19%) 150 85% 244 80% 244 88% ArgDel 66 (21%) 37 (28%) 349 (12%) 175 85% 234 87% 234 89% ArgTrans 81 (25%) 38 (28%) 317 (11%) 115 77% 201 76% 220 77% ParamAdd 13 (4.0%) 1 (0.7%) 127 (4.2%) 110 87% 127 95% 127 90% ArgNC 17 (5.3%) 11 (8.2%) 112 (3.7%) 20 60% 58 59% 58 84% AsyncTrans 13 (4.0%) 4 (3.0%) 50 (1.7%) 35 83% 50 90% 50 92% OutTrans 16 (5.0%) 15 (11%) 49 (1.6%) 25 88% 49 78% 49 92% Cardinality One-to-One 200 (62%) 98 (73%) 1,552 (52%) 408 86% 877 84% 894 91% One-to-Zero 50 (16%) 25 (19%) 297 (10%) 99 99% 290 99% 290 100% Zero-to-One 35 (11%) 18 (13%) 198 (6.6%) 34 97% 55 87% 55 89% Higher Card.
-91 (28%) 45 (34%) 210 (7.0%) 94 70% 164 79% 164 84% One-to-Many 52 (16%) 23 (17%) 142 (4.8%) 61 69% 103 78% 103 80% Many-to-One 30 (9.3%) 24 (18%) 41 (1.4%) 16 62% 34 82% 34 91% Many-to-Many 21 (6.5%) 11 (8.2%) 27 (0.9%) 17 82% 27 78% 27 89% Total 321 134 2,989 911 89% 1,966 89% 1,996 94%
+PyMigTax category Frequency in PyMigBench Llama 3.1 GPT-4o mini GPT-4o
+Migs LPs $C$s #Eval %Cor #Eval %Cor #Eval %Cor
+Program elements
+Function call 229 (71%) 105 (78%) 1,804 (60%) 518 88% 1,093 87% 1,092 93%
+Import 318 (99%) 132 (99%) 732 (24%) 276 95% 580 94% 593 98%
+Decorator 39 (12%) 13 (10%) 411 (14%) 197 87% 335 90% 353 87%
+Attribute 54 (17%) 33 (25%) 183 (6.1%) 63 83% 103 78% 103 89%
+Type 17 (5.3%) 13 (10%) 55 (1.8%) 12 75% 34 85% 34 88%
+Exception 16 (5.0%) 13 (10%) 29 (1.0%) 17 65% 29 86% 29 93%
+Function ref 4 (1.2%) 4 (3.0%) 12 (0.4%) 3 33% 12 58% 12 100%
+Properties
+NoProps 57 (18%) 36 (27%) 262 (8.8%) 43 84% 150 88% 150 97%
+ElemNC 188 (59%) 88 (66%) 1,025 (34%) 398 84% 773 85% 771 90%
+ArgAdd 93 (29%) 49 (37%) 578 (19%) 150 85% 244 80% 244 88%
+ArgDel 66 (21%) 37 (28%) 349 (12%) 175 85% 234 87% 234 89%
+ArgTrans 81 (25%) 38 (28%) 317 (11%) 115 77% 201 76% 220 77%
+ParamAdd 13 (4.0%) 1 (0.7%) 127 (4.2%) 110 87% 127 95% 127 90%
+ArgNC 17 (5.3%) 11 (8.2%) 112 (3.7%) 20 60% 58 59% 58 84%
+AsyncTrans 13 (4.0%) 4 (3.0%) 50 (1.7%) 35 83% 50 90% 50 92%
+OutTrans 16 (5.0%) 15 (11%) 49 (1.6%) 25 88% 49 78% 49 92%
+Cardinality
+One-to-One 200 (62%) 98 (73%) 1,552 (52%) 408 86% 877 84% 894 91%
+One-to-Zero 50 (16%) 25 (19%) 297 (10%) 99 99% 290 99% 290 100%
+Zero-to-One 35 (11%) 18 (13%) 198 (6.6%) 34 97% 55 87% 55 89%
+Higher Card. 91 (28%) 45 (34%) 210 (7.0%) 94 70% 164 79% 164 84%
+One-to-Many 52 (16%) 23 (17%) 142 (4.8%) 61 69% 103 78% 103 80%
+Many-to-One 30 (9.3%) 24 (18%) 41 (1.4%) 16 62% 34 82% 34 91%
+Many-to-Many 21 (6.5%) 11 (8.2%) 27 (0.9%) 17 82% 27 78% 27 89%
+Total 321 134 2,989 911 89% 1,966 89% 1,996 94%
 To understand if some types of code changes are more difficult to migrate, Table 3 shows the distribution of code change correctness across the different PyMigTax categories.
-The three columns in group “Frequency in PyMigBench” shows how often each PyMigTax category appears in all migrations, library pairs and code changes in PyMigBench, which provides perspective on the category’s impact in practice.
+The three columns in group “Frequency in PyMigBench” shows how often each PyMigTax category appears in all migrations (Migs), library pairs (LPs) and code changes (CCs) in PyMigBench, which provides perspective on the category’s impact in practice.
 This information is the same for each LLM since it describes the existing data in PyMigBench.
@@
 For example, the Function call row under the Program elements group shows that Llama’s syntactically correct migrations included 518 code changes that involve function calls, and that 88% of these code changes were correctly migrated.
-Program elements: Function calls are the most common program elements in terms of code changes (60%), and second most common in terms of migrations (71%) and library pairs (78%).
+A. Program elements
+Function calls are the most common program elements in terms of code changes (60%), and second most common in terms of migrations (71%) and library pairs (78%).
 All three LLMs correctly migrate a high proportion of these changes, with 88%, 87%, and 93% of the function calls being correctly migrated by the three LLMs.
@@
 While this last category shows most discrepancy between the models, this program element is present in only 0.4% of the code changes, therefore does not affect the overall performance significantly.
-Properties: The property no properties indicates that the source and target APIs are identical, and hence no changes are required.
+B. Properties
+The property no properties indicates that the source and target APIs are identical, and hence no changes are required.
 While 4o correctly leaves almost all of these APIs unchanged (97%), Llama and Mini incorrectly changed several of them, resulting in 84% and 88% correct, respectively.
@@
 The three LLMs correctly migrate only 76-77% of the argument transformation changes.
-We discuss specific failure instances in Section 3.4.
-The property parameter addition to decorated function is found in migrations in just one library, click library, where some decorators require adding a parameter to the decorated function libclick .
+The property parameter addition to decorated function is found in migrations in just one library, click library, where some decorators require adding a parameter to the decorated function Pallets (2024).
 All three LLMs perform well in migrating these changes; interestingly, Mini performs better than 4o (95% vs.
@@
 While 4o performs reasonably well in migrating these changes (84%), Llama and Mini fail frequently with them (only 60% and 59% correct).
-Cardinality: One-to-One code changes are the most common in PyMigBench, and all three LLMs perform well in migrating them.
+C. Cardinality
+One-to-One code changes are the most common in PyMigBench, and all three LLMs perform well in migrating them.
 One-to-Zero code changes are the ones where a source API needs to be removed, but no target needs to be added.
@@
 Interestingly, Llama performs the best in migrating these changes (97% correct), making this the only category where it performs better than the other models (87% and 89% correct).
-While previous research showed that higher-cardinality code changes are generally difficult compared to one-to-one code changes alrubaye2018automating ; wang2016transforming ; zhang2020deep ; HuangMappingAPI2024 , the PyMigBench dataset shows that higher cardinality changes are frequent with 34% of the library pairs requiring at least one higher-cardinality code change.
+While previous research showed that higher-cardinality code changes are generally difficult compared to one-to-one code changes Alrubaye and Mkaouer (2018); Wang et al.
+(2016); Zhang et al.
+(2020); Huang et al.
+(2024), the PyMigBench dataset shows that higher cardinality changes are frequent with 34% of the library pairs requiring at least one higher-cardinality code change.
 Overall, we find that the LLMs are reasonably successful at migrating them, with 70%, 79%, and 84% correctness rates.
-RQ-A1: Code Change Level LLMs correctly migrate a high proportion of most code changes, with 4o achieving 94% correct code changes.
+RQ-1.1: Code change level LLMs correctly migrate a high proportion of most code changes, with 4o achieving 94% correct code changes.
 The models perform reasonably well in migrating difficult higher-cardinality code changes (4o gets 84% correct), but struggle relatively more with argument transformation and argument name change.
-3.3 RQ-A2 How many migrations pass unit tests?
-We now assess the same LLM PyMigBench migrations from RQ1 but using the unit tests available in the corresponding repositories.
-A test-based evaluation that runs the code ensures that the migration achieves the expected behavior.
-Note that we should not expect that an LLM migration improves the rate of passing tests; a successful migration means that each test that previously passed on codedev must still pass on codellm.
-We now explain our evaluation set up.
-3.3.1 Preparing the code
-We make a copy of the repository after the developers’ migration (codedev).
-We then make another copy of developer’s migration, but replace all Filedev with Filellm; we refer to this as codellm.
-The idea is that codellm is basically the version of codedev where the LLM did the migration on behalf of the developer.
-3.3.2 Setting up the virtual environment
-To set up a virtual environment, we need to identify the Python version used in the client repository.
-If this information is available in the setup.py file, we use that version.
-Otherwise, we resolve the version based on the migration commit date as follows.
-We first identify the release dates of all minor versions of Python (3.6, 3.7 etc).
-Then, we find the latest release date that is before the migration commit date.
-This is the latest Python version that was available at the time of the migration; we use this version to create the virtual environment.
-Next, we install the code dependencies using pyproject.toml, setup.py and requirements files.
-In cases where specific dependency versions are not specified in those files, we look up the version history of the dependency on PyPI and install the latest version available at the migration commit date, similar to how we resolve the Python version.
-This ensures that the dependencies are compatible with the code at the time of migration.
-3.3.3 Running the tests and coverage
-Once the virtual environment is ready, we run all the unit tests while measuring coverage in the repository on codedev.
-If the run has errors, we read the error log and try to fix the errors, and run the tests again.
-We maintain a configuration file where we record any project-specific configurations or commands we used.
-We include this information in our artifact.
-We find that 218 out of the 314 migrations have at least one test.
-Among these, 172 migrations had unresolvable errors.
-The failures are primarily due to missing dependencies, specially for older projects.
-The remaining 46 migrations have tests that we are able to successfully run on codedev.
-However, for the tests to be useful to validate migration, they must cover the migration-related code changes.
-Using the coverage report, we find that only 27 migrations have at least one test that covers the code changes, and among them, 25 migrations have at least one test that passes on codedev.
-We run the entire test suite on codellm for these 25 migrations for each model using the same randomly selected LLM run used in RQ-A1.
-3.3.4 Determining migration status
-We compare the test results between codedev and codellm and assign the following statuses to the migrations:
-  • Correct: All passing tests in codedev also pass in codellm.
-  • Partially correct: Some of the tests that pass in codedev pass in codellm, but the others fail or raise run-time errors.
-  • Incorrect: None of the passing tests in codedev pass in codellm; they all fail or raise run-time errors.
-3.3.5 RQ-A2 Findings
-Table 5: Correctness of 25 PyMigBench migrations using their available unit tests (RQ2)
+3.3 RQ-1.2 How many migrations pass unit tests?
+Table 4: Correctness of 25 PyMigBench migrations using their available unit tests (RQ-1.2)
 Number (percentage) of migrations
@@
 Llama, on the other hand, has a much higher proportion of incorrect migrations (60%).
-RQ-A2 Out of 25 PyMigBench migrations with unit tests covering the migration-related code changes, the LLMs correctly migrated 36%-64%, with 4o being the highest.
-3.4 Implications and discussion
-3.4.1 Are LLMs suitable for library migration?
-We look into some specific examples to understand the promises and limitations of LLMs for library migration.
-3.4.1.1 The good
-We start with the promising side of using LLMs for library migration.
-In addition to the high overall correctness of individual code changes and migrations in RQ1, our evaluation reveals that the evaluated LLMs can migrate code changes that were traditionally known as difficult to migrate xu2019meditor .
-We discuss some examples.
-Different API styles: The library argparse libargparse provides API functions for parsing command line arguments, while click libclick provides decorators.
+RQ-1.2 Out of 25 PyMigBench migrations with unit tests covering the migration-related code changes, the LLMs correctly migrated 36%-64%, with 4o being the highest.
+3.4 Strengths of LLM for library migration
+Overall, the results of this empirical study highlight the high overall correctness of LLM migrations. We now discuss the key strengths we observe.
+Higher cardinality code changes
+As shown in RQ-1.1, Llama, Mini, and 4o successfully migrated 70%, 79%, and 84% of the higher cardinality code changes, respectively, which the literature always viewed as complex migrations Xu et al.
+(2019); Alrubaye et al.
+(2020).
+Figure 3 shows an example of a many-to-many migration from twitter to tweepy done by Llama.
+In addition to correctly replacing the functions Twitter(), and OAuth() with OAuthHandler(), set_access_token(), and API() with correct arguments, Llama also splits the code into multiple statements, making the code more readable.
+Before migration (using twitter)
+4tw_api = Twitter(
+5 auth = OAuth(oauth_token, oauth_secret,
+6 consumer_key, consumer_secret))
+7
+8 ÂX
+Llama’s migration (using tweepy)
+4auth = OAuthHandler(
+5 consumer_key, consumer_secret)
+6auth.set_access_token(
+7 oauth_token, oauth_secret)
+8tw_api = API(auth)
+Figure 3: Llama correctly migrating a many-to-many code change
+Different API styles
+We find that the LLMs are able to handle transformation between different API styles.
+For example, the library argparse provides API functions for parsing command line arguments, while click provides decorators.
 Figure 4 shows a migration example between these libraries.
 Despite the completely different API styles, the three LLMs correctly migrate 94% of code changes between this library pair.
-Figure 2: 4o correctly migrating from argparse to click, libraries using different API styles.
-Higher cardinality migrations As shown in RQ2, LLama, Mini, and 4o successfully migrated 70%, 79%, and 84% of the higher cardinality code changes, respectively, which the literature always viewed as complex migrations xu2019meditor ; alrubaye2020learning .
-Figure 3 shows an example of a many-to-many migration from twitter to tweepy done by LLama.
-In addition to correctly replacing the functions Twitter(), and OAuth() with OAuthHandler(), set_access_token(), and API() with correct arguments, LLama also splits the code into multiple statements, making the code more readable.
-twitter to tweepy, cloudbotirc/cloudbot/f8243223/twitter.py
-Figure 3: LLama correctly migrating a many-to-many code change
-Inferring changes outside of the source and target APIs The migration in Figure 4 adds the async_mode argument to the SocketIO() function.
+Before migration (using argparse)
+4def parse_commands():
+5 p = ArgumentParser()
+6 p.add_argument("src", type=Path)
+7 p.add_argument("dst", type=Path)
+8 p.add_argument("usedir",
+9 action = "store_true")
+10 return parser.parse_args()
+11
+12def main():
+13 args = parse_commands()
+14 convert(args.src, args.dst, args.usedir)
+4o’s migration (using click)
+4@argument("src",
+5 type=click.Path())
+6@argument("dst",
+7 type=click.Path())
+8@argument("usedir", type=bool)
+9def main(src, dst, usedir):
+10 convert(src, dst, usedir)
+11
+12
+13
+14ÂX
+Figure 4: 4o correctly migrating between libraries that use different API styles.
+Inferring changes outside of the source and target APIs
+We find that LLMs can infer some additional changes beyond the immediate APIs of the source or target library.
+For example, the migration in Figure 5 adds the async_mode argument to the SocketIO() function.
 This is interesting because the API is from the Flask-SocketIO library, which is not the source (eventlet) or target (gevent) libraries.
-However, the SocketIO function works with several libraries, eventlet being the default flasksocketio .
+However, the SocketIO function works with several libraries, eventlet being the default Flask (2024).
 This means, if 4o did not explicitly set the async_mode argument to gevent, the code would break after the migration.
+Before migration (using eventlet)
+3socketio = SocketIO(app)
+4bootstrap = Bootstrap(app)
+4o’s migration (using gevent)
+3socketio = SocketIO(app, async_mode="gevent")
+4bootstrap = Bootstrap(app)
 Figure 5: 4o performing a correct inferred change
-3.4.1.2 The Bad
-We also analyze cases where the LLMs did not correctly migrate the code and present notable observations.
-Failing to handle argument transformation Figure 5 shows an example where the source library retrying libretrying expects the time arguments in milliseconds, while the target library tenacity libtenacity expects them in seconds.
-The dataset includes 28 instances of this transformation, and all three LLMs failed to handle it in most cases, except for 4o, which handled three instances correctly, while the other two models correctly handled one instance each.
-This suggests that while the LLMs can map parameters, differences in expected format or units may not be properly managed.
-Figure 5: Mini filing in argument transformation.
-Replacing exception type We find cases where the LLM replaces generic exception types with specific ones, which can potentially fail to catch the same exceptions as the original code.
-This occurs in both migration-related and non-migration-related changes.
-For example, the LLM replaces an original Exception with ValueError.
-While using more specific exceptions is considered good practice cacho2014does , it does change the behavior of the code.
-Since modern development tools are often equipped to suggest better exception handling, it may be safer for an LLM-based migration tool to avoid changing existing exception types during a migration.
-3.4.1.3 The ugly
-We also observe the following problematic LLM behavior.
-Extra changes LLMs are known for not always following the instructions provided in the prompt ye2023cognitive .
-Despite our prompt explicitly asking the LLMs to avoid any refactoring or non-migration changes, we find cases where an LLM, particularly LLama, completely removes parts of the code that do not include any source APIs.
-We speculate that this behavior may stem from the LLM attempting to optimize output size due to token limitations.
-Silently deleting code can be dangerous for the overall functionality of the system, and led to test failures in RQ2 for many cases.
-Similarly, the LLMs sometimes refactor code that is not related to the migration, adopting a different style.
-While these improvements can be beneficial, they are unrelated to the migration task at hand.
-Future work could explore the use of an interactive IDE plugin that confirms the code changes before applying them, potentially allowing the developer to directly edit the changes.
-While less “dangerous”, we also find that the LLMs remove code comments and docstrings from the source code, even though the comments remain relevant after migration.
-Large file handling We observe that large files pose a challenge to LLama, given its smaller context size.
-Additionally, we observe that even the other two models often migrate an initial part of a large file but then incorrectly state that the remaining part does not require any changes.
-3.4.2 Differences in evaluation setup
-While the scope of RQ-A2 is limited due to the availability of the tests, we do find cases where the migrations are correct in RQ-A1 but incorrect in RQ-A2, or vice versa.
-To better understand these discrepancies, we manually inspect migrations with different results and provide some examples.
-3.4.2.1 Library version compatibility
-When evaluating RQ-A1, we mark a code change to be correct if it is correct as of the latest version of the library.
-However, in RQ-A2, we install the version the developers specified, and if it is not available, we resolve the library based on the migration commit date (see Section 3.3).
-Since we do not ask for a specific version number in the prompt, we notice that the LLMs usually pick APIs from the latest version of the target library.
-In some cases, this version discrepancy led to runtime errors when running tests in LLM’s code, which we report as incorrect in RQ-A2.
-Note that we consider version numbers moving forward when designing PyMigTool in Section 4.
-3.4.2.2 Function parameter shadowing import name
-ipaddr to ipaddress, google/capirca/eb768ea7/capirca/lib/nacaddr.py
-Figure 6: Example of function parameter shadowing import name.
-Notice in Figure 6 that the function parameter in the function IP on line 6 (Filepre) is named ipaddress, the same as the target library.
-The developer changed the parameter name to ip in Filedev.
-However, while Mini correctly replaces function IPNetwork with ip_network, it retained the original parameter name, causing it to shadow the new import name and leading to test failures.
-4o changed the parameter name to ipaddress_str to avoid the shadowing issue.
-3.4.2.3 Developer doing non-refactoring changes
-Recall that in RQ-A1, we find cases where the developer makes non-refactoring changes in the same commit as the migration while the LLM only does the migration it is expected to do.
-We count this as a correct migration in RQ-A1.
-However, given the changed behavior (which is often reflected in the tests), the tests fail in RQ-A2.
-Overall, the discrepancies show that both evaluation techniques provide different perspectives on the migration correctness.
-3.4.3 Understanding the cost of LLM-based migration
-The 314 migrations we use in our experiments involve a total of 258,198 lines of code, or 258.0 KLOC.
-The 10 runs per model makes it 2582.0 KLOC per model.
-The total cost of running our experiments was approximately \$15 for GPT-4o mini and \$420 for GPT-4o, which equates to \$0.01 and \$0.16 per KLOC, respectively.
-OpenAI has since reduced prices, so current costs would be lower.
-LLama 3.1 is free and open-source, though we used a pro subscription of Hugging Face (US\$9/month [^3] https://huggingface.co/pricing) to avoid any installation issues.
-While library migration is an infrequent task in a project’s lifetime, it requires significant developer effort when it happens.
-The cost of a developer’s time can easily exceed the cost of using an LLM, even GPT-4o, which can justify the cost of LLM-based tooling.
-3.5 Threats to validity
-Moha: I tried merging the two threats sections. But it because repetations of ”in phase one” and ”in phase two”. So I kept them separate for now.
-3.5.1 Internal validity
-Our evaluation involves two manual steps: validating correctness of code changes and labeling a non-migration related change as refactoring or non-refactoring.
-Manual activities are inherently subjective and may contain errors.
-We minimize this by having two authors independently perform the manual activities and resolving disagreements through discussion.
-3.5.2 Construct validity
-In RQ-A1, we analyze the code change categories based on Changedev, not on Changellm.
-The categories can differ if the LLM produces a different code change than the developer.
-Categorizing the LLM changes would require additional manual effort following the our previous work in PyMigTax pymigtax , which is beyond the scope of this work.
-That said, our observation is that the code/API differences in most cases are minor, and would not yield in a different PyMigTax category.
-Tests passing may not always guarantee migration correctness (e.g., the tests do not cover certain edge cases).
-We mitigate this, to the extent possible, by selecting repositories with high test coverage in LABEL:rq:unseen and ensuring that the tests cover the source API usages.
-3.5.3 External validity
-We use three off the shelf LLMs.
-While using an LLM specifically tuned for the library migration task could potentially yield better results, the goal of this initial empirical study is not to develop a migration tool, but to understand the challenges and opportunities of using LLMs for library migration.
-Our results may be viewed as a lower bound of the performance of potentially more specialized models.
-The libraries included in PyMigBench are all old-enough libraries where the LLMs are likely familiar with their APIs and API usage.
-This familiarity with APIs and API usage is precisely our hope when using LLMs for migration.
-However, being familiar with the APIs alone does not always guarantee being able to map them to each other or apply the right transformations on all code bases, which is why our evaluation is necessary.
-For newer libraries, additional information, e.g., API documentation, may be needed to guide the LLM in the migration process.
-This can be an interesting avenue for future work.
-4 Phase two: PyMigTool: a tool for end-to-end Python library migration
-In Section 3, we showed that LLMs can perform library migrations.
-However, we see two major challenges that make solely relying on LLMs in a practical tool ineffective: (1) We provide the LLM with only code files that directly use the source library.
-However, for a given project, there may be other files that depend on the migrated code and also require changes during the migration.
-(2) All LLMs often return only part of the code, skipping parts of the code that do not require migration.
-In this phase, we build and evaluate PyMigTool, a CLI tool that automates the full library migration process for a full project by combining the power of LLMs with pre and post processing to address the above challenges.
-PyMigTool first prepares the project that needs migration by setting up a virtual environment.
-It then runs an LLM to migrate the code, and finally performs post processing to improve the migration results.
-We run an extensive evaluation of PyMigTool on a set of 717 new, simulated migrations.
-Overall, PyMigTool is able to successfully migrate 32% of the migrations without any manual intervention.
-For a sample of the remaining not-fully-successful migrations, we find that a median of only 14% of the code lines in the project require manual changes to complete the migration, while the rest of the code is correctly migrated by PyMigTool.
-4.1 Building PyMigTool
-PyMigTool takes as input the following: (1) the path to a python project where the migration should happen, (2) the source library name, (3) the target library name and version, (4) the python version in which the code should run, (5) paths to one or more requirements files, (6) and an LLM to use for the migration.
-A key criteria for the input project is that it contains tests with high coverage, as PyMigTool uses the tests to identify the source library usage as well as to verify the migration.
+3.5 Challenges of using LLMs for library migration
+From the results of the empirical study in Phase 1 and the above strengths, we see that LLMs have strong potential for library migration.
+However, we also noted several challenges that would need to be addressed if LLMs are to be used for practical tooling for library migration.
+  1. Challenge-1
+    Files requiring migration may not be known upfront: In our experiment, we use PyMigBench to identify the files that require migration that we will send to the LLM.
+However, in a real-world scenario, the files requiring migration may not be known upfront.
+Therefore, a tool needs to first identify the files that require migration before sending them to the LLM for migration.
+  2. Challenge-2
+    LLMs do not always return complete code: A common issue we observe across all three LLMs is that they often do not return the complete code.
+Instead, they skips parts of the code, usually commenting that the next block of code remains unchanged.
+This is especially when the skipped part does not require migration, despite our prompt explicitly asking them to return the entire code.
+While this is not necessarily incorrect, as the skipped code may not require any changes, the resulting code is incomplete and cannot be provided directly to the developer.
+  3. Challenge-3
+    Dependents of migrated files may also require changes: In our experiment, we provide the LLMs only the files that directly use the source library, as marked in PyMigBench.
+However, other files that depend on the migrated files may also require changes.
+This is especially true when functions are made asynchronous during the migration, as all parts of the code that use those functions must also be made asynchronous.
+  4. Challenge-4
+    Library version compatibility: We do not mention any library version requirements in the prompt (1), and we notice that the LLMs usually pick the APIs from the latest or recent versions of the target library.
+In few cases, this caused runtime errors when running tests in LLM’s code, which we report as incorrect in RQ-1.2.
+This is because the application or the Python version used in the project may not be compatible with the latest version of the target library.
+In an actual migration scenario, the developers may want to migrate to a specific version of the target library that is compatible with their project.
+4 Phase 2: Building MigrateLib
+In Section 3, we show that LLMs can perform library migrations.
+However, we also identify several challenges when using LLMs for library migration, especially when building an end-to-end tool that developers can use.
+In this section, we describe Phase 2 of our work, where we build MigrateLib while addressing the identified challenges through pre- or post-processing steps as follows:
+  1. 1.
+    Before migration, MigrateLib identifies the files that use the source library, and thus require migration to address Challenge-1.
+  2. 2.
+    MigrateLib allows the user to specify the source and target library version, and include this in the prompt, to address Challenge-4.
+  3. 3.
+    After migration, MigrateLib automatically adds back any skipped code to address Challenge-2.
+  4. 4.
+    After migration, MigrateLib identifies the functions that become asynchronous due to the migration, and finds all other functions in the projects that use these functions, and makes them asynchronous as well to address Challenge-3.
+MigrateLib works as a Command Line Interface (CLI) and takes the following as input: (1) the path to a Python project that needs library migration, (2) the source library name (the version is inferred), (3) the target library name and version, (4) the Python version in which the code should run, (5) paths to one or more requirements files, (6) and an LLM to use for the migration.
+MigrateLib performs the migration in three main steps.
+First, it sets up a virtual environment for the project and identifies all files that require migration (Preparation step).
+In the next step, it prompts an LLM to migrate the identified files (LLM migration step).
+Finally, in the last step, it performs two post-processing steps to address Challenge-2 and Challenge-3 mentioned above to improve the migration results (Post-processing step).
+We note that a key criteria for the input project is that it contains tests with high coverage, as MigrateLib uses the tests to identify the source library usage as well as to verify the migration.
 While the migration will run regardless of the test coverage, migration results without test validation may not be as reliable.
-PyMigTool works in three main steps: Preparation round, LLM migration round, and Post-processing round.
-The preparation round sets up a virtual environment using the requirements files, and runs the tests to get the baseline results.
-The LLM migration round migrates the project to use the target library instead of the source library using the LLM the user specified.
-Finally, if the LLM migration is not fully successful, PyMigTool performs some non-LLM post processing steps to improve the migration results.
-4.1.1 Preparation round
-In the preparation round, PyMigTool first creates a virtual environment for the provided python version using the command python -m venv .venv, and then installs the required libraries using the command pip install -r <requirements-file-path> for each of the requirements file.
+4.1 Preparation step
+The preperation step handles three main preparation tasks needed before the actual migration:
+Task 1: Set up environment
+MigrateLib creates a virtual environment for the provided python version using the command python -m venv .venv, and then installs the required libraries using the pip install -r <requirements-file-path> command for each of the requirements file.
 It then installs the target library at the provided version using the pip install command.
-PyMigTool requires some additional packages, such as pytest, pytest-cov, which it also installs during this step.
-Once the environment is set up, PyMigTool runs the tests using the command python -m pytest on the project and saves the test report.
+MigrateLib requires some additional packages, such as pytest, pytest-cov, which it also installs during this step.
+After the installation, MigrateLib derives the source library version by querying the installed package using the pip show <source-library-name> command.
+Task 2: Run tests to get baseline
+Once the environment is set up, MigrateLib runs the tests using the command python -m pytest on the project and saves the test report.
 We refer to this version of the test report as TestReportpremig, which is used as the baseline for the migration.
-4.1.2 LLM migration round
-In our preliminary exploration of LLMs (Section 3), we use PyMigBench where the files that require migration were already known.
-However, in a real-world scenario for a full project, we do not know which files require migration.
-Therefore, PyMigTool first identifies the files that require migration, and then uses the LLM to migrate these files.
-4.1.2.1 Identifying files that require migration
-PyMigTool migrates one whole file at a time, therefore, it is sufficient to identify which files use the source library, without necessarily identifying the exact source library usage within these files.
-PyMigTool combines two techniques to identify the files that require migration: it uses static analysis to identify import statements and uses dynamic analysis to generate a run-time-generated call graph of the project using profiling data.
-geolib (library code)
-1class Square:
-2 def __init__(self, l):
-3 self.l = l
-4 self.name = "Square"
-5
-6 def area(self):
-7 return self.l ** 2
-serialize.py (uses library w/o import)
+Task 3: Identify files that require migration
+This task addresses Challenge-1 from Section 3, where we need to identify the files that require migration before sending them to the LLM.
+One obvious way a file can use a library is by importing it.
+For example, in Figure 6, the file main.py imports the geolib library, thus it requires migration when migrating from geolib to another library.
+To identify such files, MigrateLib reads and parses all Python files in the project and looks for import statements of the source library.
+Note that a library can be imported using a different name than the library name; a library may also have multiple import names.
+Therefore, MigrateLib first finds all import names of the source library using the Johnnydep library Jeantine-Glenn (2025), which extracts the import names from the wheel file of the library.
+A file may also use a library API without directly importing it.
+An example is shown in Figure 6.
+Notice that the to_dict() function in the serialize.py takes as input a shape object from the geolib library, even though it does not import the library.
+It uses the Shape.name attribute and Shape.area() method, which are part of the geolib library.
+To identify such files, MigrateLib builds a run-time call graph of the project by running the tests with profiling.
+It then analyzes the call graph to find all files that call any function from the source library.
+serialize.py
 1def to_dict(shape):
@@
 5 }
-main.py (uses library with import)
+6 X
+main.py
 1from serialize import to_dict
-2from geo import Square
+2from geolib import Square
 3import pprint
@@
 6pprint(to_dict(sq))
-serialize_test.py (test code)
-1from serialize import to_dict
-2from geo import Square
-3
-4def test_to_dict():
-5 square = Square(4)
-6 result = to_dict(square)
-7 assert result == {"name": "Square", "area": 16}
-Figure 7: Example of using a library with/without importing
-Detecting import statements: If a file in the project imports a library, PyMigTool considers that the file uses the library, and thus require migration.
-In Python, a library may be imported using a name different than the library name.
-For example, the library PyYAML can be installed using the command pip install pyyaml, but it has to be imported as import yaml.
-The former, pyyaml, is the library name, and the latter, yaml, is the import name.
-Furthermore, a library can have more than one import names.
-PyMigTool only takes the source library name as input, therefore, it needs to find the import names of the source library in the project.
-To do this, PyMigTool uses the Johnnydep tool libjohnnydep , which extracts the import names from the wheel file of the library.
-Next, PyMigTool uses the ast module to parse each Python file in the project and identify the import statements.
-It then compares the imported modules with the import names of the source library to identify the source library usage in the project.
-In the example in Figure 7, the file main.py directly imports the library geolib, thus PyMigTool will identify this file as requiring migration.
-However, the file serialize.py does not import the library, therefore, will not be identified as requiring migration by the static analysis.
-Detecting source library calls: In Figure 7, the file serialize.py does not import the library geolib, but both the main.py and serialize_test.py pass the geolib.Square object to the serialize.to_dict() function.
-That means when the code runs, the serialize.py file will use the geolib library, even though it does not import it.
-This shows that, identifying the import statements alone is not sufficient to identify all of the source library usage in the project.
-Accordingly, PyMigTool uses a run-time call graph to identify further source library usage in the project.
-PyMigTool runs the tests in the project with profiling using the pprofile package libpprofile .
-The profiling data includes all the calls made during the test execution, and generates a report in the CallGrind format doccallgrind .
-A call information includes the caller function and the callee function, as well as the line number where the call happened.
-A function is identified by the path of the file in which the function is defined, and the qualified name which is unique within the file.
-PyMigTool uses this data to build a call graph of the project.
-To know if a file requires migration, PyMigTool needs to determine whether any part of the file directly calls a source library function.
-A function can be defined in one of the three places: in python built-in modules, in the project, or in a third-party library.
-A Python system module file path is enclosed within angle brackets in the profile report; for example, <frozen os> indicates that the function is defined in the built-in os module.
-If a function is defined in a file under the project path (which PyMigTool takes as input), it considers this function to be in the project.
-PyMigTool also knows the path of the virtual environment where the third-party libraries are installed.
-Python virtual environments have a specific directory structure that PyMigTool leverages.
-For example, in Windows systems, the import name yaml would be in the <venv_path>/Lib/site-packages/yaml directory.
-Therefore, if a call is to a function that is defined in a file in this directory, PyMigTool can conclude that the function belongs to the yaml import name, i.e., the pyyaml library.
-In the call graph, we store the owner of each function, which can be project, system, or the import name of the library.
-Given this constructed call graph, PyMigTool iterates over each call.
-If the caller is owned by the project and the callee is owned by the source library, then the file that contains the caller function requires migration and PyMigTool adds it to the list of files that will be passed to the LLM.
-Overall, in Figure 7, PyMigTool will identify the file main.py requiring migration using static analysis, and serialize.py requiring migration using dynamic analysis.
-4.1.2.2 Migrating using an LLM
-PyMigTool can be configured to use any LLM that adheres to OpenAI’s APIs openaimodels , which has become a standard API for most models.
-For each file that requires migration, PyMigTool generates a prompt using the template in 2 and sends it to the LLM.
-The prompt is based on the prompt used in Section 4 (1), but enhanced with more specific instructions, and also includes the versions of the source and target libraries.
-Listing 2: Prompt template used by PyMigTool
-1The following Python code currently uses the library <source-lib> version<source-ver>.
-2Migrate this code to use the library <target-lib> version <target-ver> instead.
-3
-4**Instructions:**
-51. **Explain the Changes**: Begin the output with a brief explanation of the specific changes you made to migrate from "<source-lib>" to "<target-lib>".
-62. **Provide the Modified Code**: After the explanation, present the modified code. Provide the entire code after migration even if only a part of it is changed.
-7
-8**Important Guidelines**:
-9- Only make changes directly related to migrating between "<source-lib>" and "<target-lib>".
-10- Do not refactor, reformat, optimize, or alter the original coding style.
-11- The code given to you is part of a larger application. Do not change the names of classes, functions, or variables, because it can break the application.
-12
-13Original code:
-14<premig-code>
-The LLM response contains the migrated code, which PyMigTool extracts and saves.
-When the LLM finishes migrating all files, PyMigTool replaces the original files in the project with the migrated files.
+Figure 6: Example of using a library without importing
+Overall, in Figure 6, MigrateLib will identify the file main.py requiring migration using static analysis, and serialize.py requiring migration using dynamic analysis.
+4.2 LLM migration step
+MigrateLib can be configured to use any LLM that adheres to OpenAI’s APIs OpenAI (2024), which has become a standard API for most models.
+For each file that requires migration, MigrateLib generates a prompt using the template in 2 and sends the prompt to the LLM.
+The prompt is based on our original prompt used in Section 4 (1), but enhanced with more specific instructions.
+Especially, we include the versions of the source and target libraries to address Challenge-4 from Section 3.
+Listing 2: Prompt template used by MigrateLib
+The following Python code currently uses the library <source-lib> version
+<source-ver>. Migrate this code to use the library <target-lib> version
+<target-ver> instead.
+**Instructions:**
+1. **Explain the Changes**: Begin the output with a brief explanation of the
+specific changes you made to migrate from "<source-lib>" to "<target-lib>".
+2. **Provide the Modified Code**: After the explanation, present the modified
+code. Provide the entire code after migration even if only a part of it is
+changed.
+**Important Guidelines**:
+- Only make changes directly related to migrating between "<source-lib>" and
+"<target-lib>".
+- Do not refactor, reformat, optimize, or alter the original coding style.
+- The code given to you is part of a larger application. Do not change the names
+of classes, functions, or variables, because it can break the application.
+Original code:
+<premig-code>
+The LLM response contains the migrated code, which MigrateLib extracts and saves.
+When the LLM finishes migrating all files, MigrateLib replaces the original files in the project with the migrated files.
 It then runs the tests and generates a test report, which we refer to as TestReportllmmig.
-4.1.3 Post-processing round
-PyMigTool performs two post-processing steps we describe below.
-4.1.3.1 Merge skipped code
-During the the preliminary experiments (Section 3), we found that it is common for the LLMs to skip some parts of the code in the response, especially if those parts do not require migration, despite being instructed to return the full migrated code (see paragraph 3.4.1.3).
-In this post-processing step, PyMigTool identifies such skipped code, and merges it with the migrated code.
-For each file that was migrated in the LLM migration round, PyMigTool generates a diff between the original file and the migrated file using python’s difflib module.
-Then it attempts to detect which of the chunks were entirely removed from the original file.
-We make the following assumption to identify skipped code: 1.
-A skipped code must be mostly removal of lines.
-Because if the hunk has similar amount of additions and removals, it is likely that the hunk is replacing some code with other code, likely performing migration.
-2.
-The skipped code should not use source library APIs, because this makes the hunk migration-related otherwise.
-3.
-A skipped code should not be too small.
+4.3 Post-processing Step 1: Merge skipped code
+We design this post-processing step to address Challenge-2 from Section 3.
+The goal is to add back any unchanged code that LLM skipped returning in its response.
+For each file that was migrated in the LLM migration step, MigrateLib generates a diff between the original file and the migrated file using Python’s difflib module.
+Then it attempts to detect which of the hunks were entirely removed from the original file.
+We make the following assumption to identify a hunk of skipped code:
+  1. 1.
+    The hunk should not use source library APIs such that it is not a migration-related change.
+Therefore, MigrateLib checks if any of the removed lines in the hunk uses any source library API.
+If yes, the hunk is not considered skipped.
+  2. 2.
+    The hunk must mostly include removal of lines.
+Our reasoning is that if the hunk has a similar number of additions and removals, it is likely that the hunk is replacing some code with other code, likely performing migration.
+If a hunk has more than 90% of removed lines compared to total added+removed lines, we consider it mostly removal.
+  3. 3.
+    The hunk should not be too small.
 This assumption avoids incorrectly considering small migration-related removals as skipped code.
-We find the below specific baseline values for the above assumptions by trial-and-error to find a good balance between incorrectly merging migration related removals and incorrectly not merging skipped code:
-  • Small (0-9 lines of removals): this hunk is not considered skipped irrespective of the number of additions.
-  • Medium (10-19 lines of removals): if this hunk has at least one addition of line or use any source library APIs, it is not considered skipped.
-  • Large (20 or more lines of removals): if this hunk has less than 90% of removed lines compared to total added+removed lines, it is not considered skipped.
-The hunks contain the line numbers in the previous and new version of the file.
-Using the line numbers, PyMigTool inserts the removed lines in the migrated file, starting with the last removed hunk and going to the first removed hunk.
-The reverse order is important because inserting the lines would otherwise change the line numbers of the hunks that are inserted later.
-After inserting the skipped code, PyMigTool runs the tests again and generates a new test report, which we refer to as TestReportmerge.
-4.1.3.2 Infer async transformation
-In Python, a call to an asynchronous function must be awaited, i.e., preceded by the await keyword.
-Again, if a function call is awaited, the caller function must also be made asynchronous, i.e., it must be preceded by the async keyword.
-When there is a migration from a synchronous library to an asynchronous library, some functions in the migrated code may become asynchronous, which requires their transitive calls to be awaited, and transitive callers to be made asynchronous as well.
-In this step, PyMigTool identifies the functions that were made asynchronous by the LLM (asynced functions), then find all other functions that needs to be made asynchronous (to-async functions) and all calls that need to be made await (to-await calls), and applies the changes to the project.
-Additionally, PyMigTool also decorates the corresponding test functions with the pytest.mark.asyncio decorator, and installs the pytest-asyncio package, which are required to correctly run asynchronous tests.
+We consider a hunk as too small if it has less than 9 lines of code.
+We determine the above criteria thresholds through experimentation on a set of migration examples.
+If a hunk is identified as skipped code by the above criteria, MigrateLib adds back the removed lines from the original file to the migrated file.
+After finishing the merging of skipped code, MigrateLib runs the tests again and generates a new test report, which we refer to as TestReportmerge.
+4.4 Post-processing Step 2: Infer async transformation
+This post-processing step addresses Challenge-3 from Section 3 by identifying functions that become asynchronous due to the migration, and making their transitive callers asynchronous as well.
 Below we describe the steps in detail using the example in Figure 7.
 service.py
+service_test.py
+Before migration
 1import requests
 2
-3
-4def fetch_data(year):
+3def fetch_data(year):
 4 url = f"https://api.io/data/{year}"
@@
 8 return data
-service_test.py
 1from service import fetch_data
@@
 7 assert "days" in data
-8 assert isinstance(data["days"], list)
 8 assert len(data["days"]) == 365
-$\Downarrow$ LLM migration $\Downarrow$
+After LLM migration
 1from aiohttp import ClientSession
@@
 2
-3
-4def test_fetch_data():
+3def test_fetch_data():
 4 year = 2023
@@
 7 assert "days" in data
-9 assert isinstance(data["days"], list)
 8 assert len(data["days"]) == 365
-$\Downarrow$ Infer async transformation $\Downarrow$
+9
+10.
+After Infer async transformation
 1from aiohttp import ClientSession
 2
-3
-4async def fetch_data(year):
+3async def fetch_data(year):
 4 url = f"https://api.io/data/{year}"
-6
+5
 6 async with ClientSession() as session:
@@
 9 assert "days" in data
-10 assert isinstance(data["days"], list)
 10 assert len(data["days"]) == 365
@@
 Changes introduced by each migration step are highlighted in red.
-Identifying asynced functions For a given file that was migrated in the LLM migration round (Section 4.1.2), PyMigTool builds a abstract syntax tree (AST) of the code before migration and after migration.
-PyMigTool then finds the asynchronous function definitions in the code after migration.
-For each such function, PyMigTool finds functions with the same qualified name in the AST of the code before migration.
-If such a function exists in the AST before migration, but was not asynchronous, we consider the function as asynced.
-For example, in Figure 8, the function fetch_data is asynced, because it was synchronous before migration, but is asynchronous after migration.
-If a corresponding function is not found at all, it means that the function is newly added due to the migration, and we do not consider it as asynced, because it is introduced by the LLM and does not have callers in other parts of the project.
-Identifying to-async functions and to-await calls If a function a calls a function b, and b calls a function c, then a is said to transitively call b and c.
-For a given asynced function, PyMigTool finds all its transitive callers using the call graph, and marks them as to-async.
-For example, in Figure 8, the function fetch_data is asynced, and it is called by the function test_fetch_data, therefore, PyMigTool marks test_fetch_data as to-async.
-It then again uses the call graph to find all call sites of the asynced and to-async functions, and marks them as to-await.
-For example, in Figure 8, the call to fetch_data in the function test_fetch_data is marked as to-await, because it is a call to an asynced function.
-The to-async functions need the async keyword to be added to their signature, and the to-await calls need the await keyword to be added to them.
-Applying the changes PyMigTool uses LibCST liblibcst to parse the code, search code elements to be modified, and applies the following changes:
-  1. 1.
-    If a function definition is marked to-async, and it is not already asynchronous, then PyMigTool adds the async keyword to the function definition, for example the function definition test_fetch_data in Figure 8.
-  2. 2.
-    If a call is marked to-await, and it is already not awaited, PyMigTool adds the await keyword to the call, for example the call to fetch_data in test_fetch_data in Figure 8.
-  3. 3.
-    If a test function is asynced or is marked to-async, PyMigTool decorates the test function with pytest.mark.asyncio decorator, and installs the pytest-asyncio package to the virtual environment, which is required to run asynchronous tests.
-For example, in Figure 8, the function test_fetch_data is decorated with pytest.mark.asyncio.
-After finishing the async transfomations, PyMigTool runs the tests again and generates a new test report, which we refer to as TestReportasync.
-4.2 Evaluating PyMigTool
-We evaluate PyMigTool on a set of 717 simulated migrations in 175 unique projects.
-By simulated migrations, we mean that this migration did not exist in the project’s history; instead, we find source libraries used in a project and migrate the project to an analogous target library.
-We ensure that the projects we choose have high test coverage, so that the evaluation is reliable.
-We run each migration just one time, unlike in Section 3, because we found in that section that the LLM does not produce highly divergent migration results when run multiple times on the code.
-The setup we use is same as the setup in LABEL:rq:unseen (LABEL:sec:rq-llm-unseen), where we evaluate the LLMs on a set of migrations that were not used in training the LLMs.
-We aim to answer the following research questions:
-  1. RQ-B1
-    How many migrations can PyMigTool perform correctly? We measure how well PyMigTool performs on the simulated migrations without any manual intervention.
-  2. RQ-B2
+Identify asynced functions
+We call a function asynced if it was synchronous before migration, but is asynchronous after the LLM migration.
+For a given migrated file, MigrateLib creates an abstract syntax tree (AST) of the code before migration and another AST after migration.
+It then finds the function definitions that are not asynchronous in the AST before migration, but are asynchronous in the AST after migration.
+Since the location (line number) of the function definition may change due to the migration, MigrateLib matches the functions by their qualified name.
+In Figure 7, the function fetch_data is asynced.
+Identify to-async functions
+If a function foo calls a function bar, and bar calls a function fizz, then foo is said to transitively call bar and fizz.
+For a given asynced function, MigrateLib finds all its transitive callers using the call graph, and marks them as to-async, meaning that the async keyword must be added to their function definition.
+For example, in Figure 7, the function fetch_data in service.py is asynced after the LLM migration.
+This function is called by the function test_fetch_data in service_test.py, therefore, MigrateLib marks test_fetch_data as to-async.
+Identify to-await calls
+MigrateLib again uses the call graph to find all call sites of the asynced and to-async functions, and marks them as to-await, meaning that an await keyword must be added to these calls.
+For example, in the call to fetch_data in the function test_fetch_data (line 5, after LLM migration) is marked as to-await, because it is a call to an asynced function.
+Upon finding all asynced functions, to-async functions, and to-await calls, MigrateLib applies the necessary code transformations using the LibCST library contributors (2025).
+Additionally, MigrateLib decorates any asynced and to-async test functions with the pytest.mark.asyncio decorator, and installs the pytest-asyncio package to the virtual environment, which is required to run asynchronous tests.
+After finishing the async transformations, MigrateLib runs the tests again and generates a new test report, which we refer to as TestReportasync.
+5 Phase 3: Evaluating MigrateLib
+In this phase, we want to evaluate MigrateLib on a large set of real-world projects.
+To enable such large-scale evaluation, we rely on tests to determine if the migration is successful.
+Since the majority of repositories in PyMigBench did not contain tests, we resort to the idea of simulated migrations where we find source libraries used in a project and migrate the project to use an analogous target library.
+Our evaluation aims to answer the following research questions:
+  1. RQ-2.1
+    How many migrations can MigrateLib perform correctly? We measure how well MigrateLib performs on the simulated migrations without any manual intervention.
+  2. RQ-2.2
     How much manual effort is required to complete the incomplete migrations?
-While the best case scenario is that PyMigTool can migrate the project without any issues, a more practical expectation is that PyMigTool would do most of the task correctly, but there are some parts that require manual intervention.
-4.2.1 Data collection
-4.2.1.1 Selecting repositories
-We want to find repositories that are well covered by unit tests, so that we can automatically evaluate the migration correctness using the tests.
-We start by downloading a list of 95,472 Python GitHub repository metadata using SEART seart service.
+While the best case scenario is that MigrateLib can migrate the project without any issues, a more practical expectation is that MigrateLib would do most of the task correctly, but there are some parts that require manual intervention.
+5.1 Data collection and selected model
+We want to find repositories that use some source library, and then migrate them to use an analogous target library.
+To make the test-based evaluation reliable, we want to select repositories whose tests have good code coverage.
+We start by downloading a list of 95,472 Python GitHub repository metadata using SEART Dabic et al.
+(2021) service.
 We use the filters to select Python repositories that have at least one commit after January 1, 2024 to discard repositories that are not actively maintained.
 We also filter out repositories with less than 100 lines of code to ensure that the repositories have some code to migrate.
-For each repository, SEART metadata includes many details, including the repository name, the latest commit, the lines of code in the repository and repository size in KB.
-Before cloning the repository, we use some heuristics to filter out the repositories that may not be suitable for evaluation.
+We then filter out repositories using the following criteria before cloning them for further analysis:
   1. i.
@@
   2. ii.
-    No requirements files: We use requirements files to build the virtual environment.
+    No requirements files: We use requirements files to build the virtual environment to be able to run the tests.
 We filter out 39,916 repositories that do not have a requirements.txt file in the root directory, leaving us with 24,993 repositories.
-We check this by making an HTTP head request to the GitHub raw file URL.
-If the response code is not 200, we conclude that the file does not exist.
   3. iii.
-    Filtering by test files: Python test file names usually contain the term test.
-We use GitHub code search API [^4] GitHub code search API: https://docs.github.com/en/rest/search/search#search-code to query files that contains the term test, and filter out a repository if it does not have any such files.
+    No test files: Python test file names usually contain the term test.
+We use GitHub code search API [^1] GitHub code search API: https://docs.github.com/en/rest/search/search#search-code to query files that contain the term test, and filter out a repository if it does not have any such files.
 We filter out 12,558 repositories that do not have any test files, leaving us with 12,435 repositories.
@@
   1. iv.
-    Could not run tests: We attempt to setup a virtual environment for the repository using the requirements file and run the tests following Section 3.3.2.
-Due to various repository specific requirements; e.g., python version, system requirements, incomplete requirements file; we are often unable to run the tests.
+    Could not run tests: We attempt to set up a virtual environment for the repository using the requirements file and run the tests following Section 3.1.3 (Evaluating LLM migrations using unit tests).
+Due to various repository-specific requirements, e.g., Python version, system requirements, incomplete requirements file, we are often unable to run the tests.
 We filter out 10,104 repositories that we could not run the tests for, leaving us with 2,314 repositories.
@@
   4. vii.
-    Tests require user interaction: Some repositories require user interaction for the test to continue.
-For example, one repository requires a user to select an application to run a file.
-We identify 2 repositories that require manual intervention and exclude them, leaving us with 798 repositories, which we use for evaluation.
+    Tests require user interaction: We identify 2 repositories that require manual intervention and exclude them, leaving us with 798 repositories, which we use for evaluation.
 We store the commit SHA for the version of the repository we use to ensure reproducibility.
-Selecting library pairs We collect the dependencies of the 798 selected repositories by parsing their requirements files.
-These dependent libraries are the potential source libraries.
-Then for each source library, we find a list of candidate analogous libraries using GPT 4o mini, and then manually we manually verify them.
-We describe the process below.
-Collecting source libraries For each repository, we parse the requirements file to find the dependencies.
+We then collect the dependencies of the 798 selected repositories by parsing their requirements files.
 This gives us an initial list of 1,421 unique source libraries.
-We then look up the library metadata on libraries.io libraries.io .
-We exclude 12 libraries that are not found on libraries.io, leaving us with 1,409 libraries.
+We exclude 12 libraries that are not found on libraries.io Katz (2020), leaving us with 1,409 libraries.
 To ensure that the libraries are commonly known, we filter out 602 libraries that have less than 100 dependents, and filter out 2 libraries that do not have a description, leaving us with 805 potential source libraries.
 Listing 3: Prompt template for finding candidate analogous libraries
-1Identify alternative Python libraries that offer similar functionality as <lib_name>,
-2allowing for potential replacement in applications.
-3The alternative libraries should be available on PyPI.
-4Produce the result as a JSON list, having only the library names.
-5The produced library names should match the PyPI package names.
-Finding analogous library pairs For each of the 805 source libraries, we find a list of candidate analogous libraries using gpt-4o-mini-2024-07-18 with the prompt template in 3.
+Identify alternative Python libraries that offer similar functionality as
+<lib_name>, allowing for potential replacement in applications. The alternative libraries should be available on PyPI.
+Produce the result as a JSON list, having only the library names. The produced library names should match the PyPI package names.
+For each of the 805 source libraries, we find a list of candidate analogous libraries using gpt-4o-mini-2024-07-18 with the prompt template in 3.
 This gives us a total of 4,856 candidate library pairs.
 All of these pairs may not actually be analogous, i.e., they may not be suitable for migration, so we need to verify them.
-This manual task can be very time consuming, therefore, we first automatically filter out some of the pairs.
+This manual task can be very time-consuming; therefore, we first automatically filter out some of the pairs.
 First, we filter out 4,215 pairs where the source library is used in less than 10 of the selected repositories.
-We do this to ensure that every included library pair has at least 10 repositories to evaluate on, so that we can get a good estimate of the performance.
-Then, we filter out 192 pairs with unmaintained potential target library.
+This help us to keep the manual effort of evaluating analogous pairs manageable by only selecting pairs that appear in multiple projects in our data set.
+Then, we filter out 192 pairs with an unmaintained potential target library.
 Specifically, we exclude the target libraries that have not been updated since January 1, 2024.
 This leaves us with 449 pairs to manually verify.
-During manual evaluation, we look into the library documents and confirm 60 library pairs that are indeed analogous and can be potentially migrated.
-We exclude the remaining because they fall into one of the below categories:
-  1. i.
-    Not analogous: Some libraries are not analogous, e.g., they provide different functionality or have different APIs.
-  2. ii.
-    Test related libraries: For example, pytest, mock, etc.
+During manual validation, we look into the library documentation and confirm 60 library pairs that are indeed analogous and can be potentially migrated.
+We exclude the remaining because they fall into one of the following categories (i) Not analogous:Some libraries are not analogous, e.g., they provide different functionality or have different APIs.
+(ii) Test related libraries:For example, pytest, mock, etc.
 We exclude these libraries because our evaluation process relies on the tests.
-Therefore, migrating the test libraries themselves would may break our automated evaluation.
-  3. iii.
-    Non-API tools: Some dependencies are command line interfaces. Since they are not used in the code, we ignore them. For example, black, flake8, etc.
-  4. iv.
-    Documentation tools: Some dependencies are used for documentation generation. Since they are not used in the code, we ignore them. For example, sphinx, mkdocs, etc.
-4.2.1.2 Collecting migrations
-From the above process, we have a list of 798 repositories and 60 analogous library pairs.
-For each of these repositories, we now find which of the source libraries from the analogous library pairs are used in the repository.
-To find actual usage, not just listing in the requirements file, we parse the source code of the repository and look for import statements that use the source libraries.
+Therefore, migrating the test libraries themselves may break our automated evaluation.
+(iii) Non-API tools:Some dependencies are command line interfaces.
+Since they are not used in the code, we ignore them.
+For example, black, flake8, etc.
+(iv) Documentation tools:Some dependencies are used for documentation generation.
+Since they are not used in the code, we ignore them.
+For example, sphinx, mkdocs, etc.
+Now that we have a list of 798 repositories and 60 analogous library pairs, we want to find which repositories actually use the source libraries in our list.
+It could be the case that some repositories list a requirement that is not actually used in the code.
+Accordingly, we parse the source code of each repository and look for import statements that use the source libraries.
 We find that 175 repositories use at least one of the source libraries.
-Each source library can have one or more analogous target libraries.
-A repository, a commit, a source library and a target library thus form a unique migration instance.
-We get a total of 717 unique migration instances, which is the product of 175 repositories and 48 library pairs.
-These numbers are less than the 798 repositories and 60 analogous library pairs we started with, because we find many instances where a repository declares a dependency on a library, but does not actually use it in the code.
-To summarize, our evaluation is based on 717 migrations between 48 library pairs in 175 repositories.
+Note that each source library can have one or more analogous target libraries (e.g., requests→aiohttp and requests→httpx).
+We use the term unique migration instance to refer to the combination of a repository, a commit, a source library, and a target library.
+We use the latest commit of each repository at the time of our experiment (January 2025) for this purpose.
+We find a total of 717 unique migration instances, spanning 175 repositories and 48 library pairs that we use for our experiments.
 The 48 library pairs include 24 unique source libraries and 37 unique target libraries, and 55 unique libraries in total.
-4.2.2 Running migrations
-In our initial LLM experiment (Section 3), we see that GPT-4o performs better than the other two models. Therefore, we use GPT-4o (gpt-4o-2024-11-20) to run the migrations in this evaluation.
-4.2.3 RQ-B1: How many migrations can PyMigTool perform correctly?
-4.2.3.1 RQ-B1: Approach
-We describe in Section 4.1 that PyMigTool generates TestReportpremig before the migration, and three test reports after each migration step: TestReportllmmig after the LLM migration round, TestReportmerge after the merge skipped code, and TestReportasync after the infer async transformation.
-For a given post migration test report, we define correctness of a migration after that step as the proportion of tests that pass in the test report compared to the tests that passed in the test report before the migration.
-Table 6: Example of migration correctness calculation.
-Test case TestReportpremig TestReportllmmig TestReportmerge TestReportasync
-Test 1 Pass Pass Pass Pass
-Test 2 Pass Fail Fail Fail
-Test 3 Pass Fail Fail Pass
-Test 4 Pass Fail Pass Pass
-Test 5 Fail - - -
-# passing 4 1 2 3
-correctness 25% 50% 75%
-We use Table 6 to explain our correctness calculation.
-Assume that there are fives tests in the project.
-TestReportpremig has 4 passing tests and 1 failing test.
-We only take into account the tests that passed in TestReportpremig, because the failing tests cannot cannot indicate whether the migration is correct or not.
-In TestReportllmmig, only 1 test passes out of 4, therefore, the correctness is 25%.
-Similarly, in TestReportmerge and TestReportasync, 2 and 3 tests pass respectively, therefore, the correctness is 50% and 75% respectively.
-4.2.3.2 RQ-B1: Results
+In our initial LLM empirical study (Section 3), we see that GPT-4o performs better than the other two models. Therefore, we use GPT-4o (gpt-4o-2024-11-20) to run MigrateLib in this evaluation.
+5.2 Evaluation metrics
+Recall from Section 4 that MigrateLib generates TestReportpremig before the migration, and three test reports after each migration step: TestReportllmmig after the LLM migration step, TestReportmerge after the merge skipped code, and TestReportasync after the infer async transformation.
+For a given post-migration test report, we define the correctness of a migration after that step as the proportion of tests that pass in the test report compared to the tests that passed in the TestReportpremig.
+For example, if 80 tests passed in TestReportpremig, and 60 of those tests passed in TestReportllmmig, then the correctness after the LLM migration step is 75% (60/80).
+It is possible that some migrations may not be fully correct even after all the MigrateLib steps.
+We want to see how much migration work is still needed after MigrateLib migration to get a fully correct migration.
+Therefore, we randomly select one migration instance from each of the library pair, and attempt to manually fix them.
+After fixing a migration as much as possible, we run the tests again and generate a new test report, which we refer to as TestReportmanual.
+We then compute the correctness of the manually fixed version.
+For the migrations which become fully correct, i.e., 100% correctness, we estimate the amount of manual work needed to fix them to answer RQ-2.2.
+We use line number of modified lines to estimate the amount of manual change required to complete a migration.
+For a migration, we compute a diff between the code before manual migration and the code after manual migration.
+From the diff, we calculate the number of changed lines of code by adding number of removed lines and number of added lines.
+We refer to this number as $\mathit{MigLOC}_{\mathit{manual}}$.
+Now, this number gives us the absolute number of lines changed, therefore we cannot directly compare this number across different migrations.
+For example, a migration that required 50 changed lines where the tool performed 48 of them and we only have to edit 2 lines should not be considered the same as a migration that required 4 lines of code and we still have to edit 2 lines.
+Therefore, we also measure the number of lines changed by MigrateLib during the automated migration from the diff between the code before automatic migration and the code after migration by MigrateLib.
+We refer to this number as $\mathit{MigLOC}_{\mathit{auto}}$.
+Accordingly, the total number of lines that needed to be changed for the migration is the sum of these two numbers.
+We use these to normalize the $\mathit{MigLOC}_{\mathit{manual}}$ as follows:
+$\mathit{MigLOC}_{\mathit{manual}}^{\prime}=\frac{\mathit{MigLOC}_{\mathit{manual}}}{\mathit{MigLOC}_{\mathit{auto}}+\mathit{MigLOC}_{\mathit{manual}}}\times 100$ (1)
+Similarly, normalized automatic migration changes is calculated as the ratio of automatic changes to the total changes.
+$\mathit{MigLOC}_{\mathit{auto}}^{\prime}=\frac{\mathit{MigLOC}_{\mathit{auto}}}{\mathit{MigLOC}_{\mathit{auto}}+\mathit{MigLOC}_{\mathit{manual}}}\times 100$ (2)
+5.3 RQ-2.1: How many migrations can MigrateLib perform correctly?
 Figure 8: Distribution of correctness of migrations.
-The solid line indicate the median correctness, and the dashed line indicates the mean correctness.
+The dashed lines indicate the mean, the solid lines indicate the mean.
 Table 5: Comparison between migration steps
@@
 Average improvement in correctness 46% 65%
-Table 7 summarizes our results, while comparing the different migration steps.
-The table shows multiple metrics: the percentage of fully correct migrations is the proportion of migrations where all tests that passed before migration all pass after this migration step.
-Median correctness is the median correctness value across all evaluated migrations, where correctness is calculated according to Table 6.
-Number of migrations applied to is the number of migrations where this step was applicable to and got applied.
-For example, not all migrations require merging unchanged code or applying async transformations.
+Table 5 summarizes the results of RQ-2.1, while comparing the different migration steps.
+The table shows multiple metrics: the percentage of fully correct migrations is the proportion of migrations with 100% correctness.
+Median correctness is the median of correctness value across all evaluated migrations.
+Number of migrations applied to is the number of migrations where this step was applicable and thus got applied.
 Number of migrations improved shows the number of migrations where correctness improved in comparison to the previous step.
-Average improvement in correctness is calculated for those migrations where correctness improved.
-Figure 9 also shows the distribution of correctness of migrations after the each migration step.
-All three violins have concentrated values at 0% and 100%, indicating that most of the migrations either pass all tests or fail all tests.
-From this figure and Table 7, we see that the median correctness of migration increased from 65% in the LLM migration round to 80% after the merge skipped code, and further increased to 86% after the infer async transformation.
-The distributions show that after post-processing, there are fewer migration with zero or low correctness, and more migrations with high correctness.
+Average improvement in correctness is the mean of improvement for the migrations where correctness improved after a post-processing step.
+Note that the post-processing steps are applied only when the migrations are not fully correct and when that specific post-processing step is applicable.
+For example, as shown in Table 5, merge skipped code is applied to 260 migrations, among which 88 (34%) migrations have improved correctness with an average improvement of 46% per migration.
+On the other hand, infer async transformation is applied to 129 migrations, among which 45 (35%) migrations have improved correctness with an average improvement of 65% per migration.
+The frequency curve chart in Figure 8 shows the distribution of correctness of migrations after each migration step.
+All three distributions have concentrated values at 0% and 100%, indicating that most of the migrations either fail all tests or pass all tests.
+We see that the median correctness of migration increased from 65% in the LLM migration step to 80% after the merge skipped code, and further increased to 86% after the infer async transformation.
+Also note that the LLM-only has the highest value at 0% correctness, which decreases after each post-processing step.
+Similarly, the value at 100% correctness increases after each post-processing step.
+The distributions show that after post-processing, there are fewer migrations with zero or low correctness, and more migrations with high correctness.
 This indicates that the post-processing steps are effective in improving the correctness of migrations.
-Note that the post processing steps are applied only when the migrations are not fully correct, and when that specific post processing step is applicable.
-For example, merge skipped code is applied to 260 migrations, among which 88 (34%) migrations have improved correctness with an average improvement of 46% per migration.
-On the other hand, infer async transformation is applied to 129 migrations, among which 45 (35%) migrations have improved correctness with an average improvement of 65% per migration.
-4.2.4 RQ-B2: How much manual effort is required to complete the incomplete migrations?
-In RQ-B1, we see that PyMigTool is able to fully migrate 32% of the migrations; however, this means that 68% of the migrations would still require the developer to perform edits or fixes to complete the migrations.
-In this RQ, we look into how much additional manual effort is required to correct the incomplete migrations.
-4.2.4.1 RQ-B2: Approach
-RQ-B1 shows that 233 migrations are completely correct after the automated migration process.
-The remaining 484 migrations will require fixes for the tests to pass.
-These 484 migrations are between 46 unique library pairs in 147 repositories.
-Attempting to fix all these migrations manually would be very time consuming and impractical.
-Therefore, we select one random migration from each library pair, while maximizing the number of unique repositories.
-This gives us a total of 43 migrations between 43 library pairs in 39 unique repositories.
-For each of these migrations, we open the code using the PyCharm IDE at the version after PyMigTool migration.
-We refer to library documents and other online resources to understand each test error and how to fix it.
-After applying the edits, we runs the tests again and save the test report, which we refer to as TestReportmanual.
-We estimate the development effort required for a migration by measuring the number of lines changed for the migration.
-For a migration, we compute a diff between the code before manual edit (i.e., the code after automating migration by PyMigTool) and the code after manual edits.
-From the diff, we calculate the number of changed lines of code by adding number of removed lines and number of added lines.
-We refer to this number as $MigLOC_{manual}$.
-Now, while this number gives us the absolute number of line changed, it does not give us the relative effort in comparison to the automated migration by PyMigTool.
-For example, a migration that required 50 changed lines where the tool performed 48 of them and we only have to edit 2 lines should not be considered the same as a migration that required 4 lines of code and we still have to edit 2 lines.
-Therefore, we also measure the number of lines changed by PyMigTool during the automated migration from the diff between the code before automatic migration and the code after migration by PyMigTool.
-We refer to this number as $MigLOC_{auto}$.
-From these two numbers, we calculate the manual migration effort as the ratio of manual changes to the total changes.
-$\mathit{Effort}_{\mathit{manual}}=\frac{\mathit{MigLOC}_{\mathit{manual}}}{\mathit{MigLOC}_{\mathit{auto}}+\mathit{MigLOC}_{\mathit{manual}}}$ (1)
-Similarly, automatic migration effort is calculated as the ratio of automatic changes to the total changes.
-$\mathit{Effort}_{\mathit{auto}}=\frac{\mathit{MigLOC}_{\mathit{auto}}}{\mathit{MigLOC}_{\mathit{auto}}+\mathit{MigLOC}_{\mathit{manual}}}$ (2)
-Note that the total migration effort is the sum of $\mathit{Effort}_{\mathit{manual}}$and $\mathit{Effort}_{\mathit{auto}}$, i.e., $\mathit{Effort}_{\mathit{manual}}+\mathit{Effort}_{\mathit{auto}}=1$.
-4.2.4.2 RQ-B2: Results
-Figure 10: Distribution of effort savings across manually fixed migrations.
-The yellow dotted line shows median effort saving across all 24 migrations.
-We apply manual edits to the selected 43 migrations. The manual edit effort fixes at least one failing test for 24 (56%) migrations, including 21 (49%) migrations that were completely fixed.
-For the 21 migrations that we were able to fully correct through the manual edits, we measure $\mathit{Effort}_{\mathit{manual}}$and $\mathit{Effort}_{\mathit{auto}}$using Equation 1 and Equation 2, respectively.
+RQ-2.1 MigrateLib is able to fully migrate 32% of the migrations without any manual intervention.
+The median correctness of migrations improves from 65% with only the LLM migration step to 86% after the post-processing steps.
+5.4 RQ-2.2: How much manual effort is required to complete the incomplete migrations?
+RQ-2.1 shows that 233 migrations are completely correct after the automated migration process.
+The remaining 484 migrations will therefore require the developers to fix/complete them.
+To mimic what developers would do, we attempt to fix the migrations ourselves.
+However, the 484 migrations are between 43 unique library pairs in 147 repositories.
+To reduce the amount of manual work needed, while keeping the dataset diverse, we attempt to manually fix one incomplete migration from each library pair.
+Our selection results in a total of 43 migrations between 43 library pairs in 39 unique repositories, which we proceed to fix.
+The yellow dotted line shows median automatic change across all 24 migrations. The numbers on top of each bar show the total changed lines of code (automatic + manual) for that migration.
+Figure 9: Distribution of manual and automatic changes across manually fixed migrations.
+We are able to fix 24 out of 43 migrations.
+We could not fix the remaining 19 due to one or more of the following reasons: (1) As outsiders to the project, some of the application logic was too complex for us to understand and fix the migration.
+(2) In one case, the LLM migrated to dataclasses instead of cattrs, which made the migration incorrect.
+Fixing this migration would require undoing the entire migration and re-migrating it to the correct library, which is beyond the scope of fixing an incomplete migration.
+(3) In some cases, the tests themselves need to be updated to account for the behavior of the new target library.
+In several cases, we could not figure the correct expectation in the test oracle (again due to the complexity of the application) or because the original tests had complex mocking of the source library APIs.
+(4) We did observe cases where additional dependent libraries require migration, beyond the source-target pair.
+We consider such migrations beyond scope of this work, thus we did not attempt to fix such migrations.
+For the 24 migrations we were able to completely fix, we measure $\mathit{MigLOC}_{\mathit{manual}}^{\prime}$ and $\mathit{MigLOC}_{\mathit{auto}}^{\prime}$ using Equation 1 and Equation 2, respectively.
 Figure 9 shows the proportion of automatic and manual changes made for each migration.
-The blue part of the stacked bar is the $\mathit{Effort}_{\mathit{auto}}$, while red part is $\mathit{Effort}_{\mathit{manual}}$.
-The yellow dotted line shows the median of $\mathit{Effort}_{\mathit{auto}}$  across all 21 migrations, which is 86%.
-This means that for a median migration, PyMigTool performed 86% of the changes, while the developers only needed to perform 14% of the changes.
-4.3 Implications and discussion
-We now discuss some of the issues we noticed with the LLM-based migration in PyMigTool and potential extensions of PyMigTool to resolve these issues in the section after this (Section 4.3.2).
-4.3.1 Reoccurring issues in PyMigTool migrations
-To understand current limitations of PyMigTool, we look at the manual fixes we had to perform.
-The examples we show here are from the original migrations from our experiments, but often adopted for clarity and conciseness.
-In the left-to-right or top-to-bottom comparisons, the changes are highlighted in red.
-Original code
-1class EncAlchemyField:
-2 impl = BINARY
-3
-4 def __init__(self, key):
-5 self.key = key
-6 self.crypto=Crypto(key)
-PyMigTool migration
-1class EncTortoiseField:
-2 SQL_TYPE = "BLOB"
-3
-4 def __init__(self, key):
-5 self.key=key
-6 self.crypto=Crypto(key)
-Manual fix
-1class EncAlchemyField:
-2 SQL_TYPE = "BLOB"
-3
-4 def __init__(self, key):
-5 self.key=key
-6 self.crypto=Crypto(key)
-Figure 11: Example of PyMigTool renaming a public class, and we manually fixing it back to original name in a sqlalchemy→tortoise-orm migration.
-  1. Issue-1
-    The LLM refactors public program elements: In the LLM migration round, PyMigTool migrates each file in isolation, which can lead to incorrect migrations.
-For example, in Figure 11, PyMigTool renames the public class EncryptedAlchemyBinaryField defined in the client project to EncryptedTortoiseBinaryField to match the target library’s name.
+The blue part of the stacked bar is the $\mathit{MigLOC}_{\mathit{auto}}^{\prime}$, while the red part is $\mathit{MigLOC}_{\mathit{manual}}^{\prime}$.
+The yellow dotted line shows the median of $\mathit{MigLOC}_{\mathit{auto}}^{\prime}$  across all 24 migrations, which is 73%.
+This means that for a median migration, MigrateLib performed 73% of the changes, while the developers would only need to perform 27% of the changes.
+Note that 3 bars in the figure have no red part, this is because the manual fixes required to successfully complete the migration were not source code changes but some environment change, such as installing a missing dependency.
+RQ-2.2 For the migrations that required manual fixes, MigrateLib performed a median of 73% of the migration changes, while a developers would only need to perform 27% of the changes to complete the migration.
+To understand where MigrateLib fell short, we qualitatively analyze the manual fixes we had to perform.
+We categorize them into the following observations.
+We provide (adapted) examples from the actual migrations.
+  • The LLM often incorrectly uses target APIs: Sometimes the target API call used by the LLM does not correctly replace the source API call.
+This includes incorrect import statements, incorrect arguments to API calls, missing API calls, incorrect exception types, incorrect use of async/await, missing non-dafault argument, and other API related mistakes.
+For example, in a click→typer migration, the LLM did not use a name for the Typer() call, which was required.
+The equivalent Click API does not require a name, so MigrateLib missed adding it.
+  • The LLM skips part of the code: Despite explicitly mentioning in the prompt not to skip parts of the code (2), the LLM still skips part of the code in migrations, especially when the code pattern is repetitive or does not have source API usage.
+We observe the same behavior in Section 3, which inspired us to add the Merge skipped code post-processing step in MigrateLib to merge the skipped code with the migrated code (see Section 4.3).
+However, the heuristics used in Merge skipped code still fail to identify some skipped code so we had to manually add it back in the fix.
+  • MigrateLib struggles with complex async migrations: Some sync to async migrations require complex changes in the code structure, such as changing class constructors to async methods, changing sync context managers to async context managers.
+The LLM often struggles to perform such complex changes correctly.
+Our infer async transformation in tool only performs simple transformations, such as adding async/await keywords and changing context managers.
+We manually fixed some of these complex async migrations.
+  • The LLM refactors public program elements: In the LLM migration step, MigrateLib migrates each file in isolation, which can lead to incorrect migrations.
+For example, In the sqlalchemy→tortoise-orm migration, MigrateLib renames a public class EncryptedAlchemyBinaryField defined in the client project to EncryptedTortoiseBinaryField to match the target library’s name.
 This class is used in other parts of the code.
 Renaming the file prevents its callers from finding it, causing the errors in tests.
-Since PyMigTool does not pass the code that contains usage of the class in other files to the LLM, the LLM does not understand that it should not rename the class.
+Since MigrateLib does not pass the code that contains usage of the class in other files to the LLM, the LLM does not understand that it should not rename the class.
 Note that the prompt (2) does ask not to change any API names, however, the LLMs do not always strictly follow the instructions.
@@
     2gh_app_id = get_int_env_var("GH_APP_ID")
-    PyMigTool migration
+    MigrateLib migration
     1gh_actor = settings.get("GH_ACTOR", "nb")
@@
     4)
-    5def test_get_env_vars_optional(self):
-    6 expected_result = EnvVars(
+    5def test_get_env_vars(self):
+    6 expected_var = EnvVars(
     7 "testactor",
@@
     10 actual = str(result)
-    11 expected = str(expected_result)
-    12 self.assertEqual(actual, expected)
+    11 expected = str(expected_var)
+    12 assert actual == expected
     Figure 10: Example of requiring to call an additional API in a python-dotenv→dynaconf migration.
-  2. Issue-2
-    The LLM does not have enough context: Figure 12 shows an example where the underlying LLM was not able to add an additional API call (settings.reload()) required for correct migration.
+  • The LLM does not have the dependent code: Figure 10 shows an example where the underlying LLM was not able to add an additional API call (settings.reload()) required for correct migration.
 This is required because the target library, dynaconf, loads the environment variables when the library is imported for the first time, and a reload is required if the environment variables are changed.
 The original code uses python-dotenv to load the environment variables, which reads the environment variables at runtime.
-Notice how the test, which is in another file thus does not get passed to the LLM, updates the environment variables before running the tests, which is after the libraries are imported, causing the tests to fail without the settings.reload() call in the migrated code.
-If the LLM was given this test code that affects the migration-related code, it may have been able to understand that the reload is required.
-    Original code
-    1try:
-    2 d.update(toml.loads(r))
-    3except TypeError:
-    4 try:
-    5 d.update(toml.load(r))
-    6 except TypeError:
-    7 d.update(r)
-    PyMigTool migration
-    1try:
-    2 d.update(tomli.loads(r))
-    3except TypeError:
-    4 try:
-    5 d.update(tomli.load(r))
-    6 except TypeError:
-    7 d.update(r)
-    Manual fix
-    1try:
-    2 d.update(tomli.loads(r))
-    3except TypeError:
-    4 try:
-    5 d.update(tomli.load(r))
-    6 except AttributeError:
-    7 d.update(r)
-    Figure 13: Example of manually replacing exception type in a toml→tomli migration.
-  3. Issue-3
-    The LLM does not always adopt to the target library’s behavior: Figure 13 shows an example where the LLM did not recognize that the target library raises a different exception in a scenario, and therefore it was unable to replace the exception type correctly.
-  4. Issue-4
-    The LLM skips part of the code: Despite explicitly mentioning in the prompt not to skip parts of the code (2), the LLM still skips part of the code in migrations, specially when the code pattern is repetitive or does not have source API usage.
-We observe the same behavior in Section 3, which inspired us to add the Merge skipped code post-processing step in PyMigTool to merge the skipped code with the migrated code (see paragraph 4.1.3.1).
-Because this step uses some heuristics that often cannot identify all skipped code, we still had to manually fix some skipped code.
-4.3.2 Possible extensions of PyMigTool
-The problems we discuss above can potentially be mitigated by the following extensions of PyMigTool:
-  1. Extension-1
-    Post-processing for reverse refactoring: We can add a post processing step that detects renamed public APIs, and renames them back to the original names, potentially solving Issue-1.
-Refactoring detection tools, such as RefactoringMiner RefactoringMiner2 , can be used to detect such refactoring changes, and refactoring tools, such as RefMerge refmerge , can be used to apply the changes.
-  2. Extension-2
-    Including all dependant code with the prompt: This will allow the LLMs to know whether and how the code to migrate is used in other parts of the project, including the test code, potentially resolving Issue-2.
-This approach, however, can be impractical for the files that have a large number of dependencies, due to the LLM’s context length limitations.
-  3. Extension-3
-    Reprompting with test error messages: When we manually fixed the migration in Figure 13 related to Issue-3, we were able to quickly identify the issue just by looking at the test error messages.
-This means, systematically reprompting PyMigTool with the test error messages may help it to understand the context better and produce a correct migration, similar to how SOAR ni2021soar uses error messages to improve their program synthesis process (see LABEL:sec:relatedwork:transformation).
-While this can potentially help PyMigTool to produce more correct migrations, the challenge is processing the error messages and generating a meaningful prompt for the LLM.
-Note that SOAR considers only two pairs of libraries, and therefore the error messages are not as diverse as in our case.
-  4. Extension-4
-    Designing a sophisticated merge skipped code algorithm: our merge skipped code algorithm only looks at the size of the skipped code and usages of source APIs (see paragraph 4.1.3.1).
-A more sophisticated algorithm that considers the context and the structure of the code, such as removed functions/classes, may potentially improve this step reducing the effect of Issue-4.
-While the above steps can potentially mitigate the problems we identified, the positive side is that manually fixing the incomplete migrations was also simple, so further effort in more automation may not be necessary.
-For example, in Figure 13 (Issue-3), we were able to catch the issue easily just by looking at test error messages, and the fix was just replacing the exception type in the code.
-While reprompting with error messages may potentially fix the problem, it would also increase the cost of migration.
-The skipped code cases (Issue-4) are also fairly simple to resolve with modern IDEs.
-We use PyCharm’s change viewer during our manual fix, which allows us to easily identify the skipped code, and merge them with a click of mouse.
-The case in Figure 12 (Issue-2) required us to look deep into the library documentation and was time consuming.
-However, the original developers are likely to be more familiar with the libraries and their behavior.
-Therefore, they may have been be able to fix the issues easily.
-In some cases where migrations are initially marked as incorrect, the issues are related to configuration rather than code changes.
+Notice how the test, which is in another file that does not get passed to the LLM, updates the environment variables before running the tests.
+This updating happens after the library is imported, causing the tests to fail without the explicit settings.reload() call in the migrated code.
+If the LLM knew about the dependent code, it may be able to better understand the context of the migration.
+  • The LLM is unaware of configuration requirements: In some cases, the migration requires changes in the project configuration rather than code changes.
 For example, in a migration from chardet→cchardet, adjusting the Python version from 3.11 to 3.9 resolved the test failures.
 Similarly, we made a migration from flask→sanic successful by installing an additional dependency, sanic-ext, with no changes to the migrated code.
-The developers of the projects are likely to have been aware of such configuration requirements, therefore, they may be been able to fix the issues easily.
-Overall, we find that PyMigTool automatically completes migration for a large portion of cases (32% migrations) and does most of the complex migration work (86% changes) for the remaining migrations, often leaving simple edits for the developers to perform (14% changes).
+The LLM is unaware of such configuration requirements, and therefore, it cannot perform such changes.
+  • Often test expectations need to be updated: There are cases where the tests require changes due to the target library’s different behavior.
+For example, often the source library raises a generic exception, which the tests expect, while the target library raises a more specific exception.
+Another example is coloroma uses integers to represent colors (green is 32), while rich uses strings (green is "[green]").
+  • MigrateLib does not handle mock migrations: Some tests mock the source library’s API calls in the tests, and these mocks need to be migrated to the target library’s API calls.
+This is something we did not consider when designing MigrateLib, and therefore, it does not handle such cases.
+  • MigrateLib does not support multi-library migrations: We design MigrateLib to migrate from just one source library to just one target library, which may not always be sufficient.
+The project miguelgrinberg/microblog-api is such an example, which uses flask and some other libraries that depend on flask.
+To fully migrate to fastapi (or any other target library), we also need to migrate all these dependent libraries.
+6 Discussion
+Overall, we find that MigrateLib automatically completes migration for a large portion of cases (32% migrations) and does most of the complex migration work (73% changes) for the remaining migrations, often leaving simple edits for the developers to perform (27% changes).
 Further improvements in migration process may potentially help in reducing the manual effort.
 However, the manual effort seems fairly minimal in most cases, specially for the developers who are familiar with the libraries and the project.
-We stress that the author of this thesis, who has no prior knowledge of the projects, performed all the manual edits alone without help of any other developers.
-This indicates that the manual edit part can potentially be done by someone new to the project, or even be outsourced outside of the project team.
-4.4 Threats to validity
-This section discusses the potential threats to the validity of our study and the steps we take to mitigate them.
-4.4.1 Internal validity
-One author performs all manual fixes and their evaluation, which may introduce subjectivity or bias.
-To mitigate this, we follow a consistent process for manual fixes and rely on objective test outcomes as the primary measure of correctness.
-Configuration and environment setup also pose internal threats.
-Some migrations initially fail due to environment or dependency issues rather than code errors.
-We address these by adjusting configurations when possible, but environment-specific factors, such as Python and library version, operating system, may still affect the results.
-4.4.2 External validity
-We focus on Python projects that are well-tested and actively maintained, as selected through specific filters.
+We now discuss some potential improvement and future research and development directions for MigrateLib based on our findings.
+The two post-processing steps currently included in MigrateLib shows that simple automated steps can significantly improve the correctness of migrations.
+However, we also discuss in Section 5.4 find that these implementations have limitations, therefore can be further improved.
+MigrateLib can also include a reverse-refactoring post-processing step to detect and revert unwanted refactorings.
+There are existing tools that can detect refactorings Tsantalis et al.
+(2022) and apply refactoring changes Ellis et al.
+(2022), which can be used to implement this step.
+We provide the LLM with a fairly straight-forward context – the file that uses the source library APIs.
+This is often insufficient, and on the contrary, often too much if the file is large but only a small part of it needs migration.
+Identifying and sending more relevant part of the code may help the LLM to perform better migration.
+This however, comes with the cost of more pre-processing to identify the relevant code, as well as more post-processing to replace old code with the new code in the project.
+Another potential improvement may be iteratively re-prompting the LLM with additional context, such as test error messages after running the tests on the migrated code.
+The challenge here is to identify meaningful error messages and generate a prompt that can help the LLM to understand the context better.
+Trying out different re-prompting strategies for this may be an interesting future research direction.
+7 Threats to validity
+7.1 Internal validity
+There were several manual steps involved in our evaluations.
+Manual activities are inherently subjective and may contain errors.
+In Phase 1, we manually validated the correctness of code changes and labeled any non-migration related change as refactoring or non-refactoring.
+To minimize subjectivity, two authors independently perform these steps and resolved disagreements through discussion.
+In Phase 3, only one author performs all manual migrations.
+However, we rely on objective test outcomes as the primary measure of correctness.
+The merge skipped code uses some thresholds to identify skipped code hunks, which we determine through experimentation.
+Different thresholds may yield different results.
+However, we believe our chosen thresholds strike a reasonable balance between adding back genuinely skipped code and avoiding incorrect additions.
+If the assumptions do not hold, the resulting migration may potentially fail, therefore, the results we have may be a lower bound of the actual migration correctness.
+7.2 External validity
+We use three off the shelf LLMs in Phase 1, and the best of them in Phase 3.
+Using an LLM specifically tuned for the library migration task could potentially yield better results.
+Therefore, our findings in this paper may be viewed as a lower bound of the performance of potentially more powerful and specialized models.
+In Phase 3, we focus on Python projects that are well-tested and actively maintained, as selected through specific filters.
 This focus may limit the generalizability of our findings to projects with less test coverage, different maintenance status, or other programming languages.
 The set of library pairs and migration scenarios is also limited by the availability of analogous libraries and the manual validation process.
-As a result, the observed effectiveness of PyMigTool may not fully generalize to all migration contexts.
+As a result, the observed effectiveness of MigrateLib may not fully generalize to all migration contexts.
 That said, the size of the dataset (717 migrations) and the diversity of selected projects (175 repositories) provide a reasonable basis for evaluating the tool’s effectiveness in a variety of migration scenarios.
-4.4.3 Construct validity
-A key threat to construct validity is the correctness of the automated and manual evaluation process.
-We use automated test suites to assess migration correctness, but these tests may not cover all possible behaviors or edge cases in the migrated projects.
+The libraries we consider in Phase 1 and Phase 3 are relatively popular and well-established.
+This familiarity with APIs and API usage is precisely our hope when using LLMs for migration.
+However, being familiar with the APIs alone does not always guarantee being able to map them to each other or apply the right transformations on all code bases, which is why our evaluation is necessary.
+For newer libraries, additional information, e.g., API documentation, may be needed to guide the LLM in the migration process.
+This can be an interesting avenue for future study, and thus improvement of MigrateLib.
+7.3 Construct validity
+In RQ-1.1, we analyze the code change categories based on Changedev, not on Changellm.
+The categories can differ if the LLM produces a different code change than the developer.
+PyMigBench already provides us with the Changedev categories; categorizing the LLM changes according to PyMigTax Islam et al.
+(2024) requires additional manual effort that is beyond the scope of this work.
+That said, our observation is that the code/API differences in most cases are minor, and would not yield in a different PyMigTax category.
+We use automated test suites to assess migration correctness in Phase 3, but these tests may not cover all possible behaviors or edge cases in the migrated projects.
 To minimize this, we select projects with at least 80%lines of code covered by tests, which helps ensure that the tests are comprehensive enough to provide a reliable measure of correctness.
-Our manual migration effort calculation metric (Equation 1) is based on lines of code changed, which may not capture all aspects of migration effort, such as time spent understanding or debugging the migration.
-We use it as a measurable objective proxy for the effort required.
+We use migration lines of code (Equation 1) to estimate the amount of migration work, which may not capture all aspects of migration work, such as time spent understanding or debugging the migration.
+We choose to use this metric since it is a measurable objective proxy for the effort required.
 8 Related Work
@@
 The majority of the existing library migration techniques focus on identifying API mappings between two analogous libraries.
-Some of these techniques use mine existing library migrations to identify the API mappings teyton2012mining ; teyton2013automatic ; alrubaye2019use ; alrubaye2019migrationminer ; alrubaye2020learning .
-Some techniques see the requirement of existing migrations as a limitation, and therefore use other techniques, such as machine learning alrubaye2018automating ; miningAnalogicalAPIs and natural language processing ni2021soar .
+Some of these techniques mine existing library migrations to identify the API mappings Teyton et al.
+(2012, 2013); Alrubaye et al.
+(2019b, a, 2020).
+Some techniques see the requirement of existing migrations as a limitation, and therefore use other techniques, such as machine learning Alrubaye and Mkaouer (2018); Chen et al.
+(2019) and natural language processing Ni et al.
+(2021).
 These techniques only identify the API mappings, but do not perform client code transformation, which is a separate challenge.
-SOAR ni2021soar uses API documentation to identify API mappings, and then uses the mappings and unit tests to guide program synthesis for client code transformation.
+SOAR Ni et al.
+(2021) uses API documentation to identify API mappings, and then uses the mappings and unit tests to guide program synthesis for client code transformation.
 SOAR has several limitations that prevent it from being a general solution for library migration.
@@
 Zhou et al.
-hapim proposed HaPiM, which first trains an unsupervised machine learning model capable of API mapping, and then uses the API mappings to guide an LLM in code transformation.
+Zhou et al.
+(2023b) proposed HaPiM, which first trains an unsupervised machine learning model capable of API mapping, and then uses the API mappings to guide an LLM in code transformation.
 Similar to Chen et al.
-miningAnalogicalAPIs , HaPiM learns from API usage patterns of source and target libraries in their respective client projects, therefore does not require any existing migration examples.
+Chen et al.
+(2019), HaPiM learns from API usage patterns of source and target libraries in their respective client projects, therefore does not require any existing migration examples.
 For the code transformation, they use the API mappings to prompt an LLM to replace source API calls with target API calls.
-Their approach outperforms MigrationMapper alrubaye2019use and GPT-3.5 Turbo on the BLEU bleu and CodeBLEU codebleu metrics evaluated on 5 Java library pairs.
+Their approach outperforms MigrationMapper Alrubaye et al.
+(2019b) and GPT-3.5 Turbo on the BLEU Papineni et al.
+(2002) and CodeBLEU Ren et al.
+(2020) metrics evaluated on 5 Java library pairs.
 A significant limitation of HaPiM is that it works with small independent code snippets that use the source library APIs, and results in an equivalent snippet that uses the target library APIs.
@@
 Nikolov et al.
-nikolov2025google describe Google’s experiences in using LLMs for several internal code transformation tasks, including Joda Time→Java Time migrations.
+Nikolov et al.
+(2025) describe Google’s experiences in using LLMs for several internal code transformation tasks, including Joda Time→Java Time migrations.
 They do not attempt to provide a general solution for library migration; rather, they use LLMs to perform the migration they need.
@@
 Nonetheless, they show that using LLMs reduced migration time by at least 50% compared to manual migration.
+There are also a few studies that focus on migrating between different versions of the same library.
+The study of Nikolov et al.
+Nikolov et al.
+(2025) we describe above also includes JUnit 3→JUnit 4 migrations.
+Almeida et al.
+Almeida et al.
+(2024) evaluate the performance of ChatGPT for SQAlchemy 1→SQLAlchemy 2 migrations.
+While similar, migrating between different versions of the same library is different from migrating between different libraries, because the former can leverage the evolution or release history of the libraries to identify API changes.
 The above shows that there have been different approaches to automate the library migration process.
 However, these techniques come with one or more limitations that prevent them from being a general solution for library migration.
-Our tool, PyMigTool, resolves these limitations and can automate the entire library migration process for any arbitrary analogous library pair (Section 4).
-We evaluate PyMigTool on a large and diverse dataset of 717 migrations between 48 analogous library pairs, and show that PyMigTool can successfully perform most of the migration with no or minimal manual intervention from the developer.
+Our tool, MigrateLib, resolves these limitations and can automate the entire library migration process for any arbitrary analogous library pair.
+We evaluate MigrateLib on a large and diverse dataset of 717 migrations between 48 analogous library pairs, and show that MigrateLib can successfully perform most of the migration with no or minimal manual intervention from the developer.
 9 Conclusion
-In this paper, we explore the capabilities of LLMs for automating Python library migration and build PyMigTool, a practical tool that combines LLMs with static and dynamic program analysis to automate the full migration process.
-In the first phase of the work, we conduct an empirical study using three LLMs on the PyMigBench dataset to understand how well LLMs can perform library migration.
-We find that the best performing LLM, 4o, correctly migrates 94% of the individual code changes.
-At the migration level, it perfectly migrates 57% of the migrations, and 94% of them have at least one correctly migrated code change.
-We also run unit tests for a subset of the migrations, and find that 64% of the migrations by 4o have the same sets of tests passing in the developer’s migration and the LLM’s migration.
-A much lower cost LLM Mini and the free LLama also perform relatively well, with both correctly migrating 89% of code changes.
-Overall, we find that LLMs can handle complex code changes, therefore, we conclude that LLMs are a promising approach for automating library migration.
-However, we also identify some challenges that need to be addressed to make LLM-based migration practical.
-We use these insights to build PyMigTool, a CLI tool that automates the full library migration process in phase two.
-Our evaluation on 717 simulated migrations across 175 diverse projects and 48 library pairs shows that PyMigTool can fully automate a large portion of migrations (32%), with most remaining cases requiring only minor manual intervention (median 14% changes).
+Our goal in this paper is to fully automate the Python library migration process.
+We systematically approached this goal by first conducting an empirical study to understand the capabilities of LLMs for automating Python library migration.
+Overall, we find that LLMs can handle complex code changes but we also identify some challenges that need to be addressed to make LLM-based migration practical.
+Based on these findings and insights, we built MigrateLib, a practical CLI tool that combines LLMs with static and dynamic program analysis to automate the full migration process.
+Our evaluation on 717 simulated migrations across 175 diverse projects and 48 library pairs shows that MigrateLib can fully automate a large portion of migrations (32%), with most remaining cases requiring only minor manual intervention (median 27% changes).
 The tool’s layered approach, including post-processing steps, improves LLM migration correctness (from 65% to 86%) and reduces developer effort.
-While PyMigTool is not able to handle every migration scenario perfectly, it lowers the barrier for large-scale library migration in real-world projects.
-Overall, PyMigTool represents a meaningful step toward making automated library migration accessible and effective for the software engineering community.
+Overall, MigrateLib lowers the barrier for large-scale library migration in real-world projects, demonstrating accessible and effective automated library migration.
 Acknowledgements.
-We thank Dr. Ildar Akhmetov, Northeastern University, for his contributions in to the early to the ideas of the work.
+We thank Dr. Ildar Akhmetov, Northeastern University, for his contributions to the early ideas of this work.
+Declarations
+Funding
+This work was supported by the Canada Research Chairs Program and NSERC CREATE.
+Author contributions
+Below are the contributions of each author –
+  • Mohayeminul Islam contributed to idea conceptualization, design and implementation of the MigrateLib, conducting the experiments, writing script for the experiments, manual evaluations in Phase 1, manual migration fix in Phase 3, and writing the manuscript.
+  • Ajay Kumar Jha contributed to idea conceptualization, design of the MigrateLib, manual evaluations in Phase 1, and writing the manuscript.
+  • May Mahmoud manual evaluations in Phase 1, and revising the manuscript.
+  • Sarah Nadi contributed to idea conceptualization, design of the MigrateLib, manual evaluations in Phase 1, and writing the manuscript. She is the Principal Investigator of the project.
+Data availability
+The source code of MigrateLib, the experimental data, and the detailed results of our experiments are available at the following repository: https://figshare.com/account/articles/29602517.
+Conflict of interest
+No conflict.
+Ethical approval/informed concent/clinical trial number
+Not applicable.